### v4

# A Survey of Large Language Models in Medicine: Progress, Application, and Challenge

Abstract
Large language models (LLMs), such as ChatGPT, have received substantial attention due to their capabilities for understanding and generating human language. While there has been a burgeoning trend in research focusing on the employment of LLMs in supporting different medical tasks (e.g., enhancing clinical diagnostics, and providing medical education), a comprehensive review of these efforts, particularly their development, practical applications, and outcomes in medicine, remains scarce. Therefore, this review aims to provide a detailed overview of the development and deployment of LLMs in medicine, including the challenges and opportunities they face. In terms of development, we provide a detailed introduction to the principles of existing medical LLMs, including their basic model structures, number of parameters, and sources and scales of data used for model development. It serves as a guide for practitioners in developing medical LLMs tailored to their specific needs. In terms of deployment, we offer a comparison of the performance of different LLMs across various medical tasks, and further compare them with state-of-the-art lightweight models, aiming to provide a clear understanding of the distinct advantages and limitations of LLMs in medicine. Overall, in this review, we address the following study questions: 1) What are the practices for developing medical LLMs 2) How to measure the medical task performance of LLMs in a medical setting? 3) How have medical LLMs been employed in real-world practice? 4) What challenges arise from the use of medical LLMs? and 5) How to more effectively develop and deploy medical LLMs? By answering these questions, this review aims to provide insights into the opportunities and challenges of LLMs in medicine and serve as a practical resource for constructing effective medical LLMs. We also maintain a regularly updated list of practical guides on medical LLMs at: https://github.com/AI-in-Health/MedLLMsPracticalGuide.
1 Introduction
The recently emerged general large language models (LLMs) zhao2023survey ; yang2023harnessing , such as PaLM paLM , LLaMA llama ; touvron2023llama2 , GPT-series GPT-3 ; gpt-4 , and ChatGLM du2022glm ; zeng2022glm , have advanced the state-of-the-art in various natural language processing (NLP) tasks, including text generation, text summarization, and question answering. Inspired by these successes, several endeavors have been made to adapt general LLMs to the medicine domain, leading to the emergence of medical LLMs medpalm ; medpalm2 . For example, based on PaLM paLM and GPT-4 gpt-4 , MedPaLM-2 medpalm2 and MedPrompt nori2023can have respectively achieved a competitive accuracy of 86.5 and 90.2 compared to human experts (87.0 wu2023pmc ) in the United States Medical Licensing Examination (USMLE) medqa . In particular, based on publicly available general LLMs (e.g. LLaMA llama ; touvron2023llama2 ), a wide range of medical LLMs, including ChatDoctor li2023chatdoctor , MedAlpaca han2023medalpaca , PMC-LLaMA wu2023pmc , BenTsao wang2023huatuo , and Clinical Camel toma2023clinical , have been introduced. As a result, medical LLMs have gained growing research interests in assisting medical professionals to improve patient care thirunavukarasu2023large ; patel2023chatgpt .
Although existing medical LLMs have achieved promising results, there are some key issues in their development and application that need to be addressed. First, many of these models primarily focus on medical dialogue and medical question answering tasks, but their practical utility in clinical practice is often overlooked thirunavukarasu2023large . Recent research and reviews thirunavukarasu2023large ; omiye2024large ; clusmann2023future have begun to explore the potential of medical LLMs in different clinical scenarios, including Electronic Health Records (EHRs) yang2022large , discharge summary generation patel2023chatgpt , health education abd2023large , and care planning medpalm2 . However, they primarily focus on presenting clinical applications of LLMs, especially online commercial LLMs like ChatGPT (including GPT-3.5 and GPT-4 gpt-4 ), without providing practical guidelines for the development of medical LLMs. Besides, they mainly perform case studies to conduct the human evaluation on a small number of samples, thus lacking evaluation datasets for assessing model performance in clinical scenarios. Second, most existing medical LLMs report their performances mainly on answering medical questions, neglecting other biomedical domains, such as medical language understanding and generation. These research gaps motivate this review that offers a comprehensive review of the development of LLMs and their applications in medicine. We aim to cover topics on existing medical LLMs, various medical tasks, clinical applications, and arising challenges.
As shown in Figure 1, this review seeks to answer the following questions. Section 2: What are LLMs? How can medical LLMs be effectively built? Section 3: How are the current medical LLMs evaluated? What capabilities do medical LLMs offer beyond traditional models? Section 4: How should medical LLMs be applied in clinical settings? Section 5: What challenges should be addressed when implementing medical LLMs in clinical practice? Section 6: How can we optimize the construction of medical LLMs to enhance their applicability in clinical settings, ultimately contributing to medicine and creating a positive societal impact?
For the first question, we summarize the principles of existing medical LLMs, detailing their basic structures, the number of parameters, and the data used for model development. Additionally, we provide insight into the construction process of these models. This information is valuable for researchers and medical practitioners looking to build their own medical LLMs tailored to specific needs, such as computational limits, private data, and local knowledge bases. For the second question, we have conducted an extensive review on the performances of existing medical LLMs across ten biomedical NLP (discriminative and generative) tasks. This analysis allows us to understand how medical LLMs outperform traditional medical AI models in different aspects. By showcasing their abilities, we aim to clarify the strengths that medical LLMs bring to the table when deployed in clinical settings. The third question focuses on the practical application of medical LLMs in clinical settings. We provide guidelines for seven clinical application scenarios, offering specific implementations of medical LLMs and highlighting which abilities are used for each scenario. The fourth question emphasizes the challenges that must be overcome when deploying medical LLMs in clinical practice. These challenges include hallucination (i.e. generation of coherent and contextually relevant but factually incorrect outputs) umapathi2023med ; rawte2023survey , ethical, legal, and safety concerns chatgptconcernmed . We also advocate a broader evaluation of medical LLMs, such as trustworthiness xie2023faithful for ensuring their responsible and effective use in clinical settings. For the final question, we look into future directions for developing medical LLMs, providing a guide for researchers and practitioners looking to advance this field and maximize the potential of medical LLMs in medicine.
{forest} for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center,, text width=2.7cm, s sep=6pt, calign=edge midpoint, , [Medical Large Language Models, fill=gray!45, parent [ Principles
§ 2, for tree= top_class [Pipeline, for tree=fill=red!45,encoder [Pre-training from Scratch, encoder_more [BioBERTlee2020biobert ; PubMedBERTgu2021domain ; ClinicalBERTalsentzer2019publicly et al., encoder_work] ] [Fine-tuning General LLMs, encoder_more [DoctorGLMxiong2023doctorglm ; HuatuoGPTzhang2023huatuogpt ; LLaVA-Medli2023llavamed et al., encoder_work] ] [Prompting General LLMs, encoder_more [DeID-GPTliu2023deid ; ChatCADwang2023chatcad ; Dr.Knowsgao2023leveraging ; MedPaLMmedpalm et al., encoder_work] ] ] [Data, for tree=fill=green!45,generation [Medical Knowledge Bases, generation_more [DrugBank; NHS Health; NHS Medicine et al., generation_work ] ] [Pre-training Data, generation_more [PubMedPubMed ; MIMIC-IIIjohnson2016mimic ; MedDialogchen2020meddialog et al., generation_work] ] [Fine-tuning Data, generation_more [Qilin-Medye2023qilin ; BianqueCorpuschen2023bianque ; MD-EHRwang2023clinicalgpt et al., generation_work] ] ] ] [, for tree= top_class [Downstream
§ 3, for tree=fill=pink!45,model [Discriminative Tasks, model_wide [ Question Answering; Entity Extraction; Relation Extraction; Text Classification; Natural Language Inference; Semantic Textual Similarity; Information Retrieval et al., model_work ] ] [Generative Tasks, model_wide [Text Generation; Text Summarization; Text Simplification et al., model_work] ] [Performance Comparisons, model_wide [GPT-3.5-turbo; GPT-4; SOTA Task-specific Fine-tuned Models; Human Experts, model_work] ] ] ] [, for tree= top_class [Applications
§ 4, for tree=fill=yellow!45,gpa [Medical Diagnosis, gpa_wide [Chatcadwang2023chatcad ; LLM4medicinethirunavukarasu2023large et al., gpa_work ] ] [Formatting and ICD Coding, gpa_wide [PLM-ICDhuang2022plm ; LLM-ICDong2023applying ., gpa_work ] ] [Clinical Report Generation, gpa_wide [GPT-4V for Medicalwu2023can ; Qilin-Med-VLliu2023qilin et al., gpa_work ] ] [Medical Education, gpa_wide [LLM for medical educationabd2023large et al., gpa_work ] ] [Medical Robotics, gpa_wide [GRIDni2023grid ; Nested -UISxia2023nested et al., gpa_work ] ] [Medical Language Translation, gpa_wide [Advant GLM for Medical Educationkarabacak2023advent et al., gpa_work ] ] [Mental Health Support, gpa_wide [ChatCounselorliu2023chatcounselor et al., gpa_work ] ] ] ] [, for tree= top_class [Challenges
§ 5, for tree=fill=red!45, finetuning [Hallucination, finetuning_wide [Med-haltumapathi2023med ; Selfcheckgptmanakul2023selfcheckgpt et al;, finetuning_work ] ] [Lack of Evaluation Benchmarks and Metrics, finetuning_wide [Truthfulqatruthfulqa ; HaluEvalhalueval et al., finetuning_work ] ] [Domain Data Limitations, finetuning_wide [Model Dementiashumailov2023model et al., finetuning_work ] ] [New Knowledge Adaptation, finetuning_wide [EditingLMMhoelscher2023detecting et al., finetuning_work ] ] [Behavior Alignment, finetuning_wide [Webgptwebgpt ; Chain of Hindsightchainofhindsight et al., finetuning_work ] ] [Ethical & Legal and Safety Concerns, finetuning_wide [Jailbrokenchatgptjailbreak ; Jailpromptshen2023anything et al., finetuning_work ] ] ] ] [, for tree= top_class [Future Directions
§ 6, for tree=fill=yellow!45,inference [New Benchmarks, inference_wide [Chen et al.chen2023extensive ; Chen et al.chen2023large , inference_work ] ] [Multi-modal LLM, inference_wide [Opportunitieshuang2023chatgpt , MedPaLM Mtu2023towards , LLaVA-Medli2023llavamed , Visual Med-AlpacaVisual-Med-Alpaca , Med-Flamingomoor2023medflamingo , Qilin-Med-VLliu2023qilin , inference_work ] ] [Medical Agents, inference_wide [Medagentstang2023medagents , inference_work ] ] [Interdisciplinary Collaborations, inference_wide [Collaborationthirunavukarasu2023large ; omiye2024large ; clusmann2023future ; Malpractice Riskmello2023chatgpt et al., inference_work ] ] ] ] ]
Figure 1: An overview of the practical guides for medical large language models.
By addressing these questions and providing a holistic perspective on medical LLMs, we hope to promote further research and exploration in this interdisciplinary field. With the rapid development of foundation models, the LLMs could significantly improve future clinical practice and medical discoveries for the benefit of society. It requires sustained interdisciplinary collaboration between clinicians and AI researchers, doctor-in-the-loop, and human-centered design, to address existing challenges with applying LLMs in the medical domain. The co-development of the collaboration for appropriate training data, benchmarks, metrics, and deployment strategies could enable faster and more responsible implementation of medical large language models than the current practice.
2 The Principles of Medical Large Language Models
Box 1 and Table LABEL:table:llm_summary briefly introduce the background of general LLMs zhao2023survey , e.g., GPT-4 gpt-4 . Table 1 summarizes the currently available medical LLMs according to their model development. Existing medical LLMs are mainly pre-trained from scratch, fine-tuned from existing general LLMs, or directly obtained through prompting to align the general LLMs to the medical domain. Therefore, we introduce the principles of medical LLMs in terms of these three methods: pre-training, fine-tuning, and prompting. Meanwhile, we further summarize the medical LLMs according to their model architectures in Figure 2.
2.1 Pre-training
Pre-training typically involves training an LLM on a large corpus of medical texts, including both structured and unstructured text, to learn the rich medical knowledge. The corpus may include EHRs peng2023study , clinical notes yang2022large , and medical literature alsentzer2019publicly . In particular, PubMed PubMed , MIMIC-III clinical notes johnson2016mimic , and PubMed Central (PMC) literature PMC , are three widely used medical corpora for medical LLM pre-training. A single corpus or a combination of corpora may be used for pre-training. For example, PubMedBERT gu2021domain and ClinicalBERT are pre-trained on PubMed and MIMIC-III, respectively. In contrast, BlueBERT peng2019transfer combines both corpora for pre-training; BioBERT lee2020biobert is pre-trained on both PubMed and PMC. The University of Florida (UF) Health EHRs are further introduced in pre-training GatorTron yang2022large and GatorTronGPT peng2023study . MEDITRON chen2023meditron is pre-trained on Clinical Practice Guidelines (CPGs). The CPGs are used to guide healthcare practitioners and patients in making evidence-based decisions about diagnosis, treatment, and management.
To meet the needs of the medical domain, pre-training medical LLMs typically involve refining the following commonly used training objectives in general LLMs: masked language modeling, next sentence prediction, and next token prediction (Please see Box 1 for an introduction of these three pre-training objectives). For example. BERT-series models (e.g., BioBERT lee2020biobert , PubMedBERT gu2021domain , ClinicalBERT alsentzer2019publicly , and GatorTron yang2022large ) mainly adopt the masked language modeling and the next sentence prediction for pre-training; GPT-series models (e.g., BioGPT luo2022biogpt , and GatorTronGPT peng2023study ) mainly adopt the next token prediction for pre-training. After pre-training, medical LLMs can learn rich medical knowledge that can be leveraged to achieve strong performance on different medical tasks
Table 1: Summary of existing medical LLMs, in terms of their model development, the number (#) of parameters, the scale of pre-training/fine-tuning data, and the data source. M: million, B: billion.
Domains Model Development Models # Params Data Scale Data Source
Medical-domain LLMs (Sec. 2) Pre-training (Sec. 2.1) BioBERT lee2020biobert ; raj2019saama 110M 18B tokens PubMed PubMed
PubMedBERT gu2021domain 110M/340M 3.2B tokens PubMed PubMed
SciBERT beltagy2019scibert 110M 3.17B tokens Literature ammar2018construction
ClinicalBERT alsentzer2019publicly 110M 112k clinical notes MIMIC-III johnson2016mimic
MedCPT jin2023biocpt 330M 255M articles PubMed PubMed
BioGPT luo2022biogpt 1.5B 15M articles PubMed PubMed
BioMedLM venigalla2022biomedlm 2.7B 110GB PubMed pile
OphGLM gao2023ophglm 6.2B 20k dialogues MedDialog chen2020meddialog
GatorTron yang2022gatortron ; yang2022large 8.9B
>82B tokens
6B tokens
2.5B tokens+0.5B tokens
EHRs yang2022large
PubMed PubMed
Wiki+MIMIC-III johnson2016mimic
GatorTronGPT peng2023study 5B/20B 277B tokens EHRs peng2023study
MEDITRON chen2023meditron 70B 48.1B tokens
PubMed PubMed
Clinical Guidelines chen2023meditron
Fine-tuning (Sec. 2.2) DoctorGLM xiong2023doctorglm 6.2B 323MB dialogues CMD. Chinese-medical-dialogue-data
BianQuechen2023bianque 6.2B 2.4M dialogues BianQueCorpus chen2023bianque
ClinicalGPT wang2023clinicalgpt 7B
96k EHRs
192 medical QA
100k dialogues
MD-EHR wang2023clinicalgpt
VariousMedQA zhang2018multi ; medqa
MedDialog chen2020meddialog
Qilin-Med ye2023qilin 7B 3GB ChiMed ye2023qilin
ChatDoctorli2023chatdoctor 7B 110k dialogues
HealthCareMagic healthcaremagic
iCliniq icliniq
BenTsao wang2023huatuo 7B 8k instructions CMeKG-8K byambasuren2019preliminary
HuatuoGPT zhang2023huatuogpt 7B 226k instructions&dialogues Hybrid SFT zhang2023huatuogpt
Baize-healthcare xu2023baize 7B 101K dialogues
Quora+MedQuAD abacha2019question
MedAlpaca han2023medalpaca 7B/13B 160k medical QA Medical Meadow han2023medalpaca
AlpaCare zhang2023alpacare 7B/13B 52k instructions MedInstruct-52k zhang2023alpacare
Zhongjing yang2023zhongjing 13B 70k dialogues CMtMedQA yang2023zhongjing
PMC-LLaMA wu2023pmc 13B 79.2B tokens
Books+Literature lo2019s2orc
MedC-I wu2023pmc
CPLLM shoham2023cpllm 13B 109k EHRs
eICU-CRD pollard2018eicu
MIMIC-IV johnson2020mimiciv
Clinical Camel toma2023clinical 13B/70B
70k dialogues
100k articles
4k medical QA
ShareGPT sharegpt
PubMed PubMed
MedQA medqa
MedPaLM 2 medpalm2 340B 193k medical QA MultiMedQA medpalm2
BioMedGPT luo2023biomedgpt 7B >26B tokens S2ORC lo2019s2orc
Prompting (Sec. 2.3) CodeX lievin2022can GPT-3.5 / LLama-2 Chain-of-Thought (CoT) wei2023chainofthought -
DeID-GPT liu2023deid ChatGPT / GPT-4 Chain-of-Thought (CoT) wei2023chainofthought -
ChatCAD wang2023chatcad ChatGPT Zero-shot Prompting -
Dr. Knows gao2023leveraging ChatGPT Zero-shot Prompting UMLS bodenreider2004unified ; lindberg1993unified
MedPaLM medpalm PaLM (540B) 40 instructions MultiMedQA medpalm2
MedPrompt nori2023can GPT-4
Few-shot & CoT wei2023chainofthought
Figure 2: We adopt the data from Table 1 to demonstrate the development of model sizes for medical large language models in different model architectures, i.e., BERT-like, ChatGLM/LLaMA-like, and GPT/PaLM-like.
2.2 Fine-tuning
It is high-cost and time-consuming to train a medical LLM from scratch, due to its requirement of substantial (e.g. several days or even weeks) computational power and manual labor. One solution is to fine-tune the general LLMs with medical data, and researchers have proposed different fine-tuning methods medpalm2 ; han2023medalpaca ; toma2023clinical for learning domain-specific medical knowledge and obtain medical LLMs. Current fine-tuning methods include Supervised Fine-Tuning (SFT), Instruction Fine-Tuning (IFT), and Parameter-Efficient Tuning. The resulting fine-tuned medical LLMs are summarized in Table 1.
Supervised Fine-Tuning (SFT)
aims to leverage high-quality medical corpus, which can be physician-patient conversations li2023chatdoctor , medical question-answering han2023medalpaca , and knowledge graphs ye2023qilin ; wang2023huatuo . The constructed SFT data serves as a continuation of the pre-training data to further pre-train the general LLMs with the same training objectives, e.g. next token prediction. SFT provides an additional pre-training phase that allows the general LLMs to learn rich medical knowledge and align with the medical domain, thus transforming them into specialized medical LLMs.
The diversity of SFT enables the development of diverse medical LLMs by training on different types of medical corpus. For example, DoctorGLM xiong2023doctorglm and ChatDoctor li2023chatdoctor are obtained by fine-tuning the general LLMs ChatGLM du2022glm ; zeng2022glm and LLaMA llama on the physician-patient dialogue data, respectively. MedAlpaca han2023medalpaca based on the general LLM Alpaca alpaca is fine-tuned using over 160,000 medical QA pairs sourced from diverse medical corpora. Clinicalcamel toma2023clinical combines physician-patient conversations, clinical literature, and medical QA pairs to refine the LLaMA-2 model touvron2023llama2 . In particular, Qilin-Med ye2023qilin and Zhongjing yang2023zhongjing are obtained by incorporating the knowledge graph to perform fine-tuning on the Baichuan yang2023baichuan and LLaMA llama , respectively.
In summary, existing studies have demonstrated the efficacy of SFT in adapting general LLMs to the medical domain. They show that SFT improves not only the model’s capability for understanding and generating medical text, but also its ability to provide accurate clinical decision support he2023survey .
Instruction Fine-Tuning (IFT)
constructs instruction-based training datasets zhang2023instruction ; he2023survey ; zhao2023survey , which typically comprise instruction-input-output triples, e.g. instruction-question-answer. The primary goal of IFT is to enhance the model’s ability to follow various human/task instructions, align their outputs with the medical domain, and thereby produce a specialized medical LLM.
Thus, the main difference between SFT and IFT is that the former focuses primarily on injecting medical knowledge into a general LLM through continued pre-training, thus improving its ability to understand the medical text and accurately predict the next token. In contrast, IFT aims to improve the model’s instruction following ability and adjust its outputs to match the given instructions, rather than accurately predicting the next token as in SFT zhang2023instruction . As a result, SFT emphasizes the quantity of training data, while IFT emphasizes their quality and diversity. Since IFT and SFT are both capable of improving model performance, there have been some recent works yang2023zhongjing ; ye2023qilin ; zhang2023alpacare attempting to combine them for obtaining robust medical LLMs.
In other words, to enhance the performance of LLMs through IFT, it is essential to ensure that the training data for IFT are of high quality and encompass a wide range of medical instructions and medical scenarios. To this end, MedPaLM-2 medpalm2 invited qualified medical professionals to develop the instruction data for fine-tuning the general PaLM. BenTsao wang2023huatuo and ChatGLM-Med ChatGLM-Med constructed the knowledge-based instruction data from the knowledge graph. Zhongjing yang2023zhongjing further incorporated the multi-turn dialogue as the instruction data to perform IFT. MedAlpaca han2023medalpaca simultaneously incorporated the medical dialogues and medical QA pairs for instruction fine-tuning.
Parameter-Efficient Tuning
aims to substantially reduce computational and memory requirements for fine-tuning general LLMs. The main idea is to keep most of the parameters in pre-trained LLMs unchanged, by fine-tuning only the smallest subset of parameters (or additional parameters) in these LLMs. Commonly used parameter-efficient tuning techniques include Low-Rank Adaptation (LoRA) lora , Prefix Tuning li2021prefixtuning , and Adapter Tuning liu2022p ; liu2021p .
In contrast to fine-tuning full-rank weight matrices, 1) LoRA preserves the parameters of the original LLMs and only adds trainable low-rank matrices into the self-attention module of each Transformer layer lora . Therefore, LoRA can significantly reduce the number of trainable parameters and improve the efficiency of fine-tuning, while still enabling the fine-tuned LLM to capture effectively the characteristics of the downstream tasks. 2) Prefix Tuning takes a different approach from LoRA by adding a small set of continuous task-specific vectors (i.e. “prefixes”) to the input of each Transformer layer li2021prefixtuning ; zhao2023survey . These prefixes serve as the additional context to guide the generation of the model without changing the original pre-trained parameter weights. 3) Adapter Tuning involves introducing small neural network modules, known as adapters, into each Transformer layer of the pre-trained LLMs houlsby2019parameter . These adapters are fine-tuned while keeping the original model parameters frozen houlsby2019parameter , thus allowing for flexible and efficient fine-tuning. The number of trainable parameters introduced by adapters is relatively small, yet they enable the LLMs to adapt to downstream scenarios or tasks effectively.
In general, parameter-efficient tuning is valuable for developing LLMs that meet unique needs in specific (e.g. medical) domains, due to its ability to reduce computational demands while maintaining the model performance. For example, medical LLMs DoctorGLM xiong2023doctorglm , MedAlpaca han2023medalpaca , Baize-Healthcare xu2023baize , Zhongjing yang2023zhongjing , CPLLM shoham2023cpllm , and Clinical Camel toma2023clinical adopted the LoRA lora to perform parameter-efficient fine-tuning to efficiently align the general LLMs to the medical domain.
2.3 Prompting
Fine-tuning considerably reduces computational costs compared to pre-training, but it requires further model training and collections of high-quality datasets for fine-tuning, thus still consuming some computational resources and manual labor. In contrast, the “prompting” methods efficiently align general LLMs (e.g. PaLM paLM ) to the medical domain (e.g., MedPaLM medpalm ), without training any model parameters. Popular prompting methods include Zero/Few-shot prompting, Chain-of-Thought (CoT) prompting, and Prompt Tuning.
Table 2: Summary of existing widely-used biomedical NLP datasets for evaluating medical LLMs.
Types Tasks Datasets Data Scale Data Source
Discriminative Tasks (Sec. 3.1) Question Answering (Sec. A.1) MedQA (USMLE) medqa 61,097 multiple-choice QA pairs 51 medical textbooks
MedMCQA Medmcqa 194k multiple-choice QA pairs AIIMS and NEET PG entrance exam
MMLU hendrycks2020measuring 300 multiple-choice QA pairs Clinical Knowledge
PubMedQA PubmedQA 273,500 question-context-answer triples PubMed PubMed ; sen2008collective
BioASQ-QA krithara2023bioasq 4,721 QA pairs BioASQ tsatsaronis2015overview
EMRQA pampari2018emrqa 400k+ QA pairs i2b2
CliCR vsuster2018clicr 105k QA pairs 12k clinical case reports
COVID-QA moller2020covid 2,019 QA pairs 147 articles (CORD-19 wang2020cord )
MASH-QA zhu2020question 34,808 QA pairs WebMD website
Health-QA zhu2019hierarchical 7,517 questions and 7,355 articles 1,235 website Patient’s articles
Entity Extraction (Sec. A.2) NCBI Disease dougan2014ncbi 6,892 entity mentions 793 PubMed abstracts
JNLPBA collier2004introduction 59,963 entity mentions 2,404 abstracts (GENIA kim2003genia )
GENIA kim2003genia 96,582 entity mentions 2,000 MEDLINE abstracts
BC5CDR li2016biocreative 28,785 entity mentions 1,500 PubMed articles
BC4CHEMD krallinger2015chemdner 84,355 entity mentions 10,000 PubMed abstracts
BioRED luo2022biored 20,419 entity mentions 600 PubMed abstracts
CMeEE zhang2021cblue 24,850 entity mentions 938 files
ADE gurulingappa2012development 21,160 entity mentions 2,972 documents
2012 i2b2 sun2013evaluating 30,690 entity mentions 310 discharge summaries
2014 i2b2/UTHealth stubbs2015annotating 57,744 entity mentions 1,304 longitudinal medical records
2018 n2c2 henry20202018 83,869 entity mentions 505 discharge summariesjohnson2016mimic
Cadec karimi2015cadec 9,111 entity mentions 1,253 posts
PGR sousa2019silver 1712 abstracts 19511 entity mentions
EU-ADR van2012eu 7,011 entity mentions 100 MEDLINE abstracts
Relation Extraction (Sec. A.3) BC5CDR li2016biocreative 3,116 relations 1,500 PubMed articles
BioRED luo2022biored 6,503 relations 600 PubMed abstracts
ADE gurulingappa2012development 13,806 relations 2,972 documents
2018 n2c2 henry20202018 59,810 relations 505 discharge summaries
2010 i2b2/VA uzuner20112010 1,748 reports clinical reports
GDA wu2019renet 869,152 relations 30,192 PubMed titles and abstracts
DDI segura2013semeval 5,021 relations 1,017 biomedical texts
GAD becker2004genetic 5,937 records genetic records
2012 i2b2 sun2013evaluating 54,560 relations 310 discharge summaries
PGR sousa2019silver 4,283 relations 1,712 abstracts
EU-ADR van2012eu 2,436 relations 100 abstracts
Text Classification (Sec. A.4) ADE gurulingappa2012development 20,967 sentences 2,972 documents
2014 i2b2/UTHealth stubbs2015annotating 1,304 records longitudinal medical records
HoC baker2016automatic 5,310 sentences 1,499 PubMed abstracts
OHSUMED hersh1994ohsumed 50,216 abstracts 50,216 medical abstracts
WNUT-2020 Task 2 nguyen2020wnut 10,000 tweets 10,000 tweets
Medical Abstracts schopf2022evaluating 14,438 abstracts 28,880 medical abstracts
MIMIC-III johnson2016mimic 52,724 discharge summaries 53,423 hospital admissions
Natural Language MedNLI romanov2018lessons 14,049 sentence pairs MIMIC-III johnson2016mimic
Inference (Sec. A.5) BioNLI bastan2022bionli 40,459 sentence pairs PubMed abstracts
Semantic Textual Similarity (Sec. A.6) MedSTS wang2020medsts 174,629 sentence pairs 3M clinical notes
2019 n2c2/OHNLP wang20202019 2,054 sentence pairs 113k clinical notes and MedSTS wang2020medsts
BIOSSES souganciouglu2017biosses 100 sentence pairs Text Analysis Conference (TAC)
Information Retrieval (Sec. A.7) TREC-COVID voorhees2021trec 8,691 3-level query-document relations CORD-19 wang2020cord
NFCorpus boteva2016full 712k 3-level query-document relations NutritionFacts.org (NF) website
BioASQ (BEIR) thakur2021beir 32,916 binary query-document relations BioASQ tsatsaronis2015overview
Generative Tasks (Sec. 3.2) Text Summarization /Generation (Sec. B.1/Sec. B.3) MIMIC-CXR johnson2019mimic 128k Findings-Impression pairs radiology reports
MIMIC-III johnson2016mimic 73k Findings-Impression pairs radiology reports
PubMed PubMed ; sen2008collective 36M+ citations and abstracts biomedical literature
PMC PMC 9,407,149 papers biomedical and life sciences literature
CORD-19 wang2020cord 140k+ papers COVID-19 literature
MentSum sotudeh2022mentsum 24,119 post-TLDR summary pairs 24,119 Reddit posts
MeQSum abacha2019summarization 1,000 question-summary pairs U.S. National Library of Medicine
Text Simplification MultiCochrane joseph2023multilingual 7,755 abstract-summary pairs Cochrane Library systematic reviews
(Sec. B.2) AutoMeTS van2020automets 3,300 sentence pairs English Wikipedia kauchak2013improving
Zero/Few-shot Prompting
aims to directly give an instruction to prompt the LLM to efficiently perform a task following the given instruction. Zero-shot prompting does not provide an example. Few-shot prompting presents the LLMs with a small number of examples or task demonstrations before requiring them to perform a task. It allows the LLMs to learn from these examples or demonstrations to accurately perform the downstream task and follow the given examples to give corresponding answers GPT-3 . Therefore, few-shot prompting allows LLMs to accurately understand and respond to medical queries. For example, MedPaLM medpalm significantly improves the downstream performance by providing the general LLM, PaLM paLM , with a small number of downstream examples such as medical QA pairs.
Chain-of-Thought (CoT) Prompting
further improves the accuracy and logic of model output, compared with Zero/Few-shot Prompting. Specifically, through prompting words, CoT aims to prompt the model to generate intermediate steps or paths of reasoning when dealing with downstream (complex) problems wei2023chainofthought . Moreover, CoT can be combined with few-shot prompting by giving reasoning examples, thus enabling medical LLMs to give reasoning processes when generating responses. For tasks involving complex reasoning, such as medical QA, CoT has been shown to effectively improve model performance medpalm ; medpalm2 . Medical LLMs, such as DeID-GPT liu2023deid , MedPaLM medpalm , and MedPrompt nori2023can , use CoT prompting to assist them in simulating a diagnostic thought process, thus providing more transparent and interpretable predictions or diagnoses. In particular, MedPrompt nori2023can directly prompts a general LLM, GPT-4 gpt-4 , to outperform the fine-tuned medical LLMs on medical QA without training any model parameters.
Prompt Tuning
aims to improve the downstream model performance by employing both prompting and fine-tuning techniques lester2021power ; liu2021p . The prompt tuning method introduces learnable prompts, i.e. trainable continuous vectors, which can be optimized or adjusted during the fine-tuning process to better adapt to different downstream scenarios and tasks. Therefore, they provide a more flexible way of prompting LLMs than the “prompting alone” methods that use discrete and fixed prompts, as described above. In contrast to traditional fine-tuning methods that train all model parameters, prompt tuning only tunes a very small set of parameters associated with the prompts themselves, instead of extensively training the model parameters. Thus, prompt tuning effectively and accurately responds to medical problems nori2023can , with minimal incurring computational cost.
Existing medical LLMs that employ the prompting techniques are listed in Table 1. Recently, MedPaLM medpalm and MedPaLM-2 medpalm2 propose to combine all the above prompting methods, resulting in Instruction Prompt Tuning, to achieve strong performances on various medical question-answering datasets. In particular, using the MedQA dataset for the US Medical Licensing Examination (USMLE), MedPaLM-2 medpalm2 achieves a competitive overall accuracy of 86.5% compared to human experts (87.0%), surpassing previous state-of-the-art method MedPaLM medpalm by a large margin (19%).
Figure 3: Performance (Dataset-Metric (Task)) comparison between the GPT-3.5 turbo, GPT-4, state-of-the-art task-specific lightweight models (Fine-tuned), and human experts, on seven medical tasks across eleven datasets. All data presented in our Figures originates from published and peer-reviewed literature. Please refer to the appendix for the detailed data.
Figure 4: We demonstrate the development of medical large language models over time in different model development types through the scores of the United States Medical Licensing Examination (USMLE) from the MedQA dataset.
3 Medical Tasks
In this section, we will introduce two popular types of downstream tasks: generative and discriminative tasks, including ten representative downstream tasks that further build up clinical applications. Table 2 summarizes the details of widely used evaluation datasets for each downstream task. Figure 3 illustrates the performance comparisons between different LLMs. For clarity, we will only cover a general discussion of those downstream tasks. The detailed definition of the downstream task and the performance comparisons can be found in our supplementary material.
3.1 Discriminative Tasks
Discriminative tasks are for categorizing or differentiating data into specific classes or categories based on given input data. They involve making distinctions between different types of data, often to categorize, classify, or extract relevant information from structured text or unstructured text. The representative tasks (see Table 2) include Question Answering, Entity Extraction, Relation Extraction, Text Classification, Natural Language Inference, Semantic Textual Similarity, and Information Retrieval.
The typical input for discriminative tasks can be medical questions, clinical notes, medical documents, research papers, and patient EHRs. The output can be labels, categories, extracted entities, relationships, or answers to specific questions, which are often structured and categorized information derived from the input text. In existing LLMs, the discriminative tasks are widely studied, and used to make predictions and extract information from input text.
For example, based on medical knowledge, medical literature, or patient EHRs, the medical question answering (QA) task can provide precise answers to clinical questions, such as symptoms, treatment options, and drug interactions. This can help clinicians with making more efficient and accurate diagnoses medpalm ; medpalm2 ; thirunavukarasu2023large . Entity extraction can automatically identify and categorize critical information (i.e. entities) such as symptoms, medications, diseases, diagnoses, and lab results from patient EHRs, thus assisting in organizing and managing patient data. The following entity linking task aims to link the identified entities in a structured knowledge base or a standardized terminology system, e.g., SNOMED CT donnelly2006snomed , UMLS bodenreider2004unified , or ICD codes world1978international . This task is critical in clinical decision support or management systems, for better diagnosis, treatment planning, and patient care.
3.2 Generative Tasks
Different from discriminative tasks that focus on understanding and categorizing the input text, generative tasks require a model to accurately generate fluent and appropriate new text based on given inputs. These tasks include medical text summarization tang2023evaluating ; van2023clinical , medical text generation luo2022biogpt , and medical text simplification ondov2022survey .
For medical text summarization, the input and output are typically long and detailed medical text (e.g. “Findings” in radiology reports), and a concise summarized text (e.g., the “Impression” in radiology reports). Such text contains important medical information that enables clinicians and patients to efficiently capture the key points without going through the entire text. It can also help medical professionals to draft clinical notes by summarizing patient information or medical histories.
In medical text generation, e.g. discharge instruction generation liu2022retrieve , the input can be medical conditions, symptoms, patient demographics, or even a set of medical notes or test results. The output can be a diagnosis recommendation of a medical condition, personalized instructional information, or health advice for the patient to manage their condition outside the hospital.
Medical text simplification ondov2022survey aims to generate a simplified version of the complex medical text by, for example, clarifying and explaining medical terms. Different from text summarization, which concentrates on giving shortened text while maintaining most of the original text meanings, text simplification focuses more on the readability part. In particular, complicated or opaque words will be replaced; complex syntactic structures will be improved; and rare concepts will be explained joseph2023multilingual . Thus, one example application is to generate easy-to-understand educational materials for patients from complex medical texts. It is useful for making medical information accessible to a general audience, without altering the essential meaning of the texts.
3.3 Performance Comparisons
Figure 3 shows that some existing general LLMs (e.g. GPT-3.5-turbo and GPT-4 gpt-4 ) have achieved strong performance on existing downstream tasks. This is most noticeable for the QA task where GPT-4 (shown in the blue line in Figure 3) consistently outperforms existing task-specific fine-tuned models, and is even comparable to human experts (shown in purple line). The QA datasets of evaluation include MedQA (USMLE) medqa , PubMedQA PubmedQA , and MedMCQA Medmcqa . To better understand the QA performance of existing medical LLMs, in Figure 4, we further demonstrate the QA performance of medical LLMs on the MedQA dataset over time in different model development types. It also clearly shows that current works, e.g., MedPrompt nori2023can , have successfully proposed several prompting methods to enable LLMs to outperform human experts.
However, on the non-QA tasks, as shown in Figure 3, the existing general LLMs perform worse than the task-specific fine-tuned models. For example, on the entity extraction task using the NCBI disease dataset dougan2014ncbi , the state-of-the-art task-specific fine-tuned model BioBERT lee2020biobert achieves an F1 score of 89.36, substantially exceeding the F1 score of 56.73 by GPT-4. We hypothesize that the reason for the strong QA capability of the current general LLMs is that the QA task is close-ended; i.e. the correct answer is already provided in multiple candidates. In contrast, most non-QA tasks are open-ended where the model has to predict the correct answer from a large pool of possible candidates, or even without any candidates provided.
Overall, the comparison proves that the current general LLMs have a strong question answering capability, however, the capability on other tasks still needs to be improved. Therefore, we advocate that the evaluation of medical LLMs should be extended to a broad range of tasks including non-QA tasks, instead of being limited mainly to medical QA tasks. Hereafter, we will discuss specific clinical applications of LLMs, followed by their challenges and future directions.
4 Clinical Applications
As shown in Figure 5, this section discusses the clinical applications of LLMs. Each subsection contains a specific application and discusses how LLMs perform this task.
4.1 Medical Diagnosis
Medical diagnosis involves the medical practitioner using objective medical data from tests and self-described subjective symptoms to conclude the most likely health problem occurring in the patient AImed . It is important to diagnose a patient in an accurate and timely manner because the effectiveness of treatment for most diseases is time-sensitive. A missed or inaccurate diagnosis often has negative consequences, ranging from a minor inconvenience to death. For example, breast cancer has the highest mortality rate in the world because many communities lack trained personnel to perform proper checks gao2023designing . It may increase the accessibility of professional healthcare AImed ; thirunavukarasu2023large to incorporate LLMs into the medical diagnosis pipeline.
Discussion
One distinct limitation of using LLMs as the sole tool for medical diagnosis is the heavy reliance on subjective text inputs from the patient. Since LLMs are text-based, they lack the inherent capability to analyze medical diagnostic imagery. Given that objective medical diagnoses frequently depend on visual images, LLMs are often unable to directly conduct diagnostic assessments as they lack concrete visual evidence to support disease diagnosis liu2021PPKED . However, they can help with diagnosis as a logical reasoning tool for improving accuracy in other vision-based models. One such example is ChatCAD wang2023chatcad , where images are first fed into an existing computer-aided diagnosis (CAD) model to obtain tensor outputs. These outputs are translated into natural language, which is subsequently fed into ChatCAD to summarize results and formulate diagnoses. ChatCAD achieves a recall score of 0.781, significantly higher than that (0.382) of the state-of-the-art task-specific model. We note that although the more recent GPT-4V(ision) is capable of interpreting images in the general domain, its extension to the medical image domain is yet to materialize. Nevertheless, all aforementioned methods of implementing LLMs rely on image transformation into text beforehand.
Figure 5: Integrated overview of potential applications gao2023leveraging ; dong2022automated ; d2022emotion ; biri2023assessing ; vaidyam2019chatbots of large language models in medicine.
4.2 Formatting and ICD-Coding
The international classification of diseases (ICD) world1978international is a method of standardizing diagnostic and procedural information of a clinical session. These ICD codes are recorded in the patient’s EHRs every doctor visit. They are also used for tracking health metrics, treatment outcomes, and billing. There is a need to automate the ICD labeling process because its manual entry process is very time consuming for doctors.
LLMs can help automate ICD coding by extracting medical terms from clinical notes and assigning corresponding ICD codes ong2023applying ; dong2022automated . For example, PLM-ICD huang2022plm is a multi-class classification model fine-tuned for automatic ICD coding. The base model used in PLM-ICD is domain-specific with medicine-specific knowledge to enhance the ability to understand medical terms. PLM-ICD uses segment pooling, the algorithm that divides long input texts into shorter representations using LLMs when the input surpasses the maximum allowable length. Lastly, it relates the encoding to the augmented labels to output ICD codes for each clinical input. PLM-ICD produced a higher AUC score than previous state-of-the-art lightweight models huang2022plm .
Discussion
One important issue with LLMs is the potential biases and hallucinations. In particular, traditional multi-label classification models can easily constrain their outputs to a predefined list of (usually >1000) ICD candidate codes through a classification neural network. In contrast, generative LLMs struggle to effectively handle this extensive list through prompting techniques. As a result, the LLM may assign an ICD code that is not in the candidate list or a non-existent ICD code to the input text. It leads to confusion when interpreting medical records yang2022large and is therefore crucial to establish a proactive mechanism to detect and rectify errors before they enter patient EHRs.
4.3 Clinical Report Generation
Clinical reports, such as radiology reports wu2023towards , discharge summaries patel2023chatgpt , discharge instruction liu2022retrieve , and patient clinic letters ali2023using , refer to standardized documentation that healthcare workers complete after each patient visit liu2019reporting . A large portion of the report is often medical diagnostic results. It is typically tedious for overworked clinicians to write clinical reports, and thus they are often incomplete or error-prone. Adopting LLMs for clinical report generation can provide an objective means of increasing their completeness and accuracy, as well as reducing clinical workload. The generated report is intended to be a document that the clinician can review, modify, and approve as necessary patel2023chatgpt ; ali2023using ; liu2022retrieve .
Meanwhile, LLMs can be used intuitively as a summarization tool wang2023chatcad to help with clinical report generation. Given the diagnosis as input, an LLM can use its capabilities of text summarization to produce a clear and concise conclusion. In this instance, LLMs do not directly contribute to improving the accuracy of the conclusion. Instead, they act as a tool for the work that otherwise would have been undertaken by doctors.
Another popular utilization of LLMs for generating clinical reports relies on a vision-based model or manual input acting as a precursor in the pipeline wang2023chatcad ; wu2023towards ; wu2023can . This precursor step analyzes the input medical image and feeds the LLM with some type of annotation. The LLM uses the annotation alongside some other text prompt (e.g. report format, additional input by medical personnel) to generate an accurate and fluent report that follows the parameters. This has been shown to greatly reduce the workload on doctors wang2023chatcad .
Discussion
Although LLMs have generated more complete and accurate clinical reports than human counterparts van2023clinical , they are still susceptible to hallucinations as well as tend to approach inputs with a literal view instead of an assumption-based perspective that is often taken by human doctors. Additionally, human-written reports are generally more concise than reports generated by LLMs van2023clinical .
4.4 Medical Robotics
Medical robotics are revolutionizing healthcare, offering precision in tasks ranging from medication distributioncheng2023multi to diagnostic analyses and multi-agent surgical planning xia2023nested . These intelligent systems enhance the medical workforce, bridge gaps in staffing, and perform tasks that extend human physical limits.
Incorporating LLMs, robots interpret environmental data to navigate and execute complex tasks. The Graph-based Robotic Instruction Decomposer ni2023grid exemplifies this by utilizing natural language for route planning. This method has been shown to outperform GPT-4 by over 25.4% accuracy in simultaneously predicting the correct action and object, as well as 43.6% accuracy in correctly predicting instruction tasks ni2023grid .
Moreover, robotics in surgery now involve multi-agent planning systems wang2024large . Such systems involve a symphony of robotic units working collaboratively, each programmed to perform specific tasks that complement one another. For example, one robotic arm may be responsible for precise incisions while another manages real-time imaging, providing surgeons with a dynamic view of the operation. This orchestrated approach enhances surgical accuracy, reduces procedure times, and minimizes patient recovery periods. In diagnostics, robots leverage the power of data analytics to sift through extensive EHRs, identifying patterns and anomalies that may elude the human eye pee2019artificial .
Using LLMs in medical robotics can also improve human-computer interactions. Robots with improved interactivity may recognize human emotions and requests through natural language inputs. This makes patient communication with robots less intimidating and more user-friendly than existing implementationsqiu2023large ; d2022emotion .
Discussion
Some challenges with implementing medical robotics are similar to those with collaborative robots (cobots), as both cases involve robots operating alongside humans that require trust between human and computeremaminejad2023trust . Different from cobots, when implementing LLM into algorithms for route planning and robotic motion, errors in judgment due to the effects of bias and hallucination may occur. This concern can be mitigated by the better ability of traditional robots to inhibit damage than cobots. The low prediction accuracy score (< 0.5) of current LLM-powered medical robots means they are still unsuitable for implementation in real clinical settings cheng2023multi ; xia2023nested ; robinson2023brief , which also indicates research opportunities.
4.5 Medical Language Translation
There are two main areas of medical language translation; the translation of medical terminology from one language to another noll2023machine and the translation of professional medical dialogue into expressions that are easily interpreted by non-professional personnel karabacak2023advent . Both areas are important for seamless communication between different groups. The current language barrier to global collaboration in both research and medical techniques can be largely reduced with the help of LLMs. Machine translation has been proven to be 7% more accurate than traditional human services karabacak2023advent . Language translation also improves accuracy in education resources and research articles, making knowledge accessible worldwide ahn2023impending . For the second use case, LLMs can improve medical education by identifying the skill levels of students and providing corresponding terminology and structures they will understand. This application will also help patients understand professional medical speech karabacak2023advent .
Discussion
One ethical consideration of using LLMs to perform medical translation is the potential for discriminatory verbiage to be inserted inadvertently into the output. Such verbiage is difficult to prevent due to the nature of the pipeline. This may cause miscommunications and even have legal consequences. Additionally, potential misinformation caused by translation errors may confuse patients and, in the worst case, may cause them to unintentionally inflict harm to themselves karabacak2023advent .
4.6 Medical Education
LLMs can be incorporated into the medical education system in different ways, including facilitating study through explanations, aiding in language translation, answering questions, assisting with medical exam preparation, and providing Socratic-style tutoring qiu2023large ; biri2023assessing . It has been suggested that medical education can be augmented by generating scenarios, problems, and corresponding answers by an LLM. Students will gain a richer educational experience through personalized study modules and case-based assessments, encountering a wider array of challenges and scenarios beyond those found in standard textbooks karabacak2023advent . LLMs can also generate feedback on the students’ responses to practical problems, allowing students to know their areas of weakness in real-time. Inherently, these will better prepare these medical students for the real world since they would have been exposed to more scenariosahn2023impending .
Another use of LLMs in the medical field is educating the public. Medical dialogues are often complex and difficult to understand for the average patient. LLMs can tune the textual output of prompts to use varying degrees of medical terminology for different audiences. This will make medical information easy to understand for the average person while ensuring medical professionals have access to the most precise information karabacak2023advent .
Discussion
Potential downsides of using LLMs in medical education include the current lack of ethical training and biases in training datasetsabd2023large . In addition, misinformation (e.g. hallucination) presents a challenge in utilizing LLMs for medical education.
4.7 Mental Health Support
Mental health support involves both diagnosis and treatment. For example, depression is treated through a variety of psychotherapies, including cognitive behavior therapy, interpersonal psychotherapy, psychodynamic therapy, etc. vaidyam2019chatbots . Many of these techniques are primarily dominated by patient-doctor conversations, with lengthy treatment plans that are cost-prohibitive for many. The ability of LLMs to serve as conversation partners and companions may significantly lower the barrier to entry for patients with financial or physical constraints stock2023tell , increasing the accessibility to mental health treatments liu2023chatcounselor . There have been various research works on the effects of incorporating LLMs into the treatment plan liu2023chatcounselor ; de2023benefits .
The level of self-disclosure has a heavy impact on the effectiveness of mental health diagnosis and treatment. The degree of willingness to share has a direct impact on the diagnosis results and treatment plan. Studies have shown that patient willingness to discuss mental health-related topics with a robot is high robinson2023brief ; de2023benefits . Alongside the convenience and lower financial stakes, mental health support by LLMs has the potential to be more effective than human counterparts in many scenarios.
Discussion
One major difficulty in employing LLMs for mental health support is the difference between written and spoken communication. Hill et al. hill2015real found that people answered questions differently when asked to write the answer down when compared to when verbally expressing their answers. Future studies could include longer-term research to analyze how social penetration (i.e., the development of interpersonal relationships) over time affects information disclosure stock2023tell .
Figure 6: The challenges and potential solutions of deploying medical LLMs in clinical practice, aiming to inspire further research and development in this area.
5 Challenges
We address the challenges and discuss solutions to the adoption of LLMs in an array of medical applications (Figure 6).
5.1 Hallucination
Hallucination of LLMs refers to the phenomenon where the generated output contains inaccurate or nonfactual information. It can be categorized into intrinsic and extrinsic hallucinations umapathi2023med ; rawte2023survey . Intrinsic hallucination generates outputs logically contradicting factual information, such as wrong calculations of mathematical formulas rawte2023survey . Extrinsic hallucination happens when the generated output cannot be verified, typical examples include LLMs ‘faking’ citations that do not exist or ‘dodging’ the question. When integrating LLMs into the medical domain, fluent but nonfactual LLM hallucinations can lead to the dissemination of incorrect medical information, causing misdiagnoses, inappropriate treatments, and harmful patient education. It is therefore vital to ensure the accuracy of LLM outputs in the medical domain.
Potential Solutions
Current solutions to mitigate LLM hallucination can be categorized into training-time correction, generation-time correction, and retrieval-augmented correction. The first (i.e. training-time correction) adjusts model parameter weights, thus reducing the probability of generating hallucinated outputs. Its examples include factually consistent reinforcement learning factualRL and contrastive learningcontrastivelearning-factual . The second (i.e. generation-time correction) adds a ‘reasoning’ process to the LLM inference to ensure reliability, using drawing multiple samples manakul2023selfcheckgpt or a confidence score to identify hallucination before the final generation. The third approach (i.e. retrieval-augmented correction) utilizes external resources to mitigate hallucination, for example, using factual documents as prompts factualdocumenthallucination or chain-of-retrieval prompting technique chain-of-verification .
5.2 Lack of Evaluation Benchmarks and Metrics
Current benchmarks and metrics often fail to evaluate LLM’s overall capabilities, especially in the medical domain. For example, MedQA (USMLE) medqa and MedMCQA Medmcqa offer extensive coverage on QA tasks but fail to evaluate important LLM-specific metrics, including trustworthiness, helpfulness, explainability, and faithfulness xie2023faithful . It is therefore imperative to develop domain and LLM-specific benchmarks and metrics.
Potential Solutions
Singhal et al. medpalm proposed HealthSearchQA consisting of commonly searched health queries, offering a more human-aligned benchmark for evaluating LLM’s capabilities in the medical domain. Benchmarks such as TruthfulQA truthfulqa and HaluEval halueval evaluate more LLM-specific metrics, such as truthfulness, but do not cover the medical domain. Future research is necessary to meet the need for more medical and LLM-specific benchmarks and metrics than what are currently available.
5.3 Domain Data Limitations
Current datasets in the medical domain (Table 1) remain relatively small compared to datasets for training general-purpose LLMs (Table LABEL:table:llm_summary). These limited small datasets only cover a small space medpalm of the vase domain of medical knowledge. This results in LLMs exhibiting extraordinary performance on open benchmarks with extensive data coverage, yet falling short on real-life tasks such as differential diagnosis and personalized treatment planning medpalm2 .
Although the volume of medical and health data is large, most require extensive ethical, legal, and privacy procedures to be accessed. In addition, these data are often unlabeled, and solutions to leverage these data, such as human labeling and unsupervised learning liu2021KGAE , face challenges due to the lack of human expert resources and small margins of error.
Potential Solutions
Current state-of-the-art approaches medpalm2 ; li2023chatdoctor typically fine-tune the LLMs on smaller open-sourced datasets to improve their domain-specific performance. Another solution is to generate high-quality synthetic datasets using LLMs to broaden the knowledge coverage; however, it has been discovered that training on generated datasets causes models to forget shumailov2023model . Future research is needed to validate the effectiveness of using synthetic data for LLMs in the medical field.
5.4 New Knowledge Adaptation
LLMs are trained on extensive data to learn knowledge. Once trained, it is expensive and inefficient to inject new knowledge into an LLM through re-training. However, it is sometimes necessary to update the knowledge of the LLM, for example, on a new adverse effect of a medication or a novel disease. Two problems occur during such knowledge updates. The first problem is how to make LLMs appropriately ‘forget’ the old knowledge, as it is almost impossible to remove all ‘old knowledge’ from the training data, and the discrepancy between new and old knowledge can cause unintended association and bias hoelscher2023detecting . The second problem is the timeliness of the additional knowledge - how do we ensure the model is updated in real-time liu2023medical ? Both problems pose significant barriers to using LLMs in medical fields, where accurate and timely updates of medical knowledge are crucial in real-world implementations.
Potential Solutions
Current solutions to knowledge adaptation can be categorized into model editing and retrieval-augmented generation. Model editing modelediting alters the knowledge of the model by modifying its parameters. However, this method does not generalize well, with their effectiveness varying across different model architectures. In contrast, retrieval-augmented generation provides external knowledge sources as prompts during model inference; for example, Lewis et al. RAGnewknowledge enabled model knowledge updates by updating the model’s external knowledge memory.
5.5 Behavior Alignment
Behavior alignment refers to the process of ensuring that the LLM’s behaviors align with the objectives of its task. Development efforts have been spent on aligning LLMs with general human behavior, but the behavior discrepancy between general humans and medical professionals remains challenging for adopting LLMs in the medical domain. For example, ChatGPT is well aligned with general human behavior, but their answers to medical consultations are not as concise and professional as those by human experts instructgpt . In addition, misalignment in the medical domain introduces unnecessary harm and ethical concerns alignmentAI that lead to undesirable consequences.
Potential Solutions
Current solutions include instruction fine-tuning, reinforcement learning from human feedback (RLHF) instructgpt , and prompt tuning lester2021power ; liu2021p . Instruction fine-tuning zhang2023instruction refers to improving the performance of LLMs on specific tasks based on explicit instructions. For example, Ouyang et al. instructgpt used it to help LLMs generate less toxic and more suitable outputs. RLHF uses human feedback to evaluate and align the outputs of LLMs. It has been shown to be effective in multiple tasks, including becoming helpful chatbots rlhfdialogue and decision-making agents webgpt . Prompt tuning can also align LLMs to the expected output format. For example, Liu et al. chainofhindsight uses a prompting strategy, chain of hindsight, to enable the model to detect and correct its errors, thus aligning the generated output with human expectations.
5.6 Ethical, Legal and Safety Concerns
Concerns have been raised regarding using LLMs (e.g. ChatGPT) in the medical domain chatgptconcernmed , with a focus on ethics, accountability, and safety. For example, the scientific community has disapproved of using ChatGPT in writing biomedical research papers howthefuckisthisonNature due to ethical concerns. The accountability of using LLMs as assistants to practice medicine is challenging he2023survey ; tian2024opportunities . Li et al.chatgptjailbreak and Shen et al.shen2023anything found that prompt injection can cause the LLM to leak personally identifiable information (PII), e.g. email addresses, from its training data, which is a significant vulnerability when implementing LLM in the medical domain.
Potential Solutions
With no immediate solutions available, we have nevertheless observed research efforts to understand the cause of these ethical and legal concerns. For example, Wei et al. understandsafety propose that PII leakage is attributed to the mismatched generalization between safety and capability objectives (i.e., the pre-training of LLMs utilizes a larger and more varied dataset compared to the dataset used for safety training, resulting in many of the model’s capabilities are not covered by safety training).
Figure 7: Future directions of LLMs in clinical medicine in terms of both development and deployment.
6 Future Directions
Although LLMs have already made a significant impact on people’s lives through chatbots and search engines, their integration into medicine is still in the infant stage. As shown in Figure 7, numerous new avenues of medical LLMs await researchers and practitioners to explore how to better serve the general public and patients.
6.1 Introduction of New Benchmarks
Recent studies have underscored the shortcomings of existing benchmarks in evaluating LLMs for clinical applications chen2023extensive ; chen2023large . Traditional benchmarks, which primarily gauge accuracy in medical question-answering, inadequately capture the full spectrum of clinical skills necessary for LLMs medpalm . Criticisms have been leveled against the use of human-centric standardized medical exams for LLM evaluation, arguing that passing these tests does not necessarily reflect an LLM’s proficiency in the nuanced expertise required in real-world clinical settings medpalm . In response, there is an emerging consensus on the need for more comprehensive benchmarks. These should include capabilities like sourcing from authoritative medical references, adapting to the evolving landscape of medical knowledge, and clearly communicating uncertainties thirunavukarasu2023large ; medpalm . Additionally, considering the sensitive nature of healthcare, these benchmarks should also assess factors such as fairness, ethics, and equity, which, though crucial, pose quantification challenges medpalm . The aim is to create benchmarks that more effectively mirror actual clinical scenarios, thus providing a more accurate measure of LLMs’ suitability for medical advisory roles.
6.2 Multimodal LLM Integrated with Time-Series, Visual, and Audio Data
Multimodal LLMs (MLLMs), or Large Multimodal Models (LMMs), are LLM-based models designed to perform multimodal (e.g. involving both visual and textual) tasks yin2023survey . While LLMs primarily address NLP tasks, MLLMs support a broader range of tasks, such as comprehending the underlying meaning of a meme and generating website codes from images. This versatility suggests promising applications of MLLMs in medicine. Several MLLM-based frameworks integrating vision and language, e.g. MedPaLM M tu2023towards , LLaVA-Med li2023llavamed , Visual Med-Alpaca Visual-Med-Alpaca , Med-Flamingo moor2023medflamingo , and Qilin-Med-VL liu2023qilin , have been proposed to adopt the medical image-text pairs for fine-tuning, thus enabling the medical LLMs to efficiently understand the input medical (e.g. radiology) images. A recent study huang2023chatgpt proposes to integrate vision, audio, and language inputs for automated diagnosis in dentistry. However, there exist only very few medical LLMs that can process time series data, such as electrocardiograms (ECGs) li2023frozen and sphygmomanometers (PPGs) englhardt2023exploring , despite such data being important for medical diagnosis and monitoring.
6.3 Medical Agents
LLM-based agents xi2023rise ; wang2023survey (i.e., intelligent systems or software designed to perform specific medical tasks or functions) have achieved significant progress in solving complex tasks (e.g. software design, molecular dynamics simulation) through human-like behaviors, such as role-playing and communication li2023camel ; tang2023medagents . However, integrating these agents effectively within the medical domain remains a challenge. The medical field involves numerous roles tang2023medagents and decision-making processes, especially in disease diagnosis that often requires a series of investigations involving CT scans, ultrasounds, electrocardiograms, and blood tests. The idea of utilizing LLMs to model each of these roles, thereby creating collaborative medical agents, presents a promising direction. These agents could mimic the roles of radiologists, cardiologists, pathologists, etc., each specializing in interpreting specific types of medical data. For example, a radiologist agent could analyze CT scans, while a pathologist agent could focus on blood test results. The collaboration among these specialized agents could lead to a more holistic and accurate diagnosis. By leveraging the comprehensive knowledge base and contextual understanding capabilities of LLMs, these agents not only interpret individual medical reports, but also integrate these interpretations to form a cohesive medical opinion.
6.4 LLMs in Medical Sub-domains
Current LLM research in medicine has largely focused on general medicine, likely due to the greater availability of data in this area medpalm2 ; tian2024opportunities . This has resulted in the under-representation of LLM applications in specialized fields like ‘rehabilitation therapy’ and ‘sports medicine’. The latter, in particular, holds significant potential, given the global health challenges posed by physical inactivity. The World Health Organization identifies physical inactivity as a major risk factor for non-communicable diseases (NCDs), impacting over a quarter of the global adult population who2022physical . Despite initiatives to incorporate physical activity (PA) into healthcare systems, implementation remains challenging, particularly in developing countries with limited PA education among healthcare providers who2022physical . LLMs could play a pivotal role in these settings by disseminating accurate PA knowledge and aiding in the creation of personalized PA programs connor2023large . Such applications could significantly enhance PA levels, improving global health outcomes, especially in resource-constrained environments.
6.5 Interdisciplinary Collaborations
Just as interdisciplinary collaborations are crucial in safety-critical areas like nuclear energy production, collaborations between the medical and technology communities for developing medical LLMs are essential to ensure AI safety and efficacy in medicine. The medical community has primarily LLMs provided by technology companies without questioning their data training, which is a sub-optimal situation. Medical professionals are therefore encouraged to actively participate in creating and deploying medical LLMs by providing relevant training data, defining the desired benefits of LLMs, and conducting tests in real-world scenarios to evaluate these benefits thirunavukarasu2023large ; omiye2024large ; clusmann2023future . Such assessments would help to determine the legal and medical risks associated with LLM use in medicine and inform strategies to mitigate LLM hallucination mello2023chatgpt .
7 Conclusion
This review provides a comprehensive overview of the principles, applications, and challenges of LLMs in medicine, intended to promote further research and exploration in this interdisciplinary field. With the rapid development of foundation models, the LLMs could significantly improve future clinical practice and medical discoveries for the benefit of society. It requires sustained interdisciplinary collaboration between clinicians and AI researchers, doctor-in-the-loop, and human-centered design, to address existing challenges with applying LLMs in the medical domain. The co-development of the collaboration for appropriate training data, benchmarks, metrics, and deployment strategies could enable faster and more responsible implementation of medical large language models than the current practice.
8 Acknowledgements
This work was supported in part by the Pandemic Sciences Institute at the University of Oxford; the National Institute for Health Research (NIHR) Oxford Biomedical Research Centre (BRC); an NIHR Research Professorship; a Royal Academy of Engineering Research Chair; the Well-come Trust funded VITAL project; the UK Research and Innovation (UKRI); the Engineering and Physical Sciences Research Council (EPSRC); and the InnoHK Hong Kong Centre for Cerebro-cardiovascular Engineering (COCHE).
9 Author Contributions
FL, ZL, JL, and DC conceived the project. FL conceived and designed the study. HZ, FL, BG, XZ, and JH conducted the literature review, performed data analysis, and drafted the manuscript. All authors contributed to the interpretation and final manuscript preparation. All authors read and approved the final manuscript.
Appendix A Appendix: Discriminative Tasks
A.1 Question Answering
Task Description
Question Answering (QA) [226] aims to give answers to the given queries, by generating multiple-choice or free-text responses. A multiple-choice response happens when the model is asked a question with the material for answering the question included. For example, the QA for ’Is hypertension a risk factor for cardiovascular disease, yes or no?’ is multiple-choice-orientated. In contrast, an open question without potential answers provided gives rise to a free-text response, an example QA being ’What are the common symptoms of influenza?’.
Table 3: The performance (accuracy) on five question answering (QA) datasets, shown in percentage (%). FT: fine-tuned.
Types Models MedQA (USMLE) PubMedQA MedMCQA MMLU (Clinical Knowledge) MMLU (Professional Medicine)
Task-specific FT Models BERT [70] 44.6 51.6 43.0 - -
RoBERTa [80, 81] 43.3 52.8 - - -
BioBERT [29, 90] 30.1 60.2 - - -
SciBERT [91] 29.5 57.4 - - -
ClinicalBERT [31] 29.1 49.1 - - -
BlueBERT [121, 227, 228] - 48.4 - - -
PubMedBERT [30] 38.1 55.8 - - -
General LLMs Med-PaLM-2 [11] 86.5 81.8 72.3 88.3 95.2
FLAN-PaLM (few-shot) [10] 67.6 79.0 57.6 - -
GPT-4 (zero-shot) [7] 78.9 75.2 69.5 86.0 93.0
GPT-4 (few-shot) [7] 86.1 80.4 73.7 86.4 93.8
GPT-3.5 (zero-shot) [6] 50.8 71.6 50.1 69.8 70.2
GPT-3.5 (few-shot) [6] 60.2 78.2 62.7 68.7 69.8
Clinical Camel-13B (zero-shot) [18] 34.4 72.9 39.1 54.0 51.8
Clinical Camel-13B (5-shot) [18] 45.2 74.8 44.8 60.4 53.3
Clinical Camel-70B (zero-shot) [18] 53.4 74.3 47.0 69.8 71.3
Clinical Camel-70B (5-shot) [18] 60.7 77.9 54.2 72.8 75.0
Galactica [85] 44.4 77.6 52.9 - -
BioMedLM [95] 50.3 74.4 - - -
BioGPT [94] - 81.0 - - -
PMC-LLaMA [13] 56.4 77.9 56.0 - -
Human (expert) [123, 229] 87.0 78.0 90.0 - -
Datasets and Models
The commonly used biomedical QA datasets are MedQA (USMLE) [14], PubMedQA [133], and MedMCQA [131]. Such datasets are relatively small in size compared to general datasets, thus the model performance can be readily improved by pre-training on general data and then finetuning on the biomedical data [230, 231]. Pergola et al. [232] proposed a biomedical-specific masking method: instead of masking tokens randomly, the model identifies biomedical-related tokens and masks them to focus on in-domain learning. Wang et al. [233] defined the self-questioning prompting (SQP) and utilize it on the BioASQ dataset. To extract useful information for specific tasks, SQP lets the GPT model ask questions about the given text and then answer these questions. However, hallucination remains a threat to the quality of outputs in QA. One solution is to use biomedical search systems such as Almanac [234]. Another approach is to use extra datasets as an augmentation for the QA task. There are many chatbots based on LLM QA, such as Clinical Camel [18], DoctorGLM [32], ChatDoctor [15], HuaTuo [17], HuaTuoGPT [33], and MedAlpaca [16]. Except for a few [235], most chatbots are a black box the the consumers, and thus providing insight into them is a potential direction for further studies.
Evaluation
Existing works use accuracy as the metric for model performance in QA. We compare the performance of task-specific (fine-tuned) BERT variations [30, 236, 14, 237], Med-PaLM-2 [11, 18], FLAN-PaLM [10], GPT-4 [238, 18], GPT-3.5 [116, 18], Clinical Camel [18], Galactica [85], BioMedLM [95], BioGPT [94], and PMC-LLaMA [13] on the datasets described in Table 2. The citations in this paragraph correspond to the sources that provide data for assessing the model performance. Table 3 details their performances on five widely-used datasets, indicating that GPT-series models perform much better than BERT-series models. For GPT-series models, the observation that the few-shot setting outperforms the zero-shot setting still holds. Med-PaLM-2 has the highest accuracy among all models.  These results indicate that it is feasible to fine-tune general LLMs on medical data and improve their performance.
Table 4: The performance (F1-score) on four entity extraction datasets, , shown in percentage (%).
Types Models NCBI Disease BC5CDR Disease BC5CDR Drug/Chem. JNLPBA
Task-specific FT Models BERT [70] 85.63 82.41 91.16 74.94
BioBERT [29, 90] 89.36 86.56 93.44 77.59
SciBERT [91] 88.57 90.01 90.01 77.28
PubMedBERT [30] 87.82 85.62 93.33 79.10
BlueBERT [121, 227, 228] 88.04 83.69 91.19 77.71
General LLMs ClinicalBERT [31] 86.32 83.04 90.80 78.07
GPT-3 [6] 51.40 43.60 73.00 -
GPT-3.5 (zero-shot) [6] 24.05 - 29.25 -
GPT-3.5 (one-shot) [6] 12.73 - 18.03 -
GPT-4 (zero-shot) [7] 56.73 - 74.43 -
GPT-4 (one-shot) [7] 48.37 - 82.07 -
ChatGPT [239] 50.49 51.77 60.30 41.25
BARD (zero-shot) [87] 96.00 - 97.70 -
BARD (5-shot) [87] 95.60 - 98.30 -
A.2 Entity Extraction
Task Definition
Entity extraction, or named entity recognition (NER) [240] aims to identify named entities mentioned in unstructured text, and classify them into predefined categories such as the names of persons, organizations, locations, medical codes, time expressions, quantities, monetary values, percentages, etc. Existing task-specific fine-tuned models leverage large-scale pre-trained language representations, and then fine-tune them on NER datasets to achieve state-of-the-art performances. The self-attention mechanisms [69] are used for efficiently capturing complex patterns and dependencies.
In the healthcare domain, medical NER aims to extract medical entities such as disease names, medication, dosage, and procedures from clinical narratives, electronic health records (EHRs), and chemicals and proteins from scientific literature. This enables NER to provide a solid basis for clinical decision support systems, automated patient monitoring, etc.
Datasets and Models
Table 2 shows commonly used biomedical NER datasets, such as NCBI Disease [143], JNLPBA [144], and BC5CDR [146]. To perform the NER task, for each input token, the existing models output a dense representation, which not only embeds the tokens but also includes its relation with other tokens in the text. Therefore, with the dense representations of input tokens extracted as vectors, additional layers are applied to the last Transformer layer to fit the downstream entity extraction task. The widely-used additional layers are softmax, BiLSTM [241], CRF [242, 243], and their combinations [244, 242, 245, 243, 246, 247]. Gu et al. [248] proposed a distillation method that uses few-shot GPT-3.5 to extract the correct entities and create the training set for the student model. This method shows that GPT models can be used to label the dataset first to achieve unsupervised learning. Wang et al. [233] defined a new prompting method self-questioning prompting (SQP) for NER. Wang et al. [233] further evaluated the performances of BARD, GPT-3.5, and GPT-4 with SQP and achieved some state-of-the-art results.
Evaluation
The entity-level F1 score is widely used to evaluate the models’ performance. Recent studies on the model performance of LLMs on biomedical NER tasks include GPT-3 [249, 250, 251], GPT-3.5 [62], GPT-4 [62], and ChatGPT [61]. The citations in this paragraph correspond to the sources that provide data for assessing the model performance.
We summarize the performances of existing LLMs on the NER task in Table 4. It clearly shows that both the encoder-only BERT-series models (e.g. BioBERT [29], SciBERT [91], PubMedBERT [30], BlueBERT [121], and ClinicalBERT [31]) significantly outperform most decoder-only LLM GPT-series models (e.g., GPT-3 [6] and GPT-4 [7]), under both the zero-shot and few-shot settings. Recent work has investigated the performance of GPT-3 [249, 250, 251], GPT-3.5 [62], and GPT-4 [62] on biomedical NER tasks. These GPT models still need further development to outperform traditional task-specific fine-tuned models [251].
A.3 Relation Extraction
Task Description
Different from entity extraction, relation extraction (RE) aims to find the relation between entities in a text. It is highly related to entity extraction since identifying entities first can improve the model’s performance. A relation can be ordered or unordered. For example, in the sentence ’I work in London’ the relation ’workspace’ is order-sensitive and the pair (I, London) is of such a relation, but in the sentence ’Alice and Bob are friends’ the relation ’friend’ is not order-sensitive. In the biomedical area, relation extraction is often used to extract relations in a MedAbstrct text for further applications like text summarization.
Datasets and Models
For relation extraction, Table 2 shows representative biomedical datasets such as BC5CDR, BioRED, and DDI. In practice, most datasets indicate the potential entity for the relation instead of asking the model to find it, so RE is usually considered as a classification task. One common technique for task-specific fine-tuned models is to use the classification token (e.g., [CLS] token in Transformer [69]), which learns the information of all input tokens, and is further used to classify the relation. Some studies use only the classification token and apply softmax in the last layer [252, 253, 254, 255]. Others use not only the classification token, but also all other token representations. It has been shown that using all token representations outperforms using classification token only [256]. Specifically, Wang et al. [233] defined the self-questioning prompting (SQP) as discussed in Sec A.1 and utilize it in the DDI RE task.
Evaluation
Existing studies on RE model performance typically use the F1-score, since the model performs a straightforward classification task. Table 5 compares the performance of BioBERT [29, 257], SciBERT [91, 257], PubMedBERT [30, 257], BioGPT [94], GPT-3 [249, 250, 251], GPT-3.5 [62], GPT-4 [62], and ChatGPT [61] on the datasets introduced above. It clearly shows that at present task-specific fine-tuned models outperform GPT models. For GPT models, the F1-score of the best GPT model (GPT-4) is below 70%, while that of all the task-specific fine-tuned models based on BERT is approximately 80%. There is no appreciable difference in F1-score between zero-shot and few-shot [233].
Table 5: The performance (F1-score) on four relation extraction datasets, shown in percentage (%).
Types Models BC5CDR ChemProt DDI GAD
Task-specific FT Models BioBERT [29, 90] - 76.46 80.88 82.36
SciBERT [91] 79.00 83.64 84.08 81.34
PubMedBERT [30] - - 82.36 83.96
BioGPT [94] 46.17 - 40.76 -
General LLMs GPT-3 [6] - 25.90 16.10 66.00
GPT-3.5 (zero-shot) [6] - 57.43 33.49 -
GPT-3.5 (one-shot) [6] - 61.91 34.40 -
GPT-4 (zero-shot) [7] - 66.18 63.25 -
GPT-4 (one-shot) [7] - 65.43 65.58 -
ChatGPT [239] - 34.16 51.62 52.43
BARD (zero-shot) [87] - - 56.60 -
BARD (5-shot) [87] - - 77.20 -
Table 6: The performance (F1-score) on four text classification datasets, shown in percentage (%).
Types Models HoC i2b2 MIMIC-III OHSUMED
Task-specific FT Models PubMedBERT [30] 82.32 - - -
BioGPT [94] 85.12 - 82.30 -
General LLMs GPT-4 (5-shot) [7] - 99.00 67.40 -
ChatGPT (0-shot) [239] - 92.90 - 39.93
ChatGPT (2-shot) [239] - 92.90 - 47.05
ChatGPT (5-shot) [239] - 92.90 60.00 45.39
ChatGraph [258] - - - 60.79
Table 7: The performance (F1-score) of natural language inference, shown in percentage (%).
Types Models MedNLI
Task-specific FT Models BERT [259] 76.11
ALBERT [259] 77.84
Roberta [259] 80.14
BioBERT [259] 81.83
ClinicalBERT [259] 80.66
BlueBERT [259] 83.92
SciBERT [259] 79.43
BlueBERT+AG [259] 84.34
General LLMs GPT-3.5 [6] 82.21
GPT-3.5-Distillation [248] 80.24
GPT-4 [7] 85.69
BARD (zero-shot) [87] 76.00
BARD (5-shot) [87] 76.00
A.4 Text Classification
Task Description
Unlike entity extraction or relation extraction at the token level, text classification is a sentence or document-level task. It is a classic classification problem of assigning predefined labels to a text, where it is common for a text to have multiple labels. The task is to predict all correct labels given the input medical text, and can be used as the preprocessing of medical text simplification and summarization.
Datasets and Models
Table 2 shows commonly used datasets (e.g. HoC and OHSUMED) for text classification. Since this is a text-level task, the [CLS] vector is augmented when using task-specific fine-tuned (Transformer-based) models. Another approach for distilling the overall information is to use the weighted sum of final attention layer outputs. Recent work has shown that adding a custom attention layer after the original model (BERT) improves the model performance [260, 261]. McCreery et al. [262] double-fine-tuned the BERT model, in the sense that they first fine-tuned the model on a general dataset and then fine-tuned it on a medical-specific dataset. Another approach is to combine graph-based models with general LLMs, as in ChatGraph [258] that used ChatGPT to extract text information and apply it to a graph-based model that outperforms GPT models. CohortGPT [263] used Chain-of-Thought prompting and knowledge graph to outperform few-shot ChatGPT and GPT-4.
Evaluation
Same as above, the F1-score is used for evaluating the model performance. Table 6 compares the performance of PubMedBERT [30], BioGPT [94], GPT-4 [35, 263], ChatGPT [258, 263], and ChatGraph [258] on four benchmark datasets. It shows that GPT models perform better on the i2b2 dataset than on the OHSUMED dataset, possibly because its few-shot setting is not yet sufficiently good. Overall, few-shot GPT models outperform zero-shot ones, and the F1 scores tend to increase with more parameters.
A.5 Natural Language Inference
Task Description
Natural Language Inference (NLI) is a sentence-level task, involving two sentences: hypothesis ($H$) and premise ($P$). The task to to determine whether $H$ can be inferred from $P$. The outcome is one of the following three labels: (i) Entailment: the hypothesis can be inferred from the premise, or logically, $P\implies H$. (ii) Contradiction: the negation of the hypothesis can be inferred from the premise, or logically, $P\implies\neg H$. (iii) Neutral: all other cases, or logically, $\neg((P\implies H)\lor(P\implies\neg H))$.
Datasets and Models
The commonly used datasets, such as MedNLI [165] and BioNLI [166], are shown in Table 2 . Due to the scarcity of biomedical datasets, researchers also use general datasets (e.g. SNLI [264] and MultiNLI [265]). In task-specific fine-tuned models, similar to RE, [CLS] vector is often added to the text to show the overall information. The difference is that here we have two (instead of one) sentences, and so a [SEP] vector is also applied to indicate the separation of premise and hypothesis, yielding an overall input that looks like ‘[CLS], premise, [SEP], hypothesis’. For BERT-based models, a classification head is applied to the final [CLS] vector to predict the label. Kanakarajan et al. [90] first pre-trained BioBERT [29] on MIMIC-III [39] and then fine-tuned the model on MedNLI [165]. Cengiz et al. [266] used the so-called two-stage sequential transfer learning method. They trained the BioBERT model on SNLI [264] and MultiNLI [265] first then fine-tuned it on the MedNLI dataset [165]. The majority vote method is used to combine the prediction output of different trained models. Gu et al. [248] combined GPT-3.5 and PubMedBERT using knowledge distillation. They first fed GPT-3.5 original texts and asked it to distill the data, and the distilled data was subsequently used to further train the PubMedBERT model. Wang et al. [233] defined the self-questioning prompting (as discussed in Sec A.1) and utilize it in the MedNLI task.
Evaluation
The F1-score is used in the same manner as above. Table 7 compares the performance of task-specific fine-tuned models (e.g. BioBERT, ClinicalBERT, BlueBERT, and SciBER) on the benchmark MedNLI dataset [248], together with the performances of GPT-3.5 and GPT-3.5-Distillation. It shows that BioBERT [29] has the best performance, and the current general LLMs still need further improvement on natural language inference.
A.6 Semantic Textual Similarity
Task Description
Semantic Textual Similarity (STS) is similar to natural language inference (NLI). Their difference is that STS is a quantitative task, while NLI is a qualitative task. STS aims to produce a numerical value in a range (such as $[0,M]$) for any pair of sentences that indicates the degree of similarity, with the interpretation that 00 means two sentences are completely independent and $M$ means two sentences are completely correlated or equivalent.
Datasets and Models
As shown in Table 2, the ‘2019 n2c2/OHNLP’ [168], BIOSSES [169], and MedSTS [167] are widely-used benchmark datasets for STS. Since STS is a text-level task, [CLS] vector is naturally added to extract the overall information. Mutinda et al. [227] added a fully connected layer after the [CLS] vector as the architecture. Yang et al. [80] combined the representation of different models and added a fully connected layer. Wang et al. [267] dropped the concept of [CLS] and used the Hierarchical Convolution (HConv) layer as the added last layer. Wang et al. [233] defined the self-questioning prompting (SQP) and utilize it on the BIOSSES dataset. Xiong et al. [268] used the concatenation of character level, sentence level, and entity level representation (CSE-concate) to extract the information that is fed to the further added MLP layer.
For further training on a pre-trained model, there is a common understanding that the new dataset should have a similar distribution to that of the dataset for pre-training to improve the model performance. The iterative training is introduced for this purpose [269, 228], where the phrase "iterative" refers to the following two iterative steps: a) It first freezes the model and computes the outputs from the dataset, then it chooses a subset of the dataset so the outputs of the subset have a similar distribution to that of the pre-training dataset. b) The model is further trained using the obtained subset obtained in Step a).
Evaluation
The Pearson correlation coefficient $r_{xy}$ is used for evaluation purposes. For $n$ pairs of data $\{x_{i},y_{i}\}_{i=1}^{n}$, we define $r_{xy}=\frac{\sum_{i=1}^{n}(x_{i}-\bar{x})(y_{i}-\bar{y})}{\sqrt{\sum_{i=1}^{n}(x_{i}-\bar{x})^{2}}\sqrt{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}}\,.$ Here, $\{x_{i}\}_{i=1}^{n}$ represents the actual similarity and $\{y_{i}\}_{i=1}^{n}$ represents the predicted similarity. Table 8 compares the performance of BERT [227, 257], ClinicalBERT [227, 257], XLNet [80], RoBERTa [80, 257], HConv-BERT [267], BARD [233], GPT-3.5 [233], GPT-4 [233], BERT (CSE-concate) [268], and ClinicalBERT (iterative training) [228] on BIOSSES and ‘2019 n2c2/OHNLP’ datasets.
It shows that the task-specific fine-tuned model RoBERTa achieves the best results on the ‘2019 n2c2/OHNLP’ dataset. Although GPT-4 has the highest Pearson correlation coefficient of 91.60%, we note that, with significantly fewer model parameters, the task-specific fine-tuned model ClinicalBERT achieves a competitive result compared to GPT-4. The overall results indicate that further work on improving the current general LLMs is needed.
Table 8: The performance (sample Pearson correlation coefficient, shown in percentage (%), of semantic textual similarity.
Types Models BIOSSES 2019 n2c2/OHNLP
Task-specific FT Models BERT [70] 81.40 69.23
ClinicalBERT [31] 91.23 83.20
XLNet [80] - 84.70
RoBERTa [80, 81] 81.25 87.78
HConv-BERT [267] - 79.40
BERT (CSE-concate) [268] - 86.80
ClinicalBERT (iterative training) [31] - 87.00
General LLMs BARD (zero-shot) [87] 57.60 -
BARD (5-shot) [87] 60.10 -
GPT-3.5 (zero-shot) [6] 87.30 -
GPT-3.5 (5-shot) [6] 89.20 -
GPT-4 (zero-shot) [7] 88.90 -
GPT-4 (5-shot) [7] 91.60 -
Table 9: The performance ((NDCG@$10$) on three information retrieval datasets, shown in percentage (%).
Types Models TREC-COVID NFCorpus BioASQ
Task-specific FT Models OpenAI cpt-text-S [270, 93] 67.9 33.2 -
OpenAI cpt-text-M [270, 93] 58.5 36.7 -
OpenAI cpt-text-L [270, 93] 56.2 38.0 -
OpenAI cpt-text-XL [270, 93] 64.9 40.7 -
BioCPT [29, 90] 70.9 35.5 55.3
General LLMs GTR-Base [271] 53.9 30.8 27.1
GTR-Large [271] 55.7 32.9 32.0
GTR-XL [271] 58.4 34.3 31.7
GTR-XXL [271] 50.1 34.2 32.4
ChatGPT [239] 76.7 35.6 -
ChatGPT [239] + GPT-4 [7] 85.5 38.5 -
A.7 Information Retrieval
Task Description
Information retrieval (IR) plays an important role in the clinical area. It is the process of retrieving relevant knowledge or information related to the query from a number of unstructured data. It satisfies one’s need for searching information, such as article recommendations and literature searches. IR in the biomedical domain contains multiple tasks, among which text summarization and text simplification are discussed in Sec B.1 and Sec B.2, respectively. Question answering is also an important task included in IR, previously discussed in Sec A.1. In this subsection, we will concentrate on the query-article relevance task, also known as the ranking task. Specifically, for a query $q$ and a dataset of articles $\{d_{i}\}_{i=1}^{n}$, we aim to find the most relevant $k$ articles $(d_{1}^{q},d_{2}^{q},\cdots,d_{k}^{q})$ where the degree of relevance is defined task-specifically.
Datasets and Models
Table 2 shows commonly used datasets, such as TREC-COVID [170], NFCorpus [171], and BioASQ [172] in the biomedical area. In addition, a general dataset, BEIR (Benchmarking IR) [172], is often included in the training step of biomedical IR. Jin et al.
[93] developed a BERT-based model BioCPT that encodes the query and articles for the ranking task. They split the method into training, inference, and evaluation steps. In the training step, they introduce query-to-document loss and document-to-query loss to train the encoders for the query and articles. In the inference step, they concatenate the encoding of the query and its best-fit article $d_{1}^{q}$ together with $k-1$ non-relevant articles $(d_{2}^{q},d_{3}^{q},\cdots,d_{k}^{q})$ found by maximum inner product search (MIPS) to further train the model to rank $d_{1}^{q}$ at the top. In the evaluation step, for each input query $q$, the model evaluates over the whole dataset $\{d_{i}\}_{i=1}^{n}$ to find the best $k$ relevant articles by MIPS and used the model from the inference step to rank those $k$ articles.
Sun et al. [272] applied zero-shot ChatGPT to rank the most relevant documents without abnormal prompting. They also further used GPT-4 to re-rank the top 30 documents retrieved by ChatGPT. Abonizio et al. [273] introduced two LLM-based data augmentation methods, namely InPars and Promptagator, for IR. For InPars, they used GPT-3 and GPT-J to generate a new query for a randomly selected document. They used few-shot prompting which provides the model with good query examples or bad query examples. For Promptagator, the major difference is that a more dataset-specific prompting is applied. Ateia and Kruschwitz [274] proposed a query expansion technique that expands the current query into a more comprehensive query, which consistently improves the performance of any successive tasks. It is done purely by the GPT model with regular instructional prompting. Similarly, Wang et al. [275] used ChatGPT to generate more refined Boolean queries for systematic reviews. They showed that ChatGPT is able to generate or refine queries with higher precision.
Evaluation
[…]
We compare the performance of Google’s Generalizable T5-based dense Retrievers (GTR) [276, 93], OpenAI’s cpt-text [270, 93], BioCPT [93], ChatGPT [272], and ChatGPT+GPT-4 [272] on the datasets we introduced before.
Table 9 shows that ChatGPT and GPT-4 have the best performance. We observe that increasing parameters does not necessarily improve the model performance; for example, GTR-XXL has a lower NDCG@10 score than those of other GTR models, but performs better. We also note that the performance of all models appears unstable across different data; for example, ChatGPT+GPT-4 has only a 38.5% score for NFCorpus, but has a score of 85.5% for TREC-COVID.
Appendix B Appendix: Generative Tasks
B.1 Text Summarization
Task Description
There are two types of text summarization: extractive and abstractive summarizations. Extractive summarization aims to find the most important sentences in the text, while omitting the redundant or irrelevant sentences. In contrast, abstractive summarization generates brand-new text that summarizes the given text. Therefore, the extractive approach is more related to token-level tasks, while the abstractive approach is more relevant to high-level tasks (e.g. text generation).
Datasets and Models
Table 2 shows the commonly used biomedical datasets PubMed [38] and MentSum [174]. Some general datasets, such as GigaWord [277], CL-SciSumm [278], and S2ORC [110], are also commonly used for text summarization.
For extractive summarization, the key task is to define some scoring system that scores all sentences and hence finds the most important ones. Moradi et al. [279] used a clustering-based method to summarize medical text. They vectorize the tokens of the text by BERT and cluster the sentence vector into $k$ clusters, and then they define an informativeness score that chooses one sentence from each cluster to form a summary. Moradi et al. [280] used the graph-based model for summarization, where sentences are treated as nodes and relations as edges. The relations are measured by the cosine similarity of vectors representing the sentences. They then used different graph ranking algorithms to choose important sentences as a summary.
Du et al. [281] used purely the Transformer-based model as the scoring system. They i) tokenized the whole text with [CLS] and [SEP] augmented, ii) further augmented the corresponding sentence and token positions to each token, and iii) fed the whole vector into the model. A sigmoid layer is added to the model so the output is between zero and one, and the output is considered the score for each sentence. Since the Transformer automatically extracts relations of one sentence to others, there is no need for manually designing a score. Chen et al. [282] also relied only on the Transformer to score the sentences. Different from Du et al. [281], they proposed the AlphaBERT that takes the character-level tokens as input,and the training is split into pre-training and fine-tuning.
McInerney et al. [283] combined the summarization model with a query to generate specific summaries. Pang et al. [284] proposed a principled inference framework with top-down and bottom-up inference techniques to improve summarization models.
There are also studies on summarizing more than one piece of text at a time. They mainly relied on graph-based or Transformer-based models to extract relations between different texts [285]. Zhang et al. [286] applied one-shot GPT-3.5, using dialogues and summaries from the same category as prompts to generate abstractive summarization. More studies are needed to qualitatively analyze the performance of GPT models on biomedical text summarization [287].
Evaluation
The standard Recall-Oriented Understudy for Gisting Evaluation (ROUGE) data[288], including ROUGE-1 and ROUGE-2, is used for evaluation. Table 10 compares the performance of BioBERT variations using different training sets [280], rankings (PageRank and PPF) [280], and added last layers [281] on PubMed. It shows that models fine-tuned on Pubmed or PMC data achieve improved performance. Differences in the ranking strategy do not correspond to appreciable differences in the model performance. Overall, current BERT models are far below the expectation of true text summarization, measured by the ROUGE-2 scores.
Table 10: The performance (ROUGE-1 & ROUGE-2) of the text summarization task, shown in percentage (%).
Types Models PubMed
Task-specific FT Models BioBERT (LSTM) [281] 35.82 & 17.15
BioBERTSum [281] 37.45 & 17.59
BioBERT (Pubmed + PPF) [280] 74.21 & 32.88
BioBERT (PMC + PageRank) [280] 75.82 & 34.01
BioBERT (Pubmed + PageRank) [280] 76.09 & 34.38
BioBERT (Pubmed + PMC + PageRank) [280] 76.34 & 34.67
Table 11: The performance (BLEU score) of the text simplification task on the MultiCochrane (English) dataset, shown in percentage (%). Higher is better for BLEU.
Table 12: The performance (accuracy) of the text simplification on the AutoMeTS, shown in percentage (%). Higher is better for BLEU.
Model MultiCochrane
GPT-3 (zero-shot) [6] 2.38
Flan-T5 (zero-shot) [76] 8.12
mT5 [76] 8.82
Flan-T5 (fine-tuned) [76] 8.70
Model AutoMeTS
BERT [70] 62.40
RoBERTa [80, 81] 53.28
XLNet [80] 46.20
GPT-2 [82] 49.00
BERT [70] + RoBERTa [80, 81] + XLNet [80] + GPT-2 [82] 64.52
Table 12: The performance (accuracy) of the text simplification on the AutoMeTS, shown in percentage (%). Higher is better for BLEU.
B.2 Text Simplification
Task Description
Text simplification is to generate a new text that recovers almost all the information of the original text while improving its readability. The task has potential in biomedical education, since one major characteristic of medical education is its opaque vocabulary [182, 289].
Datasets and Models
Table 2 shows the commonly used text simplification datasets. Patel et al. [290] used NER to identify the medical terms. Lexical substitution is then applied to reduce the complexity of the text. Jeblick et al. [291] tested the performance of ChatGPT on simplifying self-collected radiology reports. They tried different prompting texts and found that the prompt ‘Explain this medical report to a child using simple language’ performs the best. However, by evaluating from a) factual correctness, b) completeness, and c) harmfulness, they concluded that GPT models may generate harmful texts which is unacceptable in the medical domain. Joseph et al. [76] evaluated the performance of zero-shot GPT-3 and Flan-T5 on MultiCochrane. They also fine-tuned mT5 [292] and Flan-T5 on this dataset. Yang et al. [293] introduced a data augmentation method for text simplification based on LLMs. For text without its simplified counterpart, they used GPT-3 to generate multiple choices for simplified text, and further trained a BERT model as the score to choose the best one as the simplification. Van et al. [81] transferred the simplification task into a prediction task. They assume they have the original text $(d_{1},d_{2},\cdots,d_{n})$ and a unfinished simplified text $(s_{1},s_{2},\cdots,s_{i})$. The task is to predict the next token $s_{i+1}$. In particular, at each time, the model receive $(d_{1},d_{2},\cdots,d_{n},s_{1},s_{2},\cdots,s_{i})$ as input and predict $s_{i+1}$. They used different models like BERT, RoBERTa, XLNet, and GPT-2. They also tried to combine the predicted token $s_{i+1}$ of all four models to outperform any of them.
Evaluation
For evaluation, we use the bilingual evaluation understudy (BLEU) score [294]. We compare the performance of GPT-3 [76], Flan-T5 [76], and mT5 [76] on MultiCochrane (English) [76] (please see Table B.1). For AutoMeTS dataset [81], we use the accuracy of the next token $s_{i+1}$ as we discussed and we compare the performance of BERT, RoBERTa [81], XLNet [81], GPT-2 [81], and their combination [81] (please see Table B.1). The citations in this paragraph correspond to the sources that provide data on the performance of the models. For MultiCochrane, T5 models have much better performance than other baselines. For AutoMeTS, a combination of different models actually outperforms any of them, but overall, none of those models have a high accuracy or BLEU score.
B.3 Text Generation
Task Description
Text generation is a broad task that includes or relates to multiple specific tasks such as question answering (Sec A.1) and text summarization (Sec B.1). Here we focus on data-to-text tasks with open (instead of well-defined) answers, which involves taking structured data (e.g. a table) and producing text as output that describes these data. For example, generating patient clinic letters, radiology reports, and medical notes [295].
Datasets, Models, and Discussion
Yermakov et al. [296] introduced a new dataset BioLeaflets and evaluated the performance of multiple LLMs on the data-to-text generation task. They found that T5 [297] and BARD [87] achieve the best performance for this task. However, multiple questions remain. Current LLMs may generate texts with typos, hallucinations, and repetitious words. Also, LLMs are not mature enough to produce coherent long text so far. Ranjit et al. [298] studied the task of generating reports for chest X-ray images. They also investigated and discussed the hallucinations that occurred in the GPT-generated texts. There are concerns about using model-generated pseudo-text for attacking since humans without expert knowledge cannot easily see the factual errors in the generated text. Rodriguez et al. [299] works on preventing attacks in the biomedical domain.