### v5

# A Survey of Large Language Models in Medicine: Progress, Application, and Challenge

Abstract
Large language models (LLMs), such as ChatGPT, have received substantial attention due to their capabilities for understanding and generating human language. While there has been a burgeoning trend in research focusing on the employment of LLMs in supporting different medical tasks (e.g., enhancing clinical diagnostics, and providing medical education), a comprehensive review of these efforts, particularly their development, practical applications, and outcomes in medicine, remains scarce. Therefore, this review aims to provide a detailed overview of the development and deployment of LLMs in medicine, including the challenges and opportunities they face. In terms of development, we provide a detailed introduction to the principles of existing medical LLMs, including their basic model structures, number of parameters, and sources and scales of data used for model development. It serves as a guide for practitioners in developing medical LLMs tailored to their specific needs. In terms of deployment, we offer a comparison of the performance of different LLMs across various medical tasks, and further compare them with state-of-the-art lightweight models, aiming to provide a clear understanding of the distinct advantages and limitations of LLMs in medicine. Overall, in this review, we address the following study questions: 1) What are the practices for developing medical LLMs 2) How to measure the medical task performance of LLMs in a medical setting? 3) How have medical LLMs been employed in real-world practice? 4) What challenges arise from the use of medical LLMs? and 5) How to more effectively develop and deploy medical LLMs? By answering these questions, this review aims to provide insights into the opportunities and challenges of LLMs in medicine and serve as a practical resource for constructing effective medical LLMs. We also maintain a regularly updated list of practical guides on medical LLMs at: https://github.com/AI-in-Health/MedLLMsPracticalGuide.
1 Introduction
Figure 1: An overview of the practical guides for medical large language models.
The recently emerged general large language models (LLMs) zhao2023survey; yang2023harnessing, such as PaLM paLM, LLaMA llama; touvron2023llama2, GPT-series GPT-3; gpt-4, and ChatGLM du2022glm; zeng2022glm, have advanced the state-of-the-art in various natural language processing (NLP) tasks, including text generation, text summarization, and question answering. Inspired by these successes, several endeavors have been made to adapt general LLMs to the medicine domain, leading to the emergence of medical LLMs medpalm; medpalm2. For example, based on PaLM paLM and GPT-4 gpt-4, MedPaLM-2 medpalm2 and MedPrompt nori2023can have respectively achieved a competitive accuracy of 86.5 and 90.2 compared to human experts (87.0 wu2023pmc) in the United States Medical Licensing Examination (USMLE) medqa. In particular, based on publicly available general LLMs (e.g. LLaMA llama; touvron2023llama2), a wide range of medical LLMs, including ChatDoctor li2023chatdoctor, MedAlpaca han2023medalpaca, PMC-LLaMA wu2023pmc, BenTsao wang2023huatuo, and Clinical Camel toma2023clinical, have been introduced. As a result, medical LLMs have gained growing research interests in assisting medical professionals to improve patient care thirunavukarasu2023large; patel2023chatgpt.
Although existing medical LLMs have achieved promising results, there are some key issues in their development and application that need to be addressed. First, many of these models primarily focus on medical dialogue and medical question-answering tasks, but their practical utility in clinical practice is often overlooked thirunavukarasu2023large. Recent research and reviews thirunavukarasu2023large; omiye2024large; clusmann2023future have begun to explore the potential of medical LLMs in different clinical scenarios, including Electronic Health Records (EHRs) yang2022large, discharge summary generation patel2023chatgpt, health education abd2023large, and care planning medpalm2. However, they primarily focus on presenting clinical applications of LLMs, especially online commercial LLMs like ChatGPT (including GPT-3.5 and GPT-4 gpt-4), without providing practical guidelines for the development of medical LLMs. Besides, they mainly perform case studies to conduct the human evaluation on a small number of samples, thus lacking evaluation datasets for assessing model performance in clinical scenarios. Second, most existing medical LLMs report their performances mainly on answering medical questions, neglecting other biomedical domains, such as medical language understanding and generation. These research gaps motivate this review which offers a comprehensive review of the development of LLMs and their applications in medicine. We aim to cover topics on existing medical LLMs, various medical tasks, clinical applications, and arising challenges.
As shown in Figure 1, this review seeks to answer the following questions. Section 2: What are LLMs? How can medical LLMs be effectively built? Section 3: How are the current medical LLMs evaluated? What capabilities do medical LLMs offer beyond traditional models? Section 4: How should medical LLMs be applied in clinical settings? Section 5: What challenges should be addressed when implementing medical LLMs in clinical practice? Section 6: How can we optimize the construction of medical LLMs to enhance their applicability in clinical settings, ultimately contributing to medicine and creating a positive societal impact?
For the first question, we analyze the foundational principles underpinning current medical LLMs, providing detailed descriptions of their architecture, parameter scales, and the datasets used during their development. This exposition aims to serve as a valuable resource for researchers and clinicians designing medical LLMs tailored to specific requirements, such as computational constraints, data privacy concerns, and the integration of local knowledge bases. For the second question, we evaluate the performance of medical LLMs across ten biomedical NLP tasks, encompassing both discriminative and generative tasks. This comparative analysis elucidates how these models outperform traditional AI approaches, offering insights into the specific capabilities that render LLMs effective in clinical environments. The third question, the practical deployment of medical LLMs in clinical settings, is explored through the development of guidelines tailored for seven distinct clinical application scenarios. This section outlines practical implementations, emphasizing specific functionalities of medical LLMs that are leveraged in each scenario. The fourth question emphasizes addressing the challenges associated with the clinical deployment of medical LLMs, such as the risk of generating factually inaccurate yet plausible outputs (hallucination), and the ethical, legal, and safety implications. Citing recent studies, we argue for a comprehensive evaluation framework that assesses the trustworthiness of medical LLMs to ensure their responsible and effective utilization in healthcare. For the last question, we propose future research directions to advance the medical LLMs field. This includes fostering interdisciplinary collaboration between AI specialists and medical professionals, advocating for a ’doctor-in-the-loop’ approach, and emphasizing human-centered design principles.
By establishing robust training data, benchmarks, metrics, and deployment strategies through co-development efforts, we aim to accelerate responsible and efficacious integration of medical LLMs into clinical practice. This study therefore seeks to stimulate continued research and development in this interdisciplinary field, with the objective of realizing the profound potential of medical LLMs in enhancing clinical practice and advancing medical science for for the betterment of society.
BOX 1: Background of Large Language Models (LLMs)
  The impressive performance of LLMs can be attributed to Transformer-based language models, large-scale pre-training, and scaling laws. Language Models    A language model bengio2000neural; mikolov2010recurrent; sundermeyer2015feedforward is a probabilistic model that models the joint probability distribution of tokens (meaningful units of text, such as words or subwords or morphemes) in a sequence, i.e., the probabilities of how words and phrases are used in sequences. Therefore, it can predict the likelihood of a sequence of tokens given the previous tokens, which can be used to predict the next token in a sequence or to generate new sequences. The Transformer architecture    The recurrent neural network (RNN) Hochreiter1997LSTM; mikolov2010recurrent has been widely used for language modeling by processing tokens sequentially and maintaining a vector named hidden state that encodes the context of previous tokens. Nonetheless, sequential processing makes it unsuitable for parallel training and limits its ability to capture long-range dependencies, making it computationally expensive and hindering its learning ability for long sequences. The strength of the Transformer transformer lies in its fully attentive mechanism, which relies exclusively on the attention mechanism and eliminates the need for recurrence. When processing each token, the attention mechanism computes a weighted sum of the other input tokens, where the weights are determined by the relevance between each input token and the current token. It allows the model to adaptively focus on different parts of the sequence to effectively learn the joint probability distribution of tokens. Therefore, Transformer not only enables efficient modeling of long-text but also allows highly paralleled training devlin2018bert, thus reducing training costs. They make the Transformer highly scalable, and therefore it is efficient to obtain LLMs through the large-scale pre-training strategy. Large-scale Pre-training    The LLMs are trained on massive corpora of unlabeled texts (e.g., CommonCrawl, Wiki, and Books) to learn rich linguistic knowledge and language patterns. The common training objectives are masked language modeling (MLM) and next token prediction (NTP). In MLM, a portion of the input text is masked, and the model is tasked with predicting the masked text based on the remaining unmasked context, encouraging the model to capture the semantic and syntactic relationships between tokens devlin2018bert; NTP is another common training objective, where the model is required to predict the next token in a sequence given the previous tokens. It helps the model to predict the next token GPT-3. Scaling Laws    LLMs are the scaled-up versions of Transformer architecture transformer with increased numbers of Transformer layers, model parameters, and volume of pre-training data. The “scaling laws” kaplan2020scaling; hoffmann2022training predict how much improvement can be expected in a model’s performance as its size increases (in terms of parameters, layers, data, or the amount of training computed). The scaling laws proposed by OpenAI kaplan2020scaling show that to achieve optimal model performance, the budget allocation for model size should be larger than the data. The scaling laws proposed by Google DeepMind hoffmann2022training show that both model and data sizes should be increased in equal scales. The scaling laws guide researchers to allocate resources and anticipate the benefits of scaling models. General Large Language Models    Existing general LLMs can be divided into three categories based on their architecture (Table 1). Encoder-only LLMs    consisting of a stack of Transformer encoder layers, employ a bidirectional training strategy that allows them to integrate context from both the left and the right of a given token in the input sequence. This bi-directionality enables the models to achieve a deep understanding of the input sentences devlin2018bert. Therefore, encoder-only LLMs are particularly suitable for language understanding tasks (e.g., sentiment analysis document classification) where the full context of the input is essential for accurate predictions. BERT devlin2018bert and DeBERTa he2021deberta are the representative encoder-only LLMs. Decoder-only LLMs    utilize a stack of Transformer decoder layers and are characterized by their uni-directional (left-to-right) processing of text, enabling them to generate language sequentially. This architecture is trained unidirectionally using the next token prediction training objective to predict the next token in a sequence, given all the previous tokens. After training, the decoder-only LLMs generate sequences autoregressively (i.e. token-by-token). The examples are the GPT-series developed by OpenAI GPT-3; gpt-4, the LLaMA-series developed by Meta llama; touvron2023llama2, and the PaLM paLM and Bard (Gemini) google2023bard developed by Google. Based on the LLaMA model, Alpaca alpaca is fine-tuned with 52k self-instructed data supervision. In addition, Baichuan yang2023baichuan is trained on approximately 1.2 trillion tokens that support bilingual communication in Chinese and English. These LLMs have been used successfully in language generation. Encoder-decoder LLMs    are designed to simultaneously process input sequences and generate output sequences. They consist of a stack of bidirectional Transformer encoder layers followed by a stack of unidirectional Transformer decoder layers. The encoder processes and understands the input sequences, while the decoder generates the output sequences du2022glm; zeng2022glm; chung2022scaling. Representative examples of encoder-decoder LLMs include Flan-T5 joseph2023multilingual, and ChatGLM du2022glm; zeng2022glm. Specifically, ChatGLM du2022glm; zeng2022glm has 6.2B parameters and is a conversational open-source LLM specially optimized for Chinese to support Chinese-English bilingual question-answering.
Table 1: Summary of existing general (large) language models, their underlying structures, numbers of parameters, and datasets used for model training. Column “# params” shows the number of parameters, M: million, B: billion. Domains Model Structures Models # Params Pre-train Data Scale General-domain (Large) Language Models Encoder-only BERT devlin2018bert 110M/340M 3.3B tokens RoBERTa liu2019roberta 355M 161GB DeBERTahe2021deberta 1.5B 160GB Decoder-only GPT-2gpt-2 1.5B 40GB Vicunavicuna 7B/13B LLaMA + 70K dialogues Alpacaalpaca 7B/13B LLaMA+ 52K IFT Mistraljiang2023mistral 7B - LLaMA llama 7B/13B/33B/65B 1.4T tokens LLaMA-2 touvron2023llama2 7B/13B/34B/70B 2T tokens LLaMA-3 meta_llama3 8B/70B 15T tokens GPT-3GPT-3 6.7B/13B/175B 300B tokens Qwenbai2023qwen 1.8B/7B/14B/72B 3T tokens PaLM paLM 8B/62B/540B 780B tokens FLAN-PaLM chung2022scaling 540B - Gemini (Bard) google2023bard - - GPT-3.5 instructgpt - - GPT-4gpt-4 - - Claude-3Claude2024 - - Encoder-Decoder BART lewis2019bart 140M/400M 160GB ChatGLM du2022glm; zeng2022glm 6.2B 1T tokens T5 joseph2023multilingual 11B 1T tokens FLAN-T5 chung2022scaling 3B/11B 780B tokens UL2 tay2022ul2 19.5B 1T tokens GLM zeng2022glm 130B 400B tokens
Table 2: Summary of existing medical-domain LLMs, in terms of their model development, the number of parameters (# params), the scale of pre-training/fine-tuning data, and the data source. M: million, B: billion.
Domains Model Development Models # Params Data Scale Data Source
Medical-domain LLMs (Sec. 2) Pre-training (Sec. 2.1) BioBERT lee2020biobert 110M 18B tokens PubMed PubMed+PMC PMC
PubMedBERT gu2021domain 110M/340M 3.2B tokens PubMed PubMed+PMC PMC
SciBERT beltagy2019scibert 110M 3.17B tokens Literature ammar2018construction
ClinicalBERT alsentzer2019publicly 110M 112k clinical notes MIMIC-III johnson2016mimic
BioM-ELECTRA alrowili2021large 110M/335M - PubMed PubMed
BioMed-RoBERTa domains 125M 7.55B tokens S2ORC lo2019s2orc
BioLinkBERT yasunaga2022linkbert 110M/340M 21GB PubMed PubMed
SciFive phan2021scifive 220M/770M - PubMed PubMed+PMC PMC
ClinicalT5 lu2022clinicalt5 220M/770M 2M clinical notes MIMIC-III johnson2016mimic
BlueBERT peng2019transfer; mutinda2020detecting; mahajan2020identification 110M/340M >4.5B tokens
PubMedPubMed+MIMIC-IIIjohnson2016mimic
MedCPT jin2023biocpt 330M 255M articles PubMed PubMed
BioGPT luo2022biogpt 1.5B 15M articles PubMed PubMed
BioMedLM venigalla2022biomedlm 2.7B 110GB Pile pile
OphGLM gao2023ophglm 6.2B 20k dialogues MedDialog chen2020meddialog
GatorTron yang2022large 8.9B
>82B tokens+6B tokens
2.5B tokens+0.5B tokens
EHRsyang2022large+PubMedPubMed
Wiki+MIMIC-III johnson2016mimic
GatorTronGPT peng2023study 5B/20B 277B tokens EHRs peng2023study
Fine-tuning (Sec. 2.2) DoctorGLM xiong2023doctorglm 6.2B 323MB dialogues CMD. Chinese-medical-dialogue-data
BianQuechen2023bianque 6.2B 2.4M dialogues BianQueCorpus chen2023bianque
ClinicalGPT wang2023clinicalgpt 7B
96k EHRs
192 medical QA
100k dialogues
MD-EHR wang2023clinicalgpt
VariousMedQA medqa
MedDialog chen2020meddialog
Qilin-Med ye2023qilin 7B 3GB ChiMed ye2023qilin
ChatDoctorli2023chatdoctor 7B 110k dialogues
HealthCareMagic healthcaremagic+iCliniq icliniq
BenTsao wang2023huatuo 7B 8k instructions CMeKG-8K byambasuren2019preliminary
HuatuoGPT zhang2023huatuogpt 7B 226k instructions&dialogues Hybrid SFT zhang2023huatuogpt
Baize-healthcare xu2023baize 7B 101K dialogues
Quora+MedQuAD abacha2019question
BioMedGPT luo2023biomedgpt 7B >26B tokens S2ORC lo2019s2orc
MedAlpaca han2023medalpaca 7B/13B 160k medical QA Medical Meadow han2023medalpaca
AlpaCare zhang2023alpacare 7B/13B 52k instructions MedInstruct-52k zhang2023alpacare
Zhongjing yang2023zhongjing 13B 70k dialogues CMtMedQA yang2023zhongjing
PMC-LLaMA wu2023pmc 13B 79.2B tokens
Books+Literature lo2019s2orc+MedC-I wu2023pmc
CPLLM shoham2023cpllm 13B 109k EHRs
eICU-CRD pollard2018eicu+MIMIC-IV johnson2020mimiciv
OpenBioLLM OpenBioLLMs 8B/70B -
MEDITRON chen2023meditron; bosselut2024meditron 7B/70B 48.1B tokens
PubMed PubMed+Guidelines chen2023meditron
Clinical Camel toma2023clinical 13B/70B
70k dialogues+100k articles
4k medical QA
ShareGPTsharegpt+PubMedPubMed
MedQA medqa
MedPaLM 2 medpalm2 340B 193k medical QA MultiMedQA medpalm2
Med-Gemini yang2024advancing; saab2024capabilities - -
MedQA-R&RS saab2024capabilities++MultiMedQA medpalm2
+MIMIC-III johnson2016mimic+MultiMedBench tanno2024consensus
Prompting (Sec. 2.3) CodeX lievin2022can GPT-3.5 / LLaMA-2 Chain-of-Thought (CoT) wei2023chainofthought -
DeID-GPT liu2023deid ChatGPT / GPT-4 Chain-of-Thought (CoT) wei2023chainofthought -
ChatCAD wang2023chatcad ChatGPT Zero-shot Prompting -
Dr. Knows gao2023leveraging ChatGPT Zero-shot Prompting UMLS bodenreider2004unified
MedPaLM medpalm PaLM (540B) 40 instructions MultiMedQA medpalm2
MedPrompt nori2023can GPT-4
Few-shot & CoT wei2023chainofthought
Chat-Orthopedist shi2023retrieval ChatGPT
Retrieval-Augmented
Generation (RAG)
PubMed+GuidelinesSRS +
UpToDateUpToDate+DynameDynamed
QA-RAG kim2024rag ChatGPT RAG FDA QA kim2024rag
Almanac zakka2024almanac ChatGPT RAG & CoT Clinical QA zakka2024almanac
Figure 2: We adopt the data from Table 2 to demonstrate the development of model sizes for medical large language models in different model architectures, i.e., BERT-like, ChatGLM/LLaMA-like, and GPT/PaLM-like.
2 The Principles of Medical Large Language Models
Box 1 and Table 1 briefly introduce the background of general LLMs zhao2023survey, e.g., GPT-4 gpt-4. Table 2 summarizes the currently available medical LLMs according to their model development. Existing medical LLMs are mainly pre-trained from scratch, fine-tuned from existing general LLMs, or directly obtained through prompting to align the general LLMs to the medical domain. Therefore, we introduce the principles of medical LLMs in terms of these three methods: pre-training, fine-tuning, and prompting. Meanwhile, we further summarize the medical LLMs according to their model architectures in Figure 2.
2.1 Pre-training
Pre-training typically involves training an LLM on a large corpus of medical texts, including both structured and unstructured text, to learn the rich medical knowledge. The corpus may include EHRs peng2023study, clinical notes yang2022large, and medical literature alsentzer2019publicly. In particular, PubMed PubMed, MIMIC-III clinical notes johnson2016mimic, and PubMed Central (PMC) literature PMC, are three widely used medical corpora for medical LLM pre-training. A single corpus or a combination of corpora may be used for pre-training. For example, PubMedBERT gu2021domain and ClinicalBERT are pre-trained on PubMed and MIMIC-III, respectively. In contrast, BlueBERT peng2019transfer combines both corpora for pre-training; BioBERT lee2020biobert is pre-trained on both PubMed and PMC. The University of Florida (UF) Health EHRs are further introduced in pre-training GatorTron yang2022large and GatorTronGPT peng2023study. MEDITRON chen2023meditron is pre-trained on Clinical Practice Guidelines (CPGs). The CPGs are used to guide healthcare practitioners and patients in making evidence-based decisions about diagnosis, treatment, and management.
To meet the needs of the medical domain, pre-training medical LLMs typically involve refining the following commonly used training objectives in general LLMs: masked language modeling, next sentence prediction, and next token prediction (Please see Box 1 for an introduction of these three pre-training objectives). For example. BERT-series models (e.g., BioBERT lee2020biobert, PubMedBERT gu2021domain, ClinicalBERT alsentzer2019publicly, and GatorTron yang2022large) mainly adopt the masked language modeling and the next sentence prediction for pre-training; GPT-series models (e.g., BioGPT luo2022biogpt, and GatorTronGPT peng2023study) mainly adopt the next token prediction for pre-training. It is worth mentioning that BERT-like Medical LLMs (e.g. BioBERT lee2020biobert, PubMedBERT gu2021domain, Clinical BERT alsentzer2019publicly) are originally derived from the general domain BERT or RoBERTa models. To clarify the differences between different models, in our Table 2, we only show the data source used to further construct medical LLMs. After pre-training, medical LLMs can learn rich medical knowledge that can be leveraged to achieve strong performance on different medical tasks.
2.2 Fine-tuning
It is high-cost and time-consuming to train a medical LLM from scratch, due to its requirement of substantial (e.g. several days or even weeks) computational power and manual labor. One solution is to fine-tune the general LLMs with medical data, and researchers have proposed different fine-tuning methods medpalm2; han2023medalpaca; toma2023clinical for learning domain-specific medical knowledge and obtaining medical LLMs. Current fine-tuning methods include Supervised Fine-Tuning (SFT), Instruction Fine-Tuning (IFT), and Parameter-Efficient Fine-Tuning (PEFT). The resulting fine-tuned medical LLMs are summarized in Table 2.
Supervised Fine-Tuning (SFT)
aims to leverage high-quality medical corpus, which can be physician-patient conversations li2023chatdoctor, medical question-answering han2023medalpaca, and knowledge graphs ye2023qilin; wang2023huatuo. The constructed SFT data serves as a continuation of the pre-training data to further pre-train the general LLMs with the same training objectives, e.g. next token prediction. SFT provides an additional pre-training phase that allows the general LLMs to learn rich medical knowledge and align with the medical domain, thus transforming them into specialized medical LLMs.
The diversity of SFT enables the development of diverse medical LLMs by training on different types of medical corpus. For example, DoctorGLM xiong2023doctorglm and ChatDoctor li2023chatdoctor are obtained by fine-tuning the general LLMs ChatGLM du2022glm; zeng2022glm and LLaMA llama on the physician-patient dialogue data, respectively. MedAlpaca han2023medalpaca based on the general LLM Alpaca alpaca is fine-tuned using over 160,000 medical QA pairs sourced from diverse medical corpora. Clinicalcamel toma2023clinical combines physician-patient conversations, clinical literature, and medical QA pairs to refine the LLaMA-2 model touvron2023llama2. In particular, Qilin-Med ye2023qilin and Zhongjing yang2023zhongjing are obtained by incorporating the knowledge graph to perform fine-tuning on the Baichuan yang2023baichuan and LLaMA llama, respectively.
In summary, existing studies have demonstrated the efficacy of SFT in adapting general LLMs to the medical domain. They show that SFT improves not only the model’s capability for understanding and generating medical text, but also its ability to provide accurate clinical decision support he2023survey.
Instruction Fine-Tuning (IFT)
constructs instruction-based training datasets zhang2023instruction; he2023survey; zhao2023survey, which typically comprise instruction-input-output triples, e.g. instruction-question-answer. The primary goal of IFT is to enhance the model’s ability to follow various human/task instructions, align their outputs with the medical domain, and thereby produce a specialized medical LLM.
Thus, the main difference between SFT and IFT is that the former focuses primarily on injecting medical knowledge into a general LLM through continued pre-training, thus improving its ability to understand the medical text and accurately predict the next token. In contrast, IFT aims to improve the model’s instruction following ability and adjust its outputs to match the given instructions, rather than accurately predicting the next token as in SFT zhang2023instruction. As a result, SFT emphasizes the quantity of training data, while IFT emphasizes their quality and diversity. Since IFT and SFT are both capable of improving model performance, there have been some recent works yang2023zhongjing; ye2023qilin; zhang2023alpacare attempting to combine them for obtaining robust medical LLMs.
In other words, to enhance the performance of LLMs through IFT, it is essential to ensure that the training data for IFT are of high quality and encompass a wide range of medical instructions and medical scenarios. To this end, MedPaLM-2 medpalm2 invited qualified medical professionals to develop the instruction data for fine-tuning the general PaLM. BenTsao wang2023huatuo and ChatGLM-Med ChatGLM-Med constructed the knowledge-based instruction data from the knowledge graph. Zhongjing yang2023zhongjing further incorporated the multi-turn dialogue as the instruction data to perform IFT. MedAlpaca han2023medalpaca simultaneously incorporated the medical dialogues and medical QA pairs for instruction fine-tuning.
Parameter-Efficient Fine-Tuning (PEFT)
aims to substantially reduce computational and memory requirements for fine-tuning general LLMs. The main idea is to keep most of the parameters in pre-trained LLMs unchanged, by fine-tuning only the smallest subset of parameters (or additional parameters) in these LLMs. Commonly used PEFT techniques include Low-Rank Adaptation (LoRA) lora, Prefix Tuning li2021prefixtuning, and Adapter Tuning liu2022p; liu2021p.
In contrast to fine-tuning full-rank weight matrices, 1) LoRA preserves the parameters of the original LLMs and only adds trainable low-rank matrices into the self-attention module of each Transformer layer lora. Therefore, LoRA can substantially reduce the number of trainable parameters and improve the efficiency of fine-tuning, while still enabling the fine-tuned LLM to capture effectively the characteristics of the tasks. 2) Prefix Tuning takes a different approach from LoRA by adding a small set of continuous task-specific vectors (i.e. “prefixes”) to the input of each Transformer layer li2021prefixtuning; zhao2023survey. These prefixes serve as the additional context to guide the generation of the model without changing the original pre-trained parameter weights. 3) Adapter Tuning involves introducing small neural network modules, known as adapters, into each Transformer layer of the pre-trained LLMs houlsby2019parameter. These adapters are fine-tuned while keeping the original model parameters frozen houlsby2019parameter, thus allowing for flexible and efficient fine-tuning. The number of trainable parameters introduced by adapters is relatively small, yet they enable the LLMs to adapt to clinical scenarios or tasks effectively.
In general, PEFT is valuable for developing LLMs that meet unique needs in specific (e.g. medical) domains, due to its ability to reduce computational demands while maintaining the model performance. For example, medical LLMs DoctorGLM xiong2023doctorglm, MedAlpaca han2023medalpaca, Baize-Healthcare xu2023baize, Zhongjing yang2023zhongjing, CPLLM shoham2023cpllm, and Clinical Camel toma2023clinical adopted the LoRA lora to perform parameter-efficient fine-tuning to efficiently align the general LLMs to the medical domain.
2.3 Prompting
Fine-tuning considerably reduces computational costs compared to pre-training, but it requires further model training and collections of high-quality datasets for fine-tuning, thus still consuming some computational resources and manual labor. In contrast, the “prompting” methods efficiently align general LLMs (e.g. PaLM paLM) to the medical domain (e.g., MedPaLM medpalm), without training any model parameters. Popular prompting methods include In-Context Learning (ICL), Chain-of-Thought (CoT) prompting, Prompt Tuning, and Retrieval-Augmented Generation (RAG).
In-Context Learning (ICL)
aims to directly give instructions to prompt the LLM to perform a task efficiently. In general, the ICL consists of four process: task understanding, context learning, knowledge reasoning, and answer generation. First, the model must understand the specific requirements and goals of the task. Second, the model learns to understand the contextual information related to the task with argument context. Then, use the model’s internal knowledge and reasoning capabilities to understand the patterns and logic in the example. Finally, the model generates the task-related answers. The advantage of ICL is that it does not require a large amount of labeled data for fine-tuning. Based on the type and number of input examples, ICL can be divided into three categories dong2022survey. 1) One-shot Prompting: Only one example and task description are allowed to be entered. 2) Few-shot Prompting: Allows the input of multiple instances and task descriptions. 3) Zero-shot Prompting: Only task descriptions are allowed to be entered. ICL presents the LLMs making task predictions based on contexts augmented with a few examples and task demonstrations. It allows the LLMs to learn from these examples or demonstrations to accurately perform the task and follow the given examples to give corresponding answers GPT-3. Therefore, ICL allows LLMs to accurately understand and respond to medical queries. For example, MedPaLM medpalm substantially improves the task performance by providing the general LLM, PaLM paLM, with a small number of task examples such as medical QA pairs.
Chain-of-Thought (CoT) Prompting
further improves the accuracy and logic of model output, compared with In-Context Learning. Specifically, through prompting words, CoT aims to prompt the model to generate intermediate steps or paths of reasoning when dealing with downstream (complex) problems wei2023chainofthought. Moreover, CoT can be combined with few-shot prompting by giving reasoning examples, thus enabling medical LLMs to give reasoning processes when generating responses. For tasks involving complex reasoning, such as medical QA, CoT has been shown to effectively improve model performance medpalm; medpalm2. Medical LLMs, such as DeID-GPT liu2023deid, MedPaLM medpalm, and MedPrompt nori2023can, use CoT prompting to assist them in simulating a diagnostic thought process, thus providing more transparent and interpretable predictions or diagnoses. In particular, MedPrompt nori2023can directly prompts a general LLM, GPT-4 gpt-4, to outperform the fine-tuned medical LLMs on medical QA without training any model parameters.
Prompt Tuning
aims to improve the model performance by employing both prompting and fine-tuning techniques lester2021power; liu2021p. The prompt tuning method introduces learnable prompts, i.e. trainable continuous vectors, which can be optimized or adjusted during the fine-tuning process to better adapt to different medical scenarios and tasks. Therefore, they provide a more flexible way of prompting LLMs than the “prompting alone” methods that use discrete and fixed prompts, as described above. In contrast to traditional fine-tuning methods that train all model parameters, prompt tuning only tunes a very small set of parameters associated with the prompts themselves, instead of extensively training the model parameters. Thus, prompt tuning effectively and accurately responds to medical problems nori2023can, with minimal incurring computational cost.
Existing medical LLMs that employ the prompting techniques are listed in Table 2. Recently, MedPaLM medpalm and MedPaLM-2 medpalm2 propose to combine all the above prompting methods, resulting in Instruction Prompt Tuning, to achieve strong performances on various medical question-answering datasets. In particular, using the MedQA dataset for the US Medical Licensing Examination (USMLE), MedPaLM-2 medpalm2 achieves a competitive overall accuracy of 86.5% compared to human experts (87.0%), surpassing previous state-of-the-art method MedPaLM medpalm by a large margin (19%).
Retrieval-Augmented Generation (RAG)
enhances the performance of LLMs by integrating external knowledge into the generation process. In detail, RAG can be used to minimize LLM’s hallucinations, obscure reasoning processes, and reliance on outdated information by incorporating external database knowledge gao2023retrieval. RAG consists of three main components: retrieval, augmentation, and generation. The retrieval component employs various indexing strategies and input query processing techniques to search and top-ranked relevant information from an external knowledge base. The retrieved external data is then augmented into the LLM’s prompt, providing additional context and grounding for the generated response. By directly updating the external knowledge base, RAG mitigates the risk of catastrophic forgetting associated with model weight modifications, making it particularly suitable for domains with low error tolerance and rapidly evolving information, such as the medical field. In contrast to traditional fine-tuning methods, RAG enables the timely incorporation of new medical information without compromising the model’s previously acquired knowledge, ensuring the generated outputs remain accurate and up-to-date in the face of evolving medical challenges. Most recently, researchers proposed the first benchmark MIRAGE xiong2024benchmarking based on medical information RAG, including 7,663 questions from five medical QA datasets, which has been established to both steer research and facilitate the practical deployment of medical RAG systems
In RAG, retrieval can be achieved by calculating the similarity between the embeddings of the question and document chunks, where the semantic representation capability of embedding models plays a key role. Recent research has introduced prominent embedding models such as AngIEli2023angle, Voyagewang2023voyager, and BGEbge_m3. In addition to embedding, the retrieval process can be optimized via various strategies such as adaptive retrieval, recursive retrieval, and iterative retrieval shao2023enhancing; trivedi2022interleaving; asai2023self. Several recent works have demonstrated the effectiveness of RAG in medical and pharmaceutical domains. Almanaczakka2024almanac is a large language framework augmented with retrieval capabilities for medical guidelines and treatment recommendations, surpassing the performance of ChatGPT on clinical scenario evaluations, particularly in terms of completeness and safety. Another work QA-RAG kim2024rag employs RAG with LLM for pharmaceutical regulatory tasks, where the model searches for relevant guideline documents and provides answers based on the retrieved guidelines. Chat-Orthopedistshi2023retrieval, a retrieval-augmented LLM, assists adolescent idiopathic scoliosis (AIS) patients and their families in preparing for meaningful discussions with clinicians by providing accurate and comprehensible responses to patient inquiries, leveraging AIS domain knowledge.
2.4 Discussion
This section discusses the principles of medical LLMs, including three types of methods for building models: pre-training, fine-tuning, and prompting. To meet the needs of practical medical applications, users can choose proper medical LLMs according to the magnitude of their own computing resources. Companies or institutes with massive computing resources can either pre-train an application-level medical LLM from scratch or fine-tune existing open-source general LLM models (e.g. LLaMA meta_llama3) using large-scale medical data. The results in existing literature (e.g. Med-PaLM2 medpalm2, MedAlpaca han2023medalpaca and Clinical Camel toma2023clinical) have shown that fine-tuning general LLMs on medical data can boost their performance of medical tasks. For example, Clinical Camel toma2023clinical, which is fine-tuned on the LLaMA-2-70B touvron2023llama2 model, even outperforms GPT-4 toma2023clinical. However, for small enterprises or individuals with certain computing resources, combining with the understanding of medical tasks and a reasonable combination of ICL, prompting engineering, and RAG, to prompt black-box LLMs may also achieve miraculous results. For example, MedPrompt nori2023can stimulates the commercial LLM GPT-4 gpt-4 through an appropriate combination of prompt strategies to achieve comparable or even better results than fine-tuned medical LLMs (e.g. Med-PaLM2 medpalm2) and human experts, suggesting that a mix of prompting strategies is an efficient and green solution in the medical domain rather than fine-tuning.
Figure 3: Performance (Dataset-Metric (Task)) comparison between the GPT-3.5 turbo, GPT-4, state-of-the-art task-specific lightweight models (Fine-tuned), and human experts, on seven medical tasks across eleven datasets. All data presented in our Figures originates from published and peer-reviewed literature. Please refer to the supplementary material for the detailed data.
Figure 4: We demonstrate the development of medical large language models over time in different model development types through the scores of the United States Medical Licensing Examination (USMLE) from the MedQA dataset.
3 Medical Tasks
In this section, we will introduce two popular types of medical machine learning tasks: generative and discriminative tasks, including ten representative tasks that further build up clinical applications. Figure 3 illustrates the performance comparisons between different LLMs. For clarity, we will only cover a general discussion of those tasks. The detailed definition of the task and the performance comparisons can be found in our supplementary material.
3.1 Discriminative Tasks
Discriminative tasks are for categorizing or differentiating data into specific classes or categories based on given input data. They involve making distinctions between different types of data, often to categorize, classify, or extract relevant information from structured text or unstructured text. The representative tasks include Question Answering, Entity Extraction, Relation Extraction, Text Classification, Natural Language Inference, Semantic Textual Similarity, and Information Retrieval.
The typical input for discriminative tasks can be medical questions, clinical notes, medical documents, research papers, and patient EHRs. The output can be labels, categories, extracted entities, relationships, or answers to specific questions, which are often structured and categorized information derived from the input text. In existing LLMs, the discriminative tasks are widely studied and used to make predictions and extract information from input text.
For example, based on medical knowledge, medical literature, or patient EHRs, the medical question answering (QA) task can provide precise answers to clinical questions, such as symptoms, treatment options, and drug interactions. This can help clinicians make more efficient and accurate diagnoses medpalm; medpalm2; thirunavukarasu2023large. Entity extraction can automatically identify and categorize critical information (i.e. entities) such as symptoms, medications, diseases, diagnoses, and lab results from patient EHRs, thus assisting in organizing and managing patient data. The following entity linking task aims to link the identified entities in a structured knowledge base or a standardized terminology system, e.g., SNOMED CT donnelly2006snomed, UMLS bodenreider2004unified, or ICD codes world1978international. This task is critical in clinical decision support or management systems, for better diagnosis, treatment planning, and patient care.
Performance Comparisons
Figure 3 shows that some existing general LLMs (e.g. GPT-3.5-turbo and GPT-4 gpt-4) have achieved strong performance on existing medical machine learning tasks. This is most noticeable for the QA task where GPT-4 (shown in the blue line in Figure 3) consistently outperforms existing task-specific fine-tuned models and is even comparable to human experts (shown in the purple line). The QA datasets of evaluation include MedQA (USMLE) medqa, PubMedQA PubmedQA, and MedMCQA Medmcqa. To better understand the QA performance of existing medical LLMs, in Figure 4, we further demonstrate the QA performance of medical LLMs on the MedQA dataset over time in different model development types. It also clearly shows that current works, e.g., MedPrompt nori2023can, have successfully proposed several prompting methods to enable LLMs to outperform human experts.
However, on the non-QA tasks, as shown in Figure 3, the existing general LLMs perform worse than the task-specific fine-tuned models. For example, on the entity extraction task using the NCBI disease dataset dougan2014ncbi, the state-of-the-art task-specific fine-tuned model BioBERT lee2020biobert achieves an F1 score of 89.36, substantially exceeding the F1 score of 56.73 by GPT-4. We hypothesize that the reason for the strong QA capability of the current general LLMs is that the QA task is close-ended; i.e. the correct answer is already provided by multiple candidates. In contrast, most non-QA tasks are open-ended where the model has to predict the correct answer from a large pool of possible candidates, or even without any candidates provided.
Overall, the comparison proves that the current general LLMs have a strong question-answering capability, however, the capability on other tasks still needs to be improved. Therefore, we advocate that the evaluation of medical LLMs should be extended to a broad range of tasks including non-QA tasks, instead of being limited mainly to medical QA tasks. Hereafter, we will discuss specific clinical applications of LLMs, followed by their challenges and future directions.
3.2 Generative Tasks
Different from discriminative tasks that focus on understanding and categorizing the input text, generative tasks require a model to accurately generate fluent and appropriate new text based on given inputs. These tasks include medical text summarization tang2023evaluating; van2023clinical, medical text generation luo2022biogpt, and medical text simplification ondov2022survey.
For medical text summarization, the input and output are typically long and detailed medical text (e.g. “Findings” in radiology reports), and a concise summarized text (e.g., the “Impression” in radiology reports). Such text contains important medical information that enables clinicians and patients to efficiently capture the key points without going through the entire text. It can also help medical professionals to draft clinical notes by summarizing patient information or medical histories.
In medical text generation, e.g. discharge instruction generation liu2022retrieve, the input can be medical conditions, symptoms, patient demographics, or even a set of medical notes or test results. The output can be a diagnosis recommendation of a medical condition, personalized instructional information, or health advice for the patient to manage their condition outside the hospital.
Medical text simplification ondov2022survey aims to generate a simplified version of the complex medical text by, for example, clarifying and explaining medical terms. Different from text summarization, which concentrates on giving shortened text while maintaining most of the original text meanings, text simplification focuses more on the readability part. In particular, complicated or opaque words will be replaced; complex syntactic structures will be improved; and rare concepts will be explained joseph2023multilingual. Thus, one example application is to generate easy-to-understand educational materials for patients from complex medical texts. It is useful for making medical information accessible to a general audience, without altering the essential meaning of the texts.
Figure 5: Integrated overview of potential applications gao2023leveraging; dong2022automated; d2022emotion; biri2023assessing; vaidyam2019chatbots of large language models in medicine.
Table 3: Summary of existing medical LLMs tailored to various clinical applications, in terms of their architecture, model development, the number of parameters, the scale of PT/FT data, and the data source. M: million, B: billion. PT: Pre-training. FT: Fine-tuning. ICL: In-Context Learning. CoT: Chain-of-Thought prompting. RAG: Retrieval-Augmented Generation.
Application Model Architecture Model Development # Params Data Scale Data Source
Medical Diagnosis (Sec. 4.1) Dr. Knowsgao2023leveraging GPT-3.5 ICL 154B 5820 notes MIMIC-IIIjohnson2016mimic+IN-HOUSE gao2023leveraging
DDx PaLM-2mcduff2023towards PaLM-2 FT & ICL 340B - MultiMedQAmedpalm2+MIMIC-IIIjohnson2016mimic
Med-Geminisaab2024capabilities Gemini FT & CoT - -
MedQA-R/RSsaab2024capabilities+MultiMedQAmedpalm2
MIMIC-IIIjohnson2016mimic+MultiMedBench tanno2024consensus
Med-Flamingomoor2023med ViT+LLaMA-7B FT - 600k pairs
Multimodal Textbookmoor2023med+PMC-OAmoor2023med
VQA-RADlau2018dataset+PathVQAhe2020pathvqa
LLaVA-Medli2024llava ViT+LLaMA-7B FT - 660k pairs
PMC-15Mli2024llava+VQA-RADlau2018dataset
SLAKEliu2021slake+PathVQAhe2020pathvqa
Formatting & ICD-Coding (Sec. 4.2) PLM-ICDhuang2022plm RoBERTa FT 355M 70,539 notes MIMIC-IIsaeed2002mimic+MIMIC-IIIjohnson2016mimic
DRG-LLaMAwang2024drg LLaMA-7B FT 7B 25k pairs MIMIC-IVjohnson2019mimic
ChatICDliu2023chaticd ChatGPT ICL - 10k pairs MIMIC-IIIjohnson2016mimic
LLM-codexyang2023surpassing ChatGPT+LSTM ICL - - MIMIC-IIIjohnson2016mimic
Clinical Report Generation (Sec. 4.3) ImpressionGPTma2024iterative ChatGPT ICL & RAG 110M 184k reports MIMIC-CXRjohnson2019mimic+IU X-rayOpeni
RadAdaptvan2023radadapt T5 FT 223M, 738M 80k reports MIMIC-IIIjohnson2016mimic
ChatCADwang2023chatcad GPT-3 ICL 175B 300 reports MIMIC-CXR johnson2019mimic
MAIRA-1hyland2023maira ViT+Vicuna-7B FT 8B 337k pairs MIMIC-CXRjohnson2019mimic
RadFMwu2023towards ViT+LLaMA-13B PT& FT 14B 32M pairs MedMDwu2023towards
Medical Robotics (Sec. 4.4) SuFIAmoghani2024sufia GPT-4 ICL - 4 tasks ORBIT-Surgical yu2024orbit
UltrasoundGPTxu2024enhancing GPT-4 ICL - 522 tasks -
Robotic X-raykilleen2024take GPT-4 ICL - - -
Medical Language Translation (Sec. 4.5) Medical mT5 garcia2024medical T5 PT 738M, 3B 4.5B pairs
PubMed PubMed+EMEA tiedemann2012parallel
ClinicalTrialsNLM2022a, etc.
Apollowang2024apollo Qwen PT & FT 1.8B-7B 2.5B pairs ApolloCorporawang2024apollo
BiMediXpieri2024bimedix Mistral FT 13B 1.3M pairs BiMed1.3Mpieri2024bimedix
Biomed-sumtang2023improving BART FT 406M 27k papers BioCiteDBtang2023improving
RALLguo2024retrieval BART FT & RAG 406M 63k pairs CELLStang2023improving
Medical Education (Sec. 4.6) ChatGPTopenai2023chatgpt GPT-3.5/GPT-4 ICL - - -
Med-Geminisaab2024capabilities Gemini FT & CoT - -
MedQA-R/RS saab2024capabilities+MultiMedQAmedpalm2
MIMIC-IIIjohnson2016mimic+MultiMedBench tanno2024consensus
Mental Health Support (Sec. 4.7) PsyChatqiu2023psychat ChatGLM FT 6B 350k pairs Xinglingqiu2023psychat+Smilechatqiu2023psychat
ChatCounselorliu2023chatcounselor Vicuna FT 7B 8k instructions Psych8Kliu2023chatcounselor
Mental-LLMxu2024mental Alpaca, FLAN-T5 FT & ICL 7B, 11B 31k pairs
Dreadditturcan2019dreaddit+DepSeveritynaseem2022early+SDCNLhaque2021deep
CSSRS-Suicidegaur2019knowledge+Red-Samsampath2022data
Twt-60Usersjamil2017monitoring+SADmauriello2021sad
Medical Inquiry and Response (Sec. 4.8) AMIEtu2024towards PaLM2 FT 340B >2M pairs
MedQAmedqa+MultiMedBenchtanno2024consensus
MIMIC-IIIjohnson2016mimic+real-world diagloguetu2024towards
Healthcare Copilotren2024healthcare ChatGPT ICL - - MedDialog ren2024healthcare
Conversational Diagnosissun2024conversational GPT-4/LLaMA ICL - 40k pairs MIMIC-IVjohnson2020mimiciv
4 Clinical Applications
As shown in Figure 5, this section discusses the clinical applications of LLMs. Each subsection contains a specific application and discusses how LLMs perform this task. Table 3 summarizes the guidelines on how to select, build, and evaluate medical LLMs for various clinical applications.
4.1 Medical Diagnosis
Medical diagnosis involves the medical practitioner using objective medical data from tests and self-described subjective symptoms to conclude the most likely health problem occurring in the patient AImed. This heavily relies on the synthesis and interpretation of vast amounts of information from various sources, including patient medical histories, clinical data, and the latest medical literature. The advent of large language models has opened up new opportunities for enhancing medical diagnostic processes. These advanced natural language processing models can rapidly process and comprehend massive volumes of medical data, literature, and legal guidelines, potentially aiding healthcare professionals in making more informed and legally sound diagnostic decisionsAImed; thirunavukarasu2023large.
Guideline
Dr. Knowsgao2023leveraging extracts knowledge graph from the unified medical language system (UMLS) bodenreider2004unified for identifying likely diagnoses from patient medical records. It is then further fine-tuned T5 models using these diagnoses as prompts and leveraged zero-shot prompting for ChatGPT, demonstrating improved diagnosis prediction accuracy by utilizing the knowledge graph. An alternative is DDx PaLM-2mcduff2023towards, which builds on Google’s PaLM-2 and is further fine-tuned with medical datasets for zero-shot prompting. It can engage in conversations and collaborate with clinicians, allowing them to interactively identify potential diagnoses for complex medical conditions while improving the doctors’ diagnostic reasoning abilities through the interactive process. Recent advancement in multimodal LLMs allows AI to understand more complex medical knowledge and data with multimodality zhao2023clip. Notable examples include Med-Flamingomoor2023med, LLaVA-Medli2024llava, and Med-Geminisaab2024capabilities. Med-Flamingomoor2023med builds on Google’s vision-language model Flamingo and is pret-rained with medical image-text data. LLaVA-Medli2024llava, proposed by Microsoft, builds on the LLaVA model and is trained with a two-stage medical concept alignment and medical instruction tuning using 660k samples. Google’s powerful Med-Geminisaab2024capabilities builds upon the Gemini foundation and is further fine-tuned with large-scale medical multimodal data, highlighting its capabilities in multimodal understanding and long-context processing. In training and evaluating these models, commonly utilized datasets include the MIMIC series, such as MIMIC-IIIjohnson2016mimic and MIMIC-IVjohnson2020mimiciv discharge summaries, as well as question answering datasets such as MultiMedQAsinghal2023large, which constructed from various resources covering patient cases from PubMed articles and textbooks. For multimodal data, VQA-RAD (radiology)lau2018dataset, SLAKE (radiology)liu2021slake, and PathVQA (pathology)he2020pathvqa are frequently employed. Most benchmarking efforts involve both quantitative evaluation metrics and human evaluations. These models have demonstrated their effectiveness and the potential for substantial improvements in medical diagnosis tasks.
Discussion
One distinct limitation of using LLMs as the sole tool for medical diagnosis is the heavy reliance on subjective text inputs from the patient. Since LLMs are text-based, they lack the inherent capability to analyze medical diagnostic imagery. Given that objective medical diagnoses frequently depend on visual images, LLMs are often unable to directly conduct diagnostic assessments as they lack concrete visual evidence to support disease diagnosis liu2021PPKED. However, they can help with diagnosis as a logical reasoning tool for improving accuracy in other vision-based models. One such example is ChatCAD wang2023chatcad, where images are first fed into an existing computer-aided diagnosis (CAD) model to obtain tensor outputs. These outputs are translated into natural language, which is subsequently fed into ChatCAD to summarize results and formulate diagnoses. ChatCAD achieves a recall score of 0.781, substantially higher than that (0.382) of the state-of-the-art task-specific model. Nevertheless, all the aforementioned methods of implementing LLMs cannot directly process images; instead, they either rely on transforming images into text beforehand or rely on an external separate vision encoder to embed images.
4.2 Formatting and ICD-Coding
The international classification of diseases (ICD) world1978international is a method of standardizing diagnostic and procedural information of a clinical session. These ICD codes are recorded in the patient’s EHRs every doctor visit. They are also used for tracking health metrics, treatment outcomes, and billing. There is a need to automate the ICD labeling process because its manual entry process is very time-consuming for doctors. Formatting and ICD-Coding usually involve entity extraction, relation extraction, text generation, and information retrieval. LLMs can help automate ICD coding by extracting medical terms from clinical notes and assigning corresponding ICD codes ong2023applying; dong2022automated.
Guideline
For example, PLM-ICDhuang2022plm builds upon the RoBERTa model liu2019roberta, fine-tuning it specifically for ICD coding and achieving strong performance on 70,539 notes from the MIMIC-II and MIMIC-III datasets johnson2016mimic, as evaluated by accuracy. The base model used in PLM-ICD is domain-specific with medicine-specific knowledge to enhance the ability to understand medical terms. PLM-ICD uses segment pooling, the algorithm that divides long input texts into shorter representations using LLMs when the input surpasses the maximum allowable length. Lastly, it relates the encoding to the augmented labels to output ICD codes for each clinical input. PLM-ICD produced a higher AUC score than previous state-of-the-art lightweight models huang2022plm. DRG-LLaMAwang2024drg leverages the LLaMA model and applies parameter-efficient fine-tuning techniques, such as LoRA, to adapt the model to this task. ChatICDliu2023chaticd and LLM-codexyang2023surpassing both utilize the ChatGPT model with prompts for ICD coding. However, LLM-codexyang2023surpassing takes this a step further by training an LSTM model on top of the ChatGPT responses, demonstrating its strong performance. ICD coding can be formulated as a multi-label classification task, and most work in this area utilizes the MIMIC-III dataset for training and evaluation. Models are typically assessed based on their F1 score, AUC, and Precision@k, considering either the top 50 most frequent labels or the full label set.
Discussion
One challenge while deploying LLMs for clinical coding is the potential biases and hallucinations. In particular, traditional multi-label classification models can easily constrain their outputs to a predefined list of (usually >1000) ICD candidate codes through a classification neural network. In contrast, generative LLMs can suffer from major hallucinations while the input text is lengthy. As a result, the LLM may assign an ICD code that is not in the candidate list or a non-existent ICD code to the input text. It leads to confusion when interpreting medical records yang2022large and is, therefore, crucial to establish a proactive mechanism to detect and rectify errors before they enter patient EHRs.
4.3 Clinical Report Generation
Clinical reports, such as radiology reports, discharge summaries, and patient clinic letters, refer to standardized documentation that healthcare workers complete after each patient visit liu2019reporting. Therefore, clinical report generation usually involves text generation/summarization, and information retrieval. A large portion of the report is often medical diagnostic results. It is typically tedious for overworked clinicians to write clinical reports, and thus they are often incomplete or error-prone. Meanwhile, LLMs can be used intuitively as a summarization tool to help with clinical report generation. In this instance, LLMs act as an assistant tool for clinicians which helps improve efficiency and reduce potential errors in lengthy reports patel2023chatgpt; ali2023using; liu2022retrieve.
Another popular approach to generating clinical reports using LLMs involves incorporating a vision-based model to provide complementary information wang2023chatcad; wu2023towards; wu2023can. The vision model analyzes the input medical image and generates an annotation, which serves as a direct and supplementary input to the LLM alongside additional text prompts. By leveraging the combination of visual and textual information, the LLM produces accurate and fluent reports that adhere to the specified parameters and structure.
Guideline
For radiology report generation, general medical vision-language models like Med-Geminisaab2024capabilities, LlaVA-Medli2024llava, and Med-Flamingomoor2023med can be utilized, but models trained specifically on radiographs, such as ChatCADwang2023chatcad, MAIRA-1hyland2023maira and RadFMwu2023towards, have shown superior performance than the general medical ones. Alternatively, language models like ImpressionGPTma2024iterative leverage textual data for report summarization, employing dynamic prompt generation and iterative optimization. RadAdaptvan2023radadapt systematically evaluates various language models and lightweight adaptation methods, achieving optimal performance through pre-training on clinical text and parameter-efficient fine-tuning with LoRA, investigating the impact of few-shot prompting. In terms of evaluation, most work uses MIMIC-III or MIMIC-IV notes for training and evaluation, as it is the largest publicly available free-text EHRs. Common automatic evaluation metrics includes lexical methods such as BLEUpapineni2002bleu, ROUGE lin2004rouge, METEORbanerjee2005meteor, semantic-based methods such as BERTScorezhang2019bertscore, and radiology-specific metrics such as CheXbert similaritysmit2020chexbert, RadGraphjain2021radgraph, RadCliQyu2023evaluating.
Discussion
While LLMs have demonstrated the ability to generate clinical reports that are more comprehensive and precise than those written by human counterpartsvan2023clinical, they still face challenges in terms of hallucinations and literal interpretation of inputs, lacking the assumption-based perspective often employed by human doctors. Moreover, LLM-generated reports tend to be less concise compared to human-written ones. The evaluation of LLMs in this domain is particularly challenging due to the specialized nature of the content and the generative nature of the task. Current automatic evaluation methods for clinical report generation primarily focus on lexical metrics, which can lead to biased and inaccurate assessments of the contextual information present in the reports xie2023faithful. For instance, consider two sentences with similar meanings but different wordings: “The patient’s blood glucose level is within normal limits” and “The patient does not exhibit signs of hyperglycemia”. While both convey the absence of hyperglycemia, lexical evaluation metrics may struggle to accurately capture their semantic equivalence, as they rely on direct word-level comparisons. This discrepancy highlights the need for more sophisticated evaluation techniques that can account for the nuances and variations in expressing clinical information. Developing evaluation methods that go beyond surface-level similarities and consider the underlying medical context is crucial for ensuring the reliability and usefulness of LLMs in generating clinical reports.
4.4 Medical Robotics
Medical robotics are revolutionizing healthcare, offering precision in ultrasound scanning, diagnostic analyses, and multi-agent surgical planning xu2024enhancing; killeen2024take; moghani2024sufia. These intelligent systems enhance the medical workforce, bridge gaps in staffing, and perform tasks that extend human physical limits. Incorporating LLMs, robots interpret environmental data to navigate and execute complex tasks.
Incorporating LLMs, robots interpret environmental data to navigate and execute complex tasks. The Graph-based Robotic Instruction Decomposer ni2023grid exemplifies this by utilizing natural language for route planning. This method has been shown to outperform GPT-4 by over 25.4% accuracy in simultaneously predicting the correct action and object, as well as 43.6% accuracy in correctly predicting instruction tasks ni2023grid.
Moreover, robotics in surgery now involve multi-agent planning systems wang2024large. Such systems involve a symphony of robotic units working collaboratively, each programmed to perform specific tasks that complement one another. For example, one robotic arm may be responsible for precise incisions while another manages real-time imaging, providing surgeons with a dynamic view of the operation. This orchestrated approach enhances surgical accuracy, reduces procedure times, and minimizes patient recovery periods. In diagnostics, robots leverage the power of data analytics to sift through extensive EHRs, identifying patterns and anomalies that may elude the human eye pee2019artificial.
Using LLMs in medical robotics can also improve human-computer interactions. Robots with improved interactivity may recognize human emotions and requests through natural language inputs. This makes patient communication with robots less intimidating and more user-friendly than existing implementationsqiu2023large; d2022emotion.
Guideline
SuFIA moghani2024sufia incorporates the strong reasoning capabilities of LLMs with perception modules to implement high-level planning and low-level control of a robot for surgical sub-task execution, with the current best results obtained via API calls to GPT-4 Turbo. While measures are incorporated to improve safety and reliability, deploying autonomous or semi-autonomous robotic surgical assistants (RSAs) in real-world scenarios still has the potential risks from unexpected circumstances. Another work UltrasoundGPT xu2024enhancing proposes an ultrasound-bodied intelligence system equipping ultrasound robots with LLMs and domain knowledge xu2024enhancing. It designs an ultrasound operation knowledge database for the LLM to enable precise motion planning. A dynamic scanning strategy based on prompt engineering allows LLMs to adjust motion planning during procedures. Their system improves ultrasound scan efficiency and quality from verbal commands, contributing to non-invasive diagnostics and streamlined workflows. Another work killeen2024take attempts to interpret domain-specific language in X-ray-guided surgery. It consists of a minimal protocol enabling an LLM, i.e., GPT-4, to control a robotic X-ray system, namely the Brainlab Loop-X device.
Discussion
Some challenges with implementing medical robotics are similar to those with collaborative robots (cobots), which are designed to work alongside and interact with human workers in shared spacesemaminejad2023trust. Integrating LLMs into medical robotics algorithms for route planning and motion control poses a critical challenge due to the risk of errors and biases inherent in LLMs. The complex and dynamic nature of shared human-robot workspaces may lead to LLM-powered cobots misjudging human intentions or making inappropriate decisions, posing safety risks. In contrast, traditional industrial robots, designed to operate at high speeds in controlled, human-free environments, could better withstand and contain errors or malfunctions due to their rigid and powerful designs, effectively mitigating human injury risks. Future research opportunities could explore safety features for cobots, such as sophisticated sensing technologies and physical design constraints, which aim to minimize the occurrence and consequences of judgment errors related to LLMs in shared human-robot environments weerarathna2023human; moglia2021systematic; xia2023nested.
4.5 Medical Language Translation
There are two main areas of medical language translation; the translation of medical terminology from one language to another noll2023machine and the translation of medical dialogue for ease of interpretation by non-professional personnel karabacak2023advent. Both areas are important for seamless communication between different groups. The language barrier to global collaboration in both research and medical techniques can be largely reduced with the help of LLMs. Language translation also improves accuracy in education resources and research articles, making knowledge accessible worldwide ahn2023impending.
Guideline
Medical mT5garcia2024medical, Apollowang2024apollo, and BiMediXpieri2024bimedix are multilingual large language models in the medical domain. Medical mT5garcia2024medical, which is based on multilingual T5 (mT5) with 738 million / 3 billion parameters, is trained on 4.5 billion tokens consisting of various languages languages, i.e., English, French, Italian, and Spanish. Apollowang2024apollo supports English, Chinese, French, Spanish, Arabic, and Hindi based on the Qwen model at various relatively small sizes (i.e., 0.5B, 1.8B, 2B, 6B, and 7B), achieving the best performance among models of equivalent size. BiMediXpieri2024bimedix is a bilingual medical mixture-of-experts language model for English and Arabic, proposing a semi-automated English-to-Arabic translation pipeline with human refinement for high-quality translations. For medical translation to lay language, a work aims to enhance the performance of language models in biomedical abstractive summarisation by aggregating knowledge from external papers cited within the source articletang2023improving. It proposes a novel attention-based citation aggregation model that integrates domain-specific knowledge from citation papers, allowing neural networks to generate summaries by leveraging both the paper content and relevant knowledge from citation paperstang2023improving. Another work introduces Retrieval-Augmented Lay Language (RALL) generation with a large and broad-ranging 63k lay language generation pairs from 12 journals, intuitively fitting the need for external knowledge beyond expert-authored source documentsguo2024retrieval. It also evaluates the ability of both an open-source LLaMA-2 and closed-source GPT-4 in background explanation, with and without retrieval augmentation.
Discussion
In both translation and simplification tasks, misinterpretation is a common occurrence that can have damaging consequences. In developing and deploying medical translation and simplification platforms, developers should prioritize professional datasets, such as textbooks and peer-reviewed journals for medical knowledge recall. This way, it will be less likely for misinformation from unreliable web sources to skew the output chen2023can. Another ethical consideration of using LLMs to perform medical translation is the potential for discriminatory verbiage to be inserted inadvertently into the output. Such verbiage is difficult to prevent due to the nature of the pipeline. This may cause miscommunications and even have legal consequences. karabacak2023advent.
4.6 Medical Education
LLMs can be incorporated into the medical education system in different ways, including facilitating study through explanations, aiding in language translation, answering questions, assisting with medical exam preparation, and providing Socratic-style tutoring qiu2023large; biri2023assessing. Therefore, medical education could involve text generation, text simplification, semantic textual similarity, information retrieval, and etc. It has been suggested that medical education can be augmented by generating scenarios, problems, and corresponding answers by an LLM. Students will gain a richer educational experience through personalized study modules and case-based assessments, encountering a wider array of challenges and scenarios beyond those found in standard textbooks karabacak2023advent. LLMs can also generate feedback on student responses to practical problems, allowing students to know their areas of weakness in real time. Inherently, these will better prepare these medical students for the real world since they would have been exposed to more scenariosahn2023impending.
Another use of LLMs in the medical field is educating the public. Medical dialogues are often complex and difficult to understand for the average patient. LLMs can tune the textual output of prompts to use varying degrees of medical terminology for different audiences. This will make medical information easy to understand for the average person while ensuring medical professionals have access to the most precise information karabacak2023advent.
Guideline
Large language models such as ChatGPTopenai2023chatgpt, and Med-Geminisaab2024capabilities are increasingly demonstrating their potential in medical education applications. These powerful models, trained on vast amounts of medical data, offer capabilities in knowledge synthesis, question answering, and content generation that can augment traditional teaching methods. For instance, ChatGPTopenai2023chatgpt can provide explanations and clarifications on complex medical concepts, facilitating self-study and reinforcing understanding. Med-Geminisaab2024capabilities, a multimodal model, can analyze medical images and generate detailed reports, aiding in the training of diagnostic skills. Institutions are exploring the integration of these language models into curricula, leveraging their strengths while ensuring proper oversight and ethical considerations. As this technology continues to advance, it holds promise for enhancing the efficiency and accessibility of medical education while complementing human expertise.
Discussion
Potential downsides of using LLMs in medical education include the current lack of ethical training and biases in training datasetsabd2023large. These biases, if not addressed, can propagate through the generated outputs, reinforcing stereotypes and potentially leading to discrimination in medical education. The lack of explicit ethical training during LLM development may also result in the generation of content that does not align with the ethical principles and guidelines of the medical profession, such as promoting unethical practices or violating patient privacy.
Furthermore, the risk of misinformation, particularly in the form of hallucinations, presents a challenge in utilizing LLMs for medical education. LLMs can generate plausible-sounding but factually incorrect information, which can mislead students and healthcare professionals if relied upon without proper verification. This can lead to the propagation of misconceptions, inappropriate treatment strategies, or misdiagnosis rawte2023survey. To mitigate these risks, it is essential to establish rigorous fact-checking and validation processes and emphasize the importance of critical thinking, evidence-based practice, and the verification of information from multiple reliable sources in medical education.
4.7 Mental Health Support
Mental health support involves both diagnosis and treatment. For example, depression is treated through a variety of psychotherapies, including cognitive behavior therapy, interpersonal psychotherapy, psychodynamic therapy, etc. vaidyam2019chatbots. Many of these techniques are primarily dominated by patient-doctor conversations, with lengthy treatment plans that are cost-prohibitive for many. The ability of LLMs to serve as conversation partners and companions may lower the barrier to entry for patients with financial or physical constraints stock2023tell, increasing the accessibility to mental health treatments liu2023chatcounselor. There have been various research works and discussions on the effects of incorporating LLMs into the treatment plan liu2023chatcounselor; de2023benefits; hua2024large.
The level of self-disclosure has a heavy impact on the effectiveness of mental health diagnosis and treatment. The degree of willingness to share has a direct impact on the diagnosis results and treatment plan. Studies have shown that patient willingness to discuss mental health-related topics with a robot is high robinson2023brief; de2023benefits. Alongside the convenience and lower financial stakes, mental health support by LLMs has the potential to be more effective than human counterparts in many scenarios.
Guideline
Development and deployment of LLMs targeted at mental health support can start with an existing LLM. Instead of pre-training or fine-tuning on general medical data, it is often better to use medical question and answer data as most of the LLM’s work will be talking to the patient, which involves back-and-forth conversation in the format of question and answering lai2023psy. PsyChat qiu2023psychat is a client-centric LLM dialogue system that provides psychological support comprising five modules: client behavior recognition, counselor strategy selection, input packer, response generator, and response selection. Specifically, the response generator is fine-tuned with ChatGLM-6B with a vast dialogue dataset. Through both automatic and human evaluations, the system has demonstrated its effectiveness and practicality in real-life mental health support scenarios. ChatCounselor is designed to provide mental health support. It initializes from Vicuna and fine-tunes from an 8k size instruct-tuning dataset collected from real-world counseling dialogue examples liu2023chatcounselor. Psy-LLM is an LLM aimed to be an assistive mental health tool to support the workflow of professional counselors, particularly to support those who might be suffering from depression or anxiety lai2023psy. Another work presents a comprehensive evaluation of prompt engineering, few-shot, and fine-tuning techniques on multiple LLMs in the mental health domainxu2024mental. The results reveal that fine-tuning on a variety of datasets can improve LLM’s capability on multiple mental-health-specific tasks across different datasets simultaneouslyxu2024mental. The work also releases their model Mental-Alpaca and Mental-FLAN-T5 as open-source LLMs targeted at multiple mental health prediction tasksxu2024mental.
Discussion
Two of the most critical difficulties in employing LLMs for mental health support are the lack of emotional understanding and the risk of inappropriate or harmful responses ma2023understanding. LLMs, being language models, may struggle to fully grasp and respond to the complex emotional states and needs of individuals seeking mental health support. They may not be able to provide the same level of empathy and human connection that is crucial in therapeutic interactions.
Moreover, if not properly trained or controlled, LLMs may generate responses that are inappropriate, insensitive, or even harmful to individuals in vulnerable emotional states chung2023challenges. They may provide advice that is not grounded in evidence-based psychological practices or that goes against established mental health guidelines. Addressing these challenges requires rigorous training of LLMs in evidence-based practices, ethical considerations, and risk assessment protocols, as well as collaboration between mental health professionals and AI researchers.
4.8 Medical Inquiry and Response
The rapid advancement of LLMs also opens up new possibilities for improving healthcare delivery and patient care. LLMs, trained on vast amounts of medical knowledge, have the potential to understand and generate human-like text, making them suitable for tasks such as answering patient inquiries and assisting physicians in documentation ren2024healthcare; wang2024jmlr. As the demand for accessible and efficient healthcare services grows, researchers are exploring the use of medical LLMs to alleviate the burden on healthcare professionals and provide patients with reliable information and support. Therefore, medical inquiry and response could involve entity extraction, information retrieval, question answering, and text generation/summarization.
Guideline
Several pioneering systems demonstrate the feasibility and effectiveness of this approach. Healthcare Copilotren2024healthcare integrates dialogue, memory, and processing components to enable safe patient-LLM interactions, enhance conversations with historical data, and summarize consultations. Google’s Articulate Medical Intelligence Explore (AMIE)tu2024towards engages in diagnostic conversations with patients, exhibiting reasoning capabilities comparable to human doctors. AMIE employs a novel self-play-based simulated environment with automated feedback mechanisms to scale learning across diverse medical contexts. Another LLM-based diagnostic systemsun2024conversational enhances planning capabilities by emulating doctors, utilizing reinforcement learning for disease screening and initial diagnoses, and leveraging LLMs to parse guidelines and conduct differential diagnoses. These systems showcase the potential of medical LLMs in providing high-quality, AI-powered medical consultations and assisting physicians in their daily practice.
Discussion
However, there is still far from deploying them in the real-world healthcare system. Several challenges must be addressed before widespread deployment in real-world healthcare settings. One major concern is the potential for biased or inaccurate outputs, which could lead to improper medical advice or misdiagnosisrawte2023survey. Rigorous testing and validation across diverse patient populations and medical contexts are essential to ensure the reliability and generalizability of these systems. Additionally, the integration of medical LLMs into existing healthcare workflows and infrastructure may require substantial technical and organizational efforts. Privacy and security concerns surrounding patient data must also be carefully considered and addressed.
Furthermore, the development and deployment of medical LLMs raise important ethical and responsible AI considerations. Ensuring transparency, explainability, and accountability in the decision-making processes of these systems is crucial to maintaining trust and facilitating informed consent from patients howthefuckisthisonNature; shen2023anything. The potential impact on the doctor-patient relationship and the role of human physicians in an AI-assisted healthcare setting must also be carefully examined. Ongoing collaboration between AI researchers, healthcare professionals, ethicists, and policymakers will be necessary to establish guidelines and best practices for the responsible development and deployment of medical LLMs in real-world healthcare settings.
5 Challenges
We address the challenges and discuss solutions to the adoption of LLMs in an array of medical applications.
5.1 Hallucination
Hallucination of LLMs refers to the phenomenon where the generated output contains inaccurate or nonfactual information. It can be categorized into intrinsic and extrinsic hallucinations umapathi2023med; rawte2023survey. Intrinsic hallucination generates outputs logically contradicting factual information, such as wrong calculations of mathematical formulas rawte2023survey. Extrinsic hallucination happens when the generated output cannot be verified, typical examples include LLMs ‘faking’ citations that do not exist or ‘dodging’ the question. When integrating LLMs into the medical domain, fluent but nonfactual LLM hallucinations can lead to the dissemination of incorrect medical information, causing misdiagnoses, inappropriate treatments, and harmful patient education. It is therefore vital to ensure the accuracy of LLM outputs in the medical domain.
Potential Solutions
Current solutions to mitigate LLM hallucination can be categorized into training-time correction, generation-time correction, and retrieval-augmented correction. The first (i.e. training-time correction) adjusts model parameter weights, thus reducing the probability of generating hallucinated outputs. Its examples include factually consistent reinforcement learning factualRL and contrastive learningcontrastivelearning-factual. The second (i.e. generation-time correction) adds a ‘reasoning’ process to the LLM inference to ensure reliability, using drawing multiple samples manakul2023selfcheckgpt or a confidence score to identify hallucination before the final generation. The third approach (i.e. retrieval-augmented correction) utilizes external resources to mitigate hallucination, for example, using factual documents as prompts factualdocumenthallucination or chain-of-retrieval prompting technique chain-of-verification.
5.2 Lack of Evaluation Benchmarks and Metrics
Current benchmarks and metrics often fail to evaluate LLM’s overall capabilities, especially in the medical domain. For example, MedQA (USMLE) medqa and MedMCQA Medmcqa offer extensive coverage on QA tasks but fail to evaluate important LLM-specific metrics, including trustworthiness, helpfulness, explainability, and faithfulness xie2023faithful. It is therefore imperative to develop domain and LLM-specific benchmarks and metrics.
Potential Solutions
Singhal et al. medpalm proposed HealthSearchQA consisting of commonly searched health queries, offering a more human-aligned benchmark for evaluating LLM’s capabilities in the medical domain. Benchmarks such as TruthfulQA truthfulqa and HaluEval halueval evaluate more LLM-specific metrics, such as truthfulness, but do not cover the medical domain. Future research is necessary to meet the need for more medical and LLM-specific benchmarks and metrics than what is currently available.
5.3 Domain Data Limitations
Current datasets in the medical domain (Table 2) remain relatively small compared to datasets for training general-purpose LLMs (Table 1). These limited small datasets only cover a small space medpalm of the vase domain of medical knowledge. This results in LLMs exhibiting extraordinary performance on open benchmarks with extensive data coverage, yet falling short on real-life tasks such as differential diagnosis and personalized treatment planning medpalm2.
Although the volume of medical and health data is large, most require extensive ethical, legal, and privacy procedures to be accessed. In addition, these data are often unlabeled, and solutions to leverage these data, such as human labeling and unsupervised learning liu2021KGAE, face challenges due to the lack of human expert resources and small margins of error.
Potential Solutions
Current state-of-the-art approaches medpalm2; li2023chatdoctor typically fine-tune the LLMs on smaller open-sourced datasets to improve their domain-specific performance. Another solution is to generate high-quality synthetic datasets using LLMs to broaden the knowledge coverage; however, it has been discovered that training on generated datasets causes models to forget shumailov2023model. Future research is needed to validate the effectiveness of using synthetic data for LLMs in the medical field.
5.4 New Knowledge Adaptation
LLMs are trained on extensive data to learn knowledge. Once trained, it is expensive and inefficient to inject new knowledge into an LLM through re-training. However, it is sometimes necessary to update the knowledge of the LLM, for example, on a new adverse effect of a medication or a novel disease. Two problems occur during such knowledge updates. The first problem is how to make LLMs appropriately ‘forget’ the old knowledge, as it is almost impossible to remove all ‘old knowledge’ from the training data, and the discrepancy between new and old knowledge can cause unintended association and bias hoelscher2023detecting. The second problem is the timeliness of the additional knowledge - how do we ensure the model is updated in real-time liu2023medical? Both problems pose substantial barriers to using LLMs in medical fields, where accurate and timely updates of medical knowledge are crucial in real-world implementations.
Potential Solutions
Current solutions to knowledge adaptation can be categorized into model editing and retrieval-augmented generation. Model editing modelediting alters the knowledge of the model by modifying its parameters. However, this method does not generalize well, with their effectiveness varying across different model architectures. In contrast, retrieval-augmented generation provides external knowledge sources as prompts during model inference; for example, Lewis et al. RAGnewknowledge enabled model knowledge updates by updating the model’s external knowledge memory.
5.5 Behavior Alignment
Behavior alignment refers to the process of ensuring that the LLM’s behaviors align with the objectives of its task. Development efforts have been spent on aligning LLMs with general human behavior, but the behavior discrepancy between general humans and medical professionals remains challenging for adopting LLMs in the medical domain. For example, ChatGPT is well aligned with general human behavior, but their answers to medical consultations are not as concise and professional as those by human experts instructgpt. In addition, misalignment in the medical domain introduces unnecessary harm and ethical concerns alignmentAI that lead to undesirable consequences.
Potential Solutions
Current solutions include instruction fine-tuning, reinforcement learning from human feedback (RLHF) instructgpt, and prompt tuning lester2021power; liu2021p. Instruction fine-tuning zhang2023instruction refers to improving the performance of LLMs on specific tasks based on explicit instructions. For example, Ouyang et al. instructgpt used it to help LLMs generate less toxic and more suitable outputs. RLHF uses human feedback to evaluate and align the outputs of LLMs. It is effective in multiple tasks, including becoming helpful chatbots rlhfdialogue and decision-making agents webgpt. Prompt tuning can also align LLMs to the expected output format. For example, Liu et al. chainofhindsight uses a prompting strategy, chain of hindsight, to enable the model to detect and correct its errors, thus aligning the generated output with human expectations.
5.6 Ethical and Safety Concerns
Concerns have been raised regarding using LLMs (e.g. ChatGPT) in the medical domain chatgptconcernmed, with a focus on ethics, accountability, and safety. For example, the scientific community has disapproved of using ChatGPT in writing biomedical research papers howthefuckisthisonNature due to ethical concerns. The accountability of using LLMs as assistants to practice medicine is challenging he2023survey; tian2024opportunities. Li et al.chatgptjailbreak and Shen et al.shen2023anything found that prompt injection can cause the LLM to leak personally identifiable information (PII), e.g. email addresses, from its training data, which is a substantial vulnerability when implementing LLM in the medical domain.
Potential Solutions
With no immediate solutions available, we have nevertheless observed research efforts to understand the cause of these ethical and legal concerns. For example, Wei et al. understandsafety propose that PII leakage is attributed to the mismatched generalization between safety and capability objectives (i.e., the pre-training of LLMs utilizes a larger and more varied dataset compared to the dataset used for safety training, resulting in many of the model’s capabilities are not covered by safety training).
5.7 Regulatory Challenges
The regulatory landscape of LLMs presents distinct challenges due to their large scale, broad applicability and varying reliability across applications. As LLMs progressively permeate the fields of medicine and healthcare, their versatility allows a single LLM family to facilitate a multitude of tasks across a broad spectrum of interest groups. This represents a substantial departure from the AI-based medical technologies of the past, which were typically tailored to meet specific medical needs and cater to particular interest groups newregulatoryLLM; AImed. In addition, the recent innovations of AI-enabled personalized approaches in areas such as oncology also present challenges to the traditional one-for-all auditing process newregulatoryoncology. This divergence and innovation necessitate regulators to develop adaptable, foresightful frameworks to ensure the safety, ethical standards, and privacy of the new family of LLMs-powered medical technologies.
Potential Solutions
To address the complex regulatory challenges without hindering innovation, regulators should devise adaptive, flexible, and robust frameworks. Drawing on the insights from Mesko and Topol newregulatoryLLM, creating a dedicated regulatory category and implementing patient design to enhance decision-making for LLMs used for medical purposes can better address their unique attributes and minimize harm. Furthermore, the insights outlined by Derraz et al. newregulatoryoncology emphasize the importance of implementing agile regulatory frameworks that can keep pace with the fast-paced advancements in personalized applications. Researchers both inside newregulatoryLLM; newregulatoryoncology and outside of healthcare regulatingLLM; auditingLLMs have proposed innovative strategies to regulate the use of LLMs involving (i) assessing LLMs-enabled applications in real-world settings, (ii) obligations of transparency of data and algorithms, (iii) adaptive risk assessment and mitigation processes, (iv) continuous testing and refinement of audited technologies. Such proactive regulatory adaptations are crucial to maintaining high standards of safety, ethics, and trustworthiness of medical technology.
Figure 6: Future directions of LLMs in clinical medicine in terms of both development and deployment.
6 Future Directions
Although LLMs have already made an impact on people’s lives through chatbots and search engines, their integration into medicine is still in the infant stage. As shown in Figure 6, numerous new avenues of medical LLMs await researchers and practitioners to explore how to better serve the general public and patients.
6.1 Introduction of New Benchmarks
Recent studies have underscored the shortcomings of existing benchmarks in evaluating LLMs for clinical applications chen2023extensive; chen2023large. Traditional benchmarks, which primarily gauge accuracy in medical question-answering, inadequately capture the full spectrum of clinical skills necessary for LLMs medpalm. Criticisms have been leveled against the use of human-centric standardized medical exams for LLM evaluation, arguing that passing these tests does not necessarily reflect an LLM’s proficiency in the nuanced expertise required in real-world clinical settings medpalm. In response, there is an emerging consensus on the need for more comprehensive benchmarks. These should include capabilities like sourcing from authoritative medical references, adapting to the evolving landscape of medical knowledge, and clearly communicating uncertainties thirunavukarasu2023large; medpalm. To further enhance the relevance of these benchmarks, new benchmarks should incorporate scenarios that test an LLM’s ability through simulation of real-world applications and adjust to feedback from clinicians while maintaining robustness. Additionally, considering the sensitive nature of healthcare, these benchmarks should also assess factors such as fairness, ethics, and equity, which, though crucial, pose quantification challenges medpalm. While efforts such as the AMIE study have advanced benchmarking by utilizing real physician evaluations and comprehensive criteria rooted in actual clinical skills and communication, as reflected in the Objective Structured Clinical Examination (OSCE), there remains a pressing need for benchmarks that are adaptive, scalable and robust for other diverse and personalized applications of LLMs. The aim is to create benchmarks that more effectively mirror diverse real-world clinical scenarios, thus providing a more accurate measure of LLMs’ suitability for their applications in medicine. Future research may focus on (i) using synthetic data along with real-world data to create benchmarks that are both comprehensive and scalable, (ii) using clinical guidelines and criteria to reflect real-world values that are not normally included in traditional benchmarks, (iii) physician-in-the-loop benchmarks to evaluate the performance of LLMs leveraging their human counterparts or users.
6.2 Multimodal LLM Integrated with Time-Series, Visual, and Audio Data
Multimodal LLMs (MLLMs), or Large Multimodal Models (LMMs), are LLM-based models designed to perform multimodal (e.g. involving both visual and textual) tasks yin2023survey. While LLMs primarily address NLP tasks, MLLMs support a broader range of tasks, such as comprehending the underlying meaning of a meme and generating website codes from images. This versatility suggests promising applications of MLLMs in medicine. Several MLLM-based frameworks integrating vision and language, e.g. MedPaLM M tu2023towards, LLaVA-Med li2023llavamed, Visual Med-Alpaca Visual-Med-Alpaca, Med-Flamingo moor2023medflamingo, and Qilin-Med-VL liu2023qilin, have been proposed to adopt the medical image-text pairs for fine-tuning, thus enabling the medical LLMs to efficiently understand the input medical (e.g. radiology) images. A recent study huang2023chatgpt proposes to integrate vision, audio, and language inputs for automated diagnosis in dentistry. However, there exist only very few medical LLMs that can process time series data, such as electrocardiograms (ECGs) li2023frozen and sphygmomanometers (PPGs) englhardt2023exploring, despite such data being important for medical diagnosis and monitoring. Although early in their proposed research stages, these studies suggest that MLLMs trained at scale have the potential to effectively generalize across various domains and modalities outside of NLP tasks. However, the training of MLLMs at scale is still costly and ineffective, resulting in the size of MLLMs being much smaller than LLMs. Moving forward, future research may focus on (i) more effective processing, representation, and learning of multi-modal data and knowledge, (ii) cost-effective training of MLLMs, especially modalities that are more resource-demanding such as videos and images, (iii) collecting or accessing safely, currently unavailable, multi-modal data in medicine and healthcare.
6.3 Medical Agents
LLM-based agents xi2023rise; wang2023survey utilize LLMs as controllers to leverage their reasoning capabilities. By integrating LLMs with external tools and multimodal perceptions, these agents can interact with environments, learn from feedback, and acquire new skills, enabling them to solve complex tasks (e.g., software design, molecular dynamics simulation) through human-like behaviors, such as role-playing and communication li2023camel; tang2023medagents.
However, integrating these agents effectively within the medical domain remains a challenge. The medical field involves numerous roles tang2023medagents and decision-making processes, especially in disease diagnosis that often requires a series of investigations involving CT scans, ultrasounds, electrocardiograms, and blood tests. The idea of utilizing LLMs to model each of these roles, thereby creating collaborative medical agents, presents a promising direction. These agents could mimic the roles of radiologists, cardiologists, pathologists, etc., each specializing in interpreting specific types of medical data. For example, a radiologist agent could analyze CT scans, while a pathologist agent could focus on blood test results. The collaboration among these specialized agents could lead to a more holistic and accurate diagnosis. By leveraging the comprehensive knowledge base and contextual understanding capabilities of LLMs, these agents not only interpret individual medical reports but also integrate these interpretations to form a cohesive medical opinion. To enhance the integration of LLMs-based agents, future research may explore (i) a seamless data pipeline that collects data from various devices and transforms them into data format compatible with LLMs (ii) effective communication and collaboration between agents, especially in areas such as ensuring truthfulness during communication, dispute resolution between agents, and role-based data security measures, (iii) real-time decision-making such as making timely decisions using data collected from remote monitoring devices, (iv) adaptive learning such as preparing for a new pandemic or learning from unseen medical conditions.
6.4 LLMs in Underrepresented Specialties
Current LLM research in medicine has largely focused on general medicine, likely due to the greater availability of data in this area medpalm2; tian2024opportunities. This has resulted in the under-representation of LLM applications in specialized fields like ‘rehabilitation therapy’ or ‘sports medicine’. The latter, in particular, holds potential, given the global health challenges posed by physical inactivity. The World Health Organization identifies physical inactivity as a major risk factor for non-communicable diseases (NCDs), impacting over a quarter of the global adult population who2022physical. Despite initiatives to incorporate physical activity (PA) into healthcare systems, implementation remains challenging, particularly in developing countries with limited PA education among healthcare providers who2022physical. LLMs could play a pivotal role in these settings by disseminating accurate PA knowledge and aiding in the creation of personalized PA programs connor2023large. Such applications could enhance PA levels, improving global health outcomes, especially in resource-constrained environments. To spark innovation in these underrepresented specialties, future research can focus on areas such as (i) effective data collection in underrepresented specialties, (ii) applications of LLMs in assisting with tasks of underrepresented specialties, (iii) using LLMs to help progress the research of these underrepresented specialties.
6.5 Interdisciplinary Collaborations
Just as interdisciplinary collaborations are crucial in safety-critical areas like nuclear energy production, collaborations between the medical and technology communities for developing medical LLMs are essential to ensure AI safety and efficacy in medicine. The medical community has primarily adopted LLMs provided by technology companies without rigorously questioning their data training, ethical protocols, or privacy protection. Medical professionals are therefore encouraged to actively participate in creating and deploying medical LLMs by providing relevant training data, defining the desired benefits of LLMs, and conducting tests in real-world scenarios to evaluate these benefits thirunavukarasu2023large; omiye2024large; clusmann2023future. Such assessments would help to determine the legal and medical risks associated with LLM use in medicine and inform strategies to mitigate LLM hallucination mello2023chatgpt. Additionally, training ‘bilingual’ professionals—those versed in both medicine and LLM technology—is increasingly vital due to the rapid integration of LLMs in healthcare. Future research may explore (i) interdisciplinary frameworks, such as frameworks to facilitate the sharing of localized data from rural clinics, (ii) ‘bilingual education programs’ that offer training from both worlds - AI and medicine, (iii) effective in-house development methods to help hospitals and physicians ‘guard’ patient data from corporations while still being able to embrace innovation.
Acknowledgements
This work was supported in part by the Pandemic Sciences Institute at the University of Oxford; the National Institute for Health Research (NIHR) Oxford Biomedical Research Centre (BRC); an NIHR Research Professorship; a Royal Academy of Engineering Research Chair; the Well-come Trust funded VITAL project; the UK Research and Innovation (UKRI); the Engineering and Physical Sciences Research Council (EPSRC); and the InnoHK Hong Kong Centre for Cerebro-cardiovascular Engineering (COCHE).
Author Contributions
FL, ZL, JL, and DC conceived the project. FL conceived and designed the study. HZ, FL, BG, XZ, and JH conducted the literature review, performed data analysis, and drafted the manuscript. All authors contributed to the interpretation and final manuscript preparation. All authors read and approved the final manuscript.
Competing Interests
The authors declare no competing interests.
Appendix A Appendix: Discriminative Tasks
Table 4: Summary of existing widely-used biomedical NLP datasets for evaluating medical LLMs.
Types Tasks Datasets Data Scale Data Source
Discriminative Tasks (Sec. A) Question Answering (Sec. A.1) MedQA (USMLE) [medqa] 61,097 multiple-choice QA pairs 51 medical textbooks
MedMCQA [Medmcqa] 194k multiple-choice QA pairs AIIMS and NEET PG entrance exam
MMLU [hendrycks2020measuring] 300 multiple-choice QA pairs Clinical Knowledge
PubMedQA [PubmedQA] 273,500 question-context-answer triples PubMed [PubMed, sen2008collective]
BioASQ-QA [krithara2023bioasq] 4,721 QA pairs BioASQ [tsatsaronis2015overview]
EMRQA [pampari2018emrqa] 400k+ QA pairs i2b2
CliCR [vsuster2018clicr] 105k QA pairs 12k clinical case reports
COVID-QA [moller2020covid] 2,019 QA pairs 147 articles (CORD-19 [wang2020cord])
MASH-QA [zhu2020question] 34,808 QA pairs WebMD website
Entity Extraction (Sec. A.2) NCBI Disease [dougan2014ncbi] 6,892 entity mentions 793 PubMed abstracts
BC5CDR [li2016biocreative] 28,785 entity mentions 1,500 PubMed articles
BC4CHEMD [krallinger2015chemdner] 84,355 entity mentions 10,000 PubMed abstracts
BioRED [luo2022biored] 20,419 entity mentions 600 PubMed abstracts
CMeEE [zhang2021cblue] 24,850 entity mentions 938 files
ADE [gurulingappa2012development] 21,160 entity mentions 2,972 documents
2012 i2b2 [sun2013evaluating] 30,690 entity mentions 310 discharge summaries
2014 i2b2/UTHealth [stubbs2015annotating] 57,744 entity mentions 1,304 longitudinal medical records
2018 n2c2 [henry20202018] 83,869 entity mentions 505 discharge summaries[johnson2016mimic]
Cadec [karimi2015cadec] 9,111 entity mentions 1,253 posts
Relation Extraction (Sec. A.3) BC5CDR [li2016biocreative] 3,116 relations 1,500 PubMed articles
BioRED [luo2022biored] 6,503 relations 600 PubMed abstracts
ADE [gurulingappa2012development] 13,806 relations 2,972 documents
2018 n2c2 [henry20202018] 59,810 relations 505 discharge summaries
2010 i2b2/VA [uzuner20112010] 1,748 reports clinical reports
DDI [segura2013semeval] 5,021 relations 1,017 biomedical texts
EU-ADR [van2012eu] 2,436 relations 100 abstracts
Text Classification (Sec. A.4) ADE [gurulingappa2012development] 20,967 sentences 2,972 documents
2014 i2b2/UTHealth [stubbs2015annotating] 1,304 records longitudinal medical records
HoC [baker2016automatic] 5,310 sentences 1,499 PubMed abstracts
OHSUMED [hersh1994ohsumed] 50,216 abstracts 50,216 medical abstracts
WNUT-2020 Task 2 [nguyen2020wnut] 10,000 tweets 10,000 tweets
Medical Abstracts [schopf2022evaluating] 14,438 abstracts 28,880 medical abstracts
MIMIC-III [johnson2016mimic] 52,724 discharge summaries 53,423 hospital admissions
Natural Language MedNLI [romanov2018lessons] 14,049 sentence pairs MIMIC-III [johnson2016mimic]
Inference (Sec. A.5) BioNLI [bastan2022bionli] 40,459 sentence pairs PubMed abstracts
Semantic Textual Similarity (Sec. A.6) MedSTS [wang2020medsts] 174,629 sentence pairs 3M clinical notes
2019 n2c2/OHNLP [wang20202019] 2,054 sentence pairs 113k clinical notes and MedSTS [wang2020medsts]
BIOSSES [souganciouglu2017biosses] 100 sentence pairs Text Analysis Conference (TAC)
Information Retrieval (Sec. A.7) TREC-COVID [voorhees2021trec] 8,691 3-level query-document relations CORD-19 [wang2020cord]
NFCorpus [boteva2016full] 712k 3-level query-document relations NutritionFacts.org (NF) website
BioASQ (BEIR) [thakur2021beir] 32,916 binary query-document relations BioASQ [tsatsaronis2015overview]
Generative Tasks (Sec. B) Text Summarization /Generation (Sec. B.1/Sec. B.3) MIMIC-CXR [johnson2019mimic] 128k Findings-Impression pairs radiology reports
PubMed [PubMed, sen2008collective] 36M+ citations and abstracts biomedical literature
PMC [PMC] 9,407,149 papers biomedical and life sciences literature
CORD-19 [wang2020cord] 140k+ papers COVID-19 literature
MentSum [sotudeh2022mentsum] 24,119 post-TLDR summary pairs 24,119 Reddit posts
MeQSum [abacha2019summarization] 1,000 question-summary pairs U.S. National Library of Medicine
Text Simplification MultiCochrane [joseph2023multilingual] 7,755 abstract-summary pairs Cochrane Library systematic reviews
(Sec. B.2) AutoMeTS [van2020automets] 3,300 sentence pairs English Wikipedia [kauchak2013improving]
Table 5: The performance (accuracy) on five question answering (QA) datasets, shown in percentage (%). FT: fine-tuned.
Types Models MedQA (USMLE) PubMedQA MedMCQA MMLU (Clinical Knowledge) MMLU (Professional Medicine)
Task-specific FT Models BERT [devlin2018bert] 44.6 51.6 43.0 - -
RoBERTa [yang2020measurement, van2020automets] 43.3 52.8 - - -
BioBERT [lee2020biobert, raj2019saama] 30.1 60.2 - - -
SciBERT [beltagy2019scibert] 29.5 57.4 - - -
ClinicalBERT [alsentzer2019publicly] 29.1 49.1 - - -
BlueBERT [peng2019transfer, mutinda2020detecting, mahajan2020identification] - 48.4 - - -
PubMedBERT [gu2021domain] 38.1 55.8 - - -
General LLMs Med-PaLM-2 [medpalm2] 86.5 81.8 72.3 88.3 95.2
FLAN-PaLM (few-shot) [medpalm] 67.6 79.0 57.6 - -
GPT-4 (zero-shot) [gpt-4] 78.9 75.2 69.5 86.0 93.0
GPT-4 (few-shot) [gpt-4] 86.1 80.4 73.7 86.4 93.8
GPT-3.5 (zero-shot) [GPT-3] 50.8 71.6 50.1 69.8 70.2
GPT-3.5 (few-shot) [GPT-3] 60.2 78.2 62.7 68.7 69.8
Clinical Camel-13B (zero-shot) [toma2023clinical] 34.4 72.9 39.1 54.0 51.8
Clinical Camel-13B (few-shot) [toma2023clinical] 45.2 74.8 44.8 60.4 53.3
Clinical Camel-70B (zero-shot) [toma2023clinical] 53.4 74.3 47.0 69.8 71.3
Clinical Camel-70B (few-shot) [toma2023clinical] 60.7 77.9 54.2 72.8 75.0
Galactica [taylor2022galactica] 44.4 77.6 52.9 - -
BioMedLM [venigalla2022biomedlm] 50.3 74.4 - - -
BioGPT [luo2022biogpt] - 81.0 - - -
PMC-LLaMA [wu2023pmc] 56.4 77.9 56.0 - -
Human (expert) [he2023survey, diao2023lmflow] 87.0 78.0 90.0 - -
A.1 Question Answering
Task Description
Question Answering (QA) [jin2022biomedical] aims to give answers to the given queries, by generating multiple-choice or free-text responses. A multiple-choice response happens when the model is asked a question with the material for answering the question included. For example, the QA for ’Is hypertension a risk factor for cardiovascular disease, yes or no?’ is multiple-choice-orientated. In contrast, an open question without potential answers provided gives rise to a free-text response, an example QA being ’What are the common symptoms of influenza?’.
Datasets and Models
The commonly used biomedical QA datasets are MedQA (USMLE) [medqa], PubMedQA [PubmedQA], and MedMCQA [Medmcqa]. Such datasets are relatively small in size compared to general datasets, thus the model performance can be readily improved by pre-training on general data and then finetuning on the biomedical data [soni2020evaluation, yoon2019pre]. Pergola et al. [pergola2021boosting] proposed a biomedical-specific masking method: instead of masking tokens randomly, the model identifies biomedical-related tokens and masks them to focus on in-domain learning. Wang et al. [wang2023large] defined the self-questioning prompting (SQP) and utilize it on the BioASQ dataset. To extract useful information for specific tasks, SQP lets the GPT model ask questions about the given text and then answer these questions. However, hallucination remains a threat to the quality of outputs in QA. One solution is to use biomedical search systems such as Almanac [hiesinger2023almanac]. Another approach is to use extra datasets as an augmentation for the QA task. There are many chatbots based on LLM QA, such as Clinical Camel [toma2023clinical], DoctorGLM [xiong2023doctorglm], ChatDoctor [li2023chatdoctor], HuaTuo [wang2023huatuo], HuaTuoGPT [zhang2023huatuogpt], and MedAlpaca [han2023medalpaca]. Except for a few [li2023huatuo], most chatbots are a black box for consumers, and thus providing insight into them is a potential direction for further studies.
Evaluation
Existing works use accuracy as the metric for model performance in QA. We compare the performance of task-specific (fine-tuned) BERT variations [gu2021domain, liu2023kbmqa, medqa, alghanmi2023interpreting], Med-PaLM-2 [medpalm2, toma2023clinical], FLAN-PaLM [medpalm], GPT-4 [nori2023capabilities, toma2023clinical], GPT-3.5 [lievin2022can, toma2023clinical], Clinical Camel [toma2023clinical], Galactica [taylor2022galactica], BioMedLM [venigalla2022biomedlm], BioGPT [luo2022biogpt], and PMC-LLaMA [wu2023pmc] on the datasets described in Table 4. The citations in this paragraph correspond to the sources that provide data for assessing the model performance. Table 5 details their performances on five widely-used datasets, indicating that GPT-series models perform much better than BERT-series models. For GPT-series models, the observation that the few-shot setting outperforms the zero-shot setting still holds. Med-PaLM-2 has the highest accuracy among all models.  These results indicate that it is feasible to fine-tune general LLMs on medical data and improve their performance.
Table 6: The performance (F1-score) on four entity extraction datasets, , shown in percentage (%).
Types Models NCBI Disease BC5CDR Disease BC5CDR Drug/Chem.
Task-specific FT Models BERT [devlin2018bert] 85.63 82.41 91.16
BioBERT [lee2020biobert, raj2019saama] 89.36 86.56 93.44
SciBERT [beltagy2019scibert] 88.57 90.01 90.01
PubMedBERT [gu2021domain] 89.86 85.62 93.50
BioLinkBERT [yasunaga2022linkbert] 88.76 86.39 94.04
BlueBERT [peng2019transfer, mutinda2020detecting, mahajan2020identification] 88.04 83.69 91.19
General LLMs ClinicalBERT [alsentzer2019publicly] 86.32 83.04 90.80
GPT-3 [GPT-3] 51.40 43.60 73.00
GPT-3.5 (zero-shot) [GPT-3] 38.02 32.26 41.25
GPT-3.5 (few-shot) [GPT-3] 42.74 - 72.10
GPT-4 (zero-shot) [gpt-4] 57.85 - 81.90
GPT-4 (few-shot) [gpt-4] 58.39 - 82.43
ChatGPT [openai2023chatgpt] 50.49 51.77 60.30
BARD (zero-shot) [manyika2023overview] 96.00 - 97.70
BARD (few-shot) [manyika2023overview] 95.60 - 98.30
A.2 Entity Extraction
Task Definition
Entity extraction, or named entity recognition (NER) [li2020survey] aims to identify named entities mentioned in unstructured text, and classify them into predefined categories such as the names of persons, organizations, locations, medical codes, time expressions, quantities, monetary values, percentages, etc. Existing task-specific fine-tuned models leverage large-scale pre-trained language representations, and then fine-tune them on NER datasets to achieve state-of-the-art performances. The self-attention mechanisms [transformer] are used for efficiently capturing complex patterns and dependencies.
In the healthcare domain, medical NER aims to extract medical entities such as disease names, medication, dosage, and procedures from clinical narratives, electronic health records (EHRs), and chemicals and proteins from scientific literature. This enables NER to provide a solid basis for clinical decision support systems, automated patient monitoring, etc.
Datasets and Models
Table 4 shows commonly used biomedical NER datasets, such as NCBI Disease [dougan2014ncbi] and BC5CDR [li2016biocreative]. To perform the NER task, for each input token, the existing models output a dense representation, which not only embeds the tokens but also includes its relation with other tokens in the text. Therefore, with the dense representations of input tokens extracted as vectors, additional layers are applied to the last Transformer layer to fit the downstream entity extraction task. The widely-used additional layers are softmax, BiLSTM [Hochreiter1997LSTM], CRF [sun2021biomedical, yu2019biobert], and their combinations [portelli2021bert, sun2021biomedical, fraser2019extracting, yu2019biobert, chen2020using, kang2021umls]. Gu et al. [gu2023distilling] proposed a distillation method that uses few-shot GPT-3.5 to extract the correct entities and create the training set for the student model. This method shows that GPT models can be used to label the dataset first to achieve unsupervised learning. Wang et al. [wang2023large] defined a new prompting method self-questioning prompting (SQP) for NER. Wang et al. [wang2023large] further evaluated the performances of BARD, GPT-3.5, and GPT-4 with SQP and achieved some state-of-the-art results.
Evaluation
The entity-level F1 score is widely used to evaluate the models’ performance. Recent studies on the model performance of LLMs on biomedical NER tasks include GPT-3 [agrawal2022large, caufield2023structured, gutierrez2022thinking], GPT-3.5 [chen2023large], GPT-4 [chen2023large], and ChatGPT [chen2023extensive]. The citations in this paragraph correspond to the sources that provide data for assessing the model performance.
We summarize the performances of existing LLMs on the NER task in Table 6. It clearly shows that both the encoder-only BERT-series models (e.g. BioBERT [lee2020biobert], SciBERT [beltagy2019scibert], PubMedBERT [gu2021domain], BlueBERT [peng2019transfer], and ClinicalBERT [alsentzer2019publicly]) significantly outperform most decoder-only LLM GPT-series models (e.g., GPT-3 [GPT-3] and GPT-4 [gpt-4]), under both the zero-shot and few-shot settings. Recent work has investigated the performance of GPT-3 [agrawal2022large, caufield2023structured, gutierrez2022thinking], GPT-3.5 [chen2023large], and GPT-4 [chen2023large] on biomedical NER tasks. These GPT models still need further development to outperform traditional task-specific fine-tuned models [gutierrez2022thinking].
A.3 Relation Extraction
Task Description
Different from entity extraction, relation extraction (RE) aims to find the relation between entities in a text. It is highly related to entity extraction since identifying entities first can improve the model’s performance. A relation can be ordered or unordered. For example, in the sentence ’I work in London’ the relation ’workspace’ is order-sensitive and the pair (I, London) is of such a relation, but in the sentence ’Alice and Bob are friends’ the relation ’friend’ is not order-sensitive. In the biomedical area, relation extraction is often used to extract relations in a MedAbstrct text for further applications like text summarization.
Datasets and Models
For relation extraction, Table 4 shows representative biomedical datasets such as BC5CDR, BioRED, and DDI. In practice, most datasets indicate the potential entity for the relation instead of asking the model to find it, so RE is usually considered as a classification task. One common technique for task-specific fine-tuned models is to use the classification token (e.g., [CLS] token in Transformer [transformer]), which learns the information of all input tokens, and is further used to classify the relation. Some studies use only the classification token and apply softmax in the last layer [wei2019relation, thillaisundaram2019biomedical, liu2020document]. Others use not only the classification token, but also all other token representations. It has been shown that using all token representations outperforms using classification token only [su2020investigation]. Specifically, Wang et al. [wang2023large] defined the self-questioning prompting (SQP) as discussed in Sec A.1 and utilize it in the DDI RE task.
Evaluation
Existing studies on RE model performance typically use the F1-score, since the model performs a straightforward classification task. Table 7 compares the performance of BioBERT [lee2020biobert, wang2023pre], SciBERT [beltagy2019scibert, wang2023pre], PubMedBERT [gu2021domain, wang2023pre], BioGPT [luo2022biogpt], GPT-3 [agrawal2022large, caufield2023structured, gutierrez2022thinking], GPT-3.5 [chen2023large], GPT-4 [chen2023large], and ChatGPT [chen2023extensive] on the datasets introduced above. It clearly shows that at present task-specific fine-tuned models outperform GPT models. For GPT models, the F1-score of the best GPT model (GPT-4) is below 70%, while that of all the task-specific fine-tuned models based on BERT is approximately 80%. There is no appreciable difference in F1-score between zero-shot and few-shot [wang2023large].
Table 7: The performance (F1-score) on four relation extraction datasets, shown in percentage (%).
Types Models BC5CDR ChemProt DDI
Task-specific FT Models BioBERT [lee2020biobert, raj2019saama] - 76.46 80.88
SciBERT [beltagy2019scibert] 79.00 83.64 84.08
PubMedBERT [gu2021domain] - 77.24 82.36
BioGPT [luo2022biogpt] 46.17 - 40.76
BioLinkBERT [yasunaga2022linkbert] - 79.98 83.35
General LLMs GPT-3 [GPT-3] - 25.90 16.10
GPT-3.5 (zero-shot) [GPT-3] 43.29 57.43 33.49
GPT-3.5 (few-shot) [GPT-3] - 57.43 34.34
GPT-4 (zero-shot) [gpt-4] - 66.18 63.25
GPT-4 (few-shot) [gpt-4] - 66.82 65.58
ChatGPT [openai2023chatgpt] - 34.16 51.62
BARD (zero-shot) [manyika2023overview] - - 56.60
BARD (few-shot) [manyika2023overview] - - 77.20
Table 8: The performance (F1-score) on four text classification datasets, shown in percentage (%).
Types Models HoC i2b2 MIMIC-III OHSUMED
Task-specific FT Models PubMedBERT [gu2021domain] 82.34 - - -
BioGPT [luo2022biogpt] 85.12 - 82.30 -
General LLMs ChatGPT [openai2023chatgpt] 51.22 - - -
ChatGPT (zero-shot) [openai2023chatgpt] 65.72 92.90 - 39.93
ChatGPT (few-shot) [openai2023chatgpt] 69.32 92.90 60.00 47.05
GPT-4 (zero-shot) [gpt-4] 74.74 - - -
GPT-4 (few-shot) [gpt-4] 74.02 99.00 67.40 -
ChatGraph [shi2023chatgraph] - - - 60.79
Table 9: The performance (F1-score) of natural language inference, shown in percentage (%).
Types Models MedNLI
Task-specific FT Models BERT [wang2022paying] 76.11
ALBERT [wang2022paying] 77.84
Roberta [wang2022paying] 80.14
BioBERT [wang2022paying] 81.83
ClinicalBERT [wang2022paying] 80.66
BlueBERT [wang2022paying] 83.92
SciBERT [wang2022paying] 79.43
BlueBERT+AG [wang2022paying] 84.34
General LLMs GPT-3.5 [GPT-3] 82.21
GPT-3.5-Distillation [gu2023distilling] 80.24
GPT-4 [gpt-4] 85.69
BARD (zero-shot) [manyika2023overview] 76.00
BARD (few-shot) [manyika2023overview] 76.00
A.4 Text Classification
Task Description
Unlike entity extraction or relation extraction at the token level, text classification is a sentence or document-level task. It is a classic classification problem of assigning predefined labels to a text, where it is common for a text to have multiple labels. The task is to predict all correct labels given the input medical text, and can be used as the preprocessing of medical text simplification and summarization.
Datasets and Models
Table 4 shows commonly used datasets (e.g. HoC and OHSUMED) for text classification. Since this is a text-level task, the [CLS] vector is augmented when using task-specific fine-tuned (Transformer-based) models. Another approach for distilling the overall information is to use the weighted sum of final attention layer outputs. Recent work has shown that adding a custom attention layer after the original model (BERT) improves the model performance [tang2019progress, wood2020automated]. McCreery et al. [mccreery2019domain] double-fine-tuned the BERT model, in the sense that they first fine-tuned the model on a general dataset and then fine-tuned it on a medical-specific dataset. Another approach is to combine graph-based models with general LLMs, as in ChatGraph [shi2023chatgraph] that used ChatGPT to extract text information and apply it to a graph-based model that outperforms GPT models. CohortGPT [guan2023cohortgpt] used Chain-of-Thought prompting and knowledge graph to outperform few-shot ChatGPT and GPT-4.
Evaluation
Same as above, the F1-score is used for evaluating the model performance. Table 8 compares the performance of PubMedBERT [gu2021domain], BioGPT [luo2022biogpt], GPT-4 [liu2023deid, guan2023cohortgpt], ChatGPT [shi2023chatgraph, guan2023cohortgpt], and ChatGraph [shi2023chatgraph] on four benchmark datasets. It shows that GPT models perform better on the i2b2 dataset than on the OHSUMED dataset, possibly because its few-shot setting is not yet sufficiently good. Overall, few-shot GPT models outperform zero-shot ones, and the F1 scores tend to increase with more parameters.
A.5 Natural Language Inference
Task Description
Natural Language Inference (NLI) is a sentence-level task, involving two sentences: hypothesis ($H$) and premise ($P$). The task to to determine whether $H$ can be inferred from $P$. The outcome is one of the following three labels: (i) Entailment: the hypothesis can be inferred from the premise, or logically, $P\implies H$. (ii) Contradiction: the negation of the hypothesis can be inferred from the premise, or logically, $P\implies\neg H$. (iii) Neutral: all other cases, or logically, $\neg((P\implies H)\lor(P\implies\neg H))$.
Datasets and Models
The commonly used datasets, such as MedNLI [romanov2018lessons] and BioNLI [bastan2022bionli], are shown in Table 4 . Due to the scarcity of biomedical datasets, researchers also use general datasets (e.g. SNLI [bowman2015large] and MultiNLI [williams2017broad]). In task-specific fine-tuned models, similar to RE, [CLS] vector is often added to the text to show the overall information. The difference is that here we have two (instead of one) sentences, and so a [SEP] vector is also applied to indicate the separation of premise and hypothesis, yielding an overall input that looks like ‘[CLS], premise, [SEP], hypothesis’. For BERT-based models, a classification head is applied to the final [CLS] vector to predict the label. Kanakarajan et al. [raj2019saama] first pre-trained BioBERT [lee2020biobert] on MIMIC-III [johnson2016mimic] and then fine-tuned the model on MedNLI [romanov2018lessons]. Cengiz et al. [cengiz2019ku_ai] used the so-called two-stage sequential transfer learning method. They trained the BioBERT model on SNLI [bowman2015large] and MultiNLI [williams2017broad] first then fine-tuned it on the MedNLI dataset [romanov2018lessons]. The majority vote method is used to combine the prediction output of different trained models. Gu et al. [gu2023distilling] combined GPT-3.5 and PubMedBERT using knowledge distillation. They first fed GPT-3.5 original texts and asked it to distill the data, and the distilled data was subsequently used to further train the PubMedBERT model. Wang et al. [wang2023large] defined the self-questioning prompting (as discussed in Sec A.1) and utilize it in the MedNLI task.
Evaluation
The F1-score is used in the same manner as above. Table 9 compares the performance of task-specific fine-tuned models (e.g. BioBERT, ClinicalBERT, BlueBERT, and SciBER) on the benchmark MedNLI dataset [gu2023distilling], together with the performances of GPT-3.5 and GPT-3.5-Distillation. It shows that BioBERT [lee2020biobert] has the best performance, and the current general LLMs still need further improvement on natural language inference.
A.6 Semantic Textual Similarity
Task Description
Semantic Textual Similarity (STS) is similar to natural language inference (NLI). Their difference is that STS is a quantitative task, while NLI is a qualitative task. STS aims to produce a numerical value in a range (such as $[0,M]$) for any pair of sentences that indicates the degree of similarity, with the interpretation that 00 means two sentences are completely independent and $M$ means two sentences are completely correlated or equivalent.
Datasets and Models
As shown in Table 4, the ‘2019 n2c2/OHNLP’ [wang20202019], BIOSSES [souganciouglu2017biosses], and MedSTS [wang2020medsts] are widely-used benchmark datasets for STS. Since STS is a text-level task, [CLS] vector is naturally added to extract the overall information. Mutinda et al. [mutinda2020detecting] added a fully connected layer after the [CLS] vector as the architecture. Yang et al. [yang2020measurement] combined the representation of different models and added a fully connected layer. Wang et al. [wang2020evaluating] dropped the concept of [CLS] and used the Hierarchical Convolution (HConv) layer as the added last layer. Wang et al. [wang2023large] defined the self-questioning prompting (SQP) and utilize it on the BIOSSES dataset. Xiong et al. [xiong2020using] used the concatenation of character level, sentence level, and entity level representation (CSE-concate) to extract the information that is fed to the further added MLP layer.
For further training on a pre-trained model, there is a common understanding that the new dataset should have a similar distribution to that of the dataset for pre-training to improve the model performance. The iterative training is introduced for this purpose [wang2020learning, mahajan2020identification], where the phrase "iterative" refers to the following two iterative steps: a) It first freezes the model and computes the outputs from the dataset, then it chooses a subset of the dataset so the outputs of the subset have a similar distribution to that of the pre-training dataset. b) The model is further trained using the obtained subset obtained in Step a).
Evaluation
The Pearson correlation coefficient $r_{xy}$ is used for evaluation purposes. For $n$ pairs of data $\{x_{i},y_{i}\}_{i=1}^{n}$, we define $r_{xy}=\frac{\sum_{i=1}^{n}(x_{i}-\bar{x})(y_{i}-\bar{y})}{\sqrt{\sum_{i=1}^{n}(x_{i}-\bar{x})^{2}}\sqrt{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}}\,.$ Here, $\{x_{i}\}_{i=1}^{n}$ represents the actual similarity and $\{y_{i}\}_{i=1}^{n}$ represents the predicted similarity. Table 10 compares the performance of BERT [mutinda2020detecting, wang2023pre], ClinicalBERT [mutinda2020detecting, wang2023pre], XLNet [yang2020measurement], RoBERTa [yang2020measurement, wang2023pre], HConv-BERT [wang2020evaluating], BARD [wang2023large], GPT-3.5 [wang2023large], GPT-4 [wang2023large], BERT (CSE-concate) [xiong2020using], and ClinicalBERT (iterative training) [mahajan2020identification] on BIOSSES and ‘2019 n2c2/OHNLP’ datasets.
It shows that the task-specific fine-tuned model RoBERTa achieves the best results on the ‘2019 n2c2/OHNLP’ dataset. Although GPT-4 has the highest Pearson correlation coefficient of 91.60%, we note that, with significantly fewer model parameters, the task-specific fine-tuned model ClinicalBERT achieves a competitive result compared to GPT-4. The overall results indicate that further work on improving the current general LLMs is needed.
Table 10: The performance (sample Pearson correlation coefficient, shown in percentage (%), of semantic textual similarity.
Types Models BIOSSES 2019 n2c2/OHNLP
Task-specific FT Models BERT [devlin2018bert] 81.40 69.23
ClinicalBERT [alsentzer2019publicly] 91.23 83.20
XLNet [yang2020measurement] - 84.70
RoBERTa [yang2020measurement, van2020automets] 81.25 87.78
HConv-BERT [wang2020evaluating] - 79.40
BERT (CSE-concate) [xiong2020using] - 86.80
ClinicalBERT (iterative training) [alsentzer2019publicly] - 87.00
PubMedBERT [gu2021domain] 92.30 -
BioLinkBERT [yasunaga2022linkbert] 93.63 -
General LLMs ChatGPT [openai2023chatgpt] 43.75 -
BARD (zero-shot) [manyika2023overview] 57.60 -
BARD (few-shot) [manyika2023overview] 60.10 -
GPT-3.5 (zero-shot) [GPT-3] 87.86 -
GPT-3.5 (few-shot) [GPT-3] 91.94 -
GPT-4 (zero-shot) [gpt-4] 88.90 -
GPT-4 (few-shot) [gpt-4] 91.60 -
Table 11: The performance ((NDCG@$10$) on three information retrieval datasets, shown in percentage (%).
Types Models TREC-COVID NFCorpus BioASQ
Task-specific FT Models OpenAI cpt-text-S [neelakantan2022text, jin2023biocpt] 67.9 33.2 -
OpenAI cpt-text-M [neelakantan2022text, jin2023biocpt] 58.5 36.7 -
OpenAI cpt-text-L [neelakantan2022text, jin2023biocpt] 56.2 38.0 -
OpenAI cpt-text-XL [neelakantan2022text, jin2023biocpt] 64.9 40.7 -
BioCPT [lee2020biobert, raj2019saama] 70.9 35.5 55.3
General LLMs GTR-Base [google2023GTR] 53.9 30.8 27.1
GTR-Large [google2023GTR] 55.7 32.9 32.0
GTR-XL [google2023GTR] 58.4 34.3 31.7
GTR-XXL [google2023GTR] 50.1 34.2 32.4
ChatGPT [openai2023chatgpt] 76.7 35.6 -
ChatGPT [openai2023chatgpt] + GPT-4 [gpt-4] 85.5 38.5 -
A.7 Information Retrieval
Task Description
Information retrieval (IR) plays an important role in the clinical area. It is the process of retrieving relevant knowledge or information related to the query from a number of unstructured data. It satisfies one’s need for searching information, such as article recommendations and literature searches. IR in the biomedical domain contains multiple tasks, among which text summarization and text simplification are discussed in Sec B.1 and Sec B.2, respectively. Question answering is also an important task included in IR, previously discussed in Sec A.1. In this subsection, we will concentrate on the query-article relevance task, also known as the ranking task. Specifically, for a query $q$ and a dataset of articles $\{d_{i}\}_{i=1}^{n}$, we aim to find the most relevant $k$ articles $(d_{1}^{q},d_{2}^{q},\cdots,d_{k}^{q})$ where the degree of relevance is defined task-specifically.
Datasets and Models
Table 4 shows commonly used datasets, such as TREC-COVID [voorhees2021trec], NFCorpus [boteva2016full], and BioASQ [thakur2021beir] in the biomedical area. In addition, a general dataset, BEIR (Benchmarking IR) [thakur2021beir], is often included in the training step of biomedical IR. Jin et al.
[jin2023biocpt] developed a BERT-based model BioCPT that encodes the query and articles for the ranking task. They split the method into training, inference, and evaluation steps. In the training step, they introduce query-to-document loss and document-to-query loss to train the encoders for the query and articles. In the inference step, they concatenate the encoding of the query and its best-fit article $d_{1}^{q}$ together with $k-1$ non-relevant articles $(d_{2}^{q},d_{3}^{q},\cdots,d_{k}^{q})$ found by maximum inner product search (MIPS) to further train the model to rank $d_{1}^{q}$ at the top. In the evaluation step, for each input query $q$, the model evaluates over the whole dataset $\{d_{i}\}_{i=1}^{n}$ to find the best $k$ relevant articles by MIPS and used the model from the inference step to rank those $k$ articles.
Sun et al. [sun2023chatgpt] applied zero-shot ChatGPT to rank the most relevant documents without abnormal prompting. They also further used GPT-4 to re-rank the top 30 documents retrieved by ChatGPT. Abonizio et al. [abonizio2023inpars] introduced two LLM-based data augmentation methods, namely InPars and Promptagator, for IR. For InPars, they used GPT-3 and GPT-J to generate a new query for a randomly selected document. They used few-shot prompting which provides the model with good query examples or bad query examples. For Promptagator, the major difference is that a more dataset-specific prompting is applied. Ateia and Kruschwitz [ateia2023chatgpt] proposed a query expansion technique that expands the current query into a more comprehensive query, which consistently improves the performance of any successive tasks. It is done purely by the GPT model with regular instructional prompting. Similarly, Wang et al. [wang2023can] used ChatGPT to generate more refined Boolean queries for systematic reviews. They showed that ChatGPT is able to generate or refine queries with higher precision.
Evaluation
[…]
We compare the performance of Google’s Generalizable T5-based dense Retrievers (GTR) [ni2021large, jin2023biocpt], OpenAI’s cpt-text [neelakantan2022text, jin2023biocpt], BioCPT [jin2023biocpt], ChatGPT [sun2023chatgpt], and ChatGPT+GPT-4 [sun2023chatgpt] on the datasets we introduced before.
Table 11 shows that ChatGPT and GPT-4 have the best performance. We observe that increasing parameters does not necessarily improve the model performance; for example, GTR-XXL has a lower NDCG@10 score than those of other GTR models, but performs better. We also note that the performance of all models appears unstable across different data; for example, ChatGPT+GPT-4 has only a 38.5% score for NFCorpus, but has a score of 85.5% for TREC-COVID.
Appendix B Appendix: Generative Tasks
B.1 Text Summarization
Task Description
There are two types of text summarization: extractive and abstractive summarizations. Extractive summarization aims to find the most important sentences in the text, while omitting the redundant or irrelevant sentences. In contrast, abstractive summarization generates brand-new text that summarizes the given text. Therefore, the extractive approach is more related to token-level tasks, while the abstractive approach is more relevant to high-level tasks (e.g. text generation).
Datasets and Models
Table 4 shows the commonly used biomedical datasets PubMed [PubMed] and MentSum [sotudeh2022mentsum]. Some general datasets, such as GigaWord [napoles2012annotated], CL-SciSumm [yasunaga2019scisummnet], and S2ORC [lo2019s2orc], are also commonly used for text summarization.
For extractive summarization, the key task is to define some scoring system that scores all sentences and hence finds the most important ones. Moradi et al. [moradi2020deep] used a clustering-based method to summarize medical text. They vectorize the tokens of the text by BERT and cluster the sentence vector into $k$ clusters, and then they define an informativeness score that chooses one sentence from each cluster to form a summary. Moradi et al. [moradi2020summarization] used the graph-based model for summarization, where sentences are treated as nodes and relations as edges. The relations are measured by the cosine similarity of vectors representing the sentences. They then used different graph ranking algorithms to choose important sentences as a summary.
Du et al. [du2020biomedical] used purely the Transformer-based model as the scoring system. They i) tokenized the whole text with [CLS] and [SEP] augmented, ii) further augmented the corresponding sentence and token positions to each token, and iii) fed the whole vector into the model. A sigmoid layer is added to the model so the output is between zero and one, and the output is considered the score for each sentence. Since the Transformer automatically extracts relations of one sentence to others, there is no need for manually designing a score. Chen et al. [chen2020modified] also relied only on the Transformer to score the sentences. Different from Du et al. [du2020biomedical], they proposed the AlphaBERT that takes the character-level tokens as input,and the training is split into pre-training and fine-tuning.
McInerney et al. [mcinerney2020query] combined the summarization model with a query to generate specific summaries. Pang et al. [pang2022long] proposed a principled inference framework with top-down and bottom-up inference techniques to improve summarization models.
There are also studies on summarizing more than one piece of text at a time. They mainly relied on graph-based or Transformer-based models to extract relations between different texts [chen2021capturing]. Zhang et al. [zhang2023ds4dh] applied one-shot GPT-3.5, using dialogues and summaries from the same category as prompts to generate abstractive summarization. More studies are needed to qualitatively analyze the performance of GPT models on biomedical text summarization [shaib2023summarizing].
Evaluation
The standard Recall-Oriented Understudy for Gisting Evaluation (ROUGE) data[lin2004rouge], including ROUGE-1 and ROUGE-2, is used for evaluation. Table 12 compares the performance of BioBERT variations using different training sets [moradi2020summarization], rankings (PageRank and PPF) [moradi2020summarization], and added last layers [du2020biomedical] on PubMed. It shows that models fine-tuned on Pubmed or PMC data achieve improved performance. Differences in the ranking strategy do not correspond to appreciable differences in the model performance. Overall, current BERT models are far below the expectation of true text summarization, measured by the ROUGE-2 scores.
Table 12: The performance (ROUGE-1 & ROUGE-2) of the text summarization task, shown in percentage (%).
Types Models PubMed
Task-specific FT Models BioBERT (LSTM) [du2020biomedical] 35.82 & 17.15
BioBERTSum [du2020biomedical] 37.45 & 17.59
BioBERT (Pubmed + PPF) [moradi2020summarization] 74.21 & 32.88
BioBERT (PMC + PageRank) [moradi2020summarization] 75.82 & 34.01
BioBERT (Pubmed + PageRank) [moradi2020summarization] 76.09 & 34.38
BioBERT (Pubmed + PMC + PageRank) [moradi2020summarization] 76.34 & 34.67
GPT-3.5 (zero-shot) [GPT-3] 6.08 & -
GPT-3.5 (few-shot) [GPT-3] 23.20 & -
GPT-4 (zero-shot) [gpt-4] 39.97 & -
GPT-4 (few-shot) [gpt-4] 40.54 & -
Table 13: The performance (BLEU score) of the text simplification task on the MultiCochrane (English) dataset, shown in percentage (%). Higher is better.
Model MultiCochrane
GPT-3 (zero-shot) [GPT-3] 2.38
Flan-T5 (zero-shot) [joseph2023multilingual] 8.12
mT5 [joseph2023multilingual] 8.82
Flan-T5 (fine-tuned) [joseph2023multilingual] 8.70
GPT-3.5 (zero-shot) [GPT-3] 13.05
GPT-3.5 (few-shot) [GPT-3] 13.18
GPT-4 (zero-shot) [gpt-4] 12.00
GPT-4 (few-shot) [gpt-4] 13.12
Table 14: The performance (accuracy) of the text simplification on the AutoMeTS, shown in percentage (%). Higher is better for BLEU.
Model AutoMeTS
BERT [devlin2018bert] 62.40
RoBERTa [yang2020measurement, van2020automets] 53.28
XLNet [yang2020measurement] 46.20
GPT-2 [gpt-2] 49.00
BERT [devlin2018bert] + RoBERTa [yang2020measurement, van2020automets] + XLNet [yang2020measurement] + GPT-2 [gpt-2] 64.52
B.2 Text Simplification
Task Description
Text simplification is to generate a new text that recovers almost all the information of the original text while improving its readability. The task has potential in biomedical education, since one major characteristic of medical education is its opaque vocabulary [ondov2022survey, van2019evaluating].
Datasets and Models
Table 4 shows the commonly used text simplification datasets. Patel et al. [patel2022summarization] used NER to identify the medical terms. Lexical substitution is then applied to reduce the complexity of the text. Jeblick et al. [jeblick2022chatgpt] tested the performance of ChatGPT on simplifying self-collected radiology reports. They tried different prompting texts and found that the prompt ‘Explain this medical report to a child using simple language’ performs the best. However, by evaluating from a) factual correctness, b) completeness, and c) harmfulness, they concluded that GPT models may generate harmful texts which is unacceptable in the medical domain. Joseph et al. [joseph2023multilingual] evaluated the performance of zero-shot GPT-3 and Flan-T5 on MultiCochrane. They also fine-tuned mT5 [xue2020mt5] and Flan-T5 on this dataset. Yang et al. [yang2023data] introduced a data augmentation method for text simplification based on LLMs. For text without its simplified counterpart, they used GPT-3 to generate multiple choices for simplified text, and further trained a BERT model as the score to choose the best one as the simplification. Van et al. [van2020automets] transferred the simplification task into a prediction task. They assume they have the original text $(d_{1},d_{2},\cdots,d_{n})$ and a unfinished simplified text $(s_{1},s_{2},\cdots,s_{i})$. The task is to predict the next token $s_{i+1}$. In particular, at each time, the model receive $(d_{1},d_{2},\cdots,d_{n},s_{1},s_{2},\cdots,s_{i})$ as input and predict $s_{i+1}$. They used different models like BERT, RoBERTa, XLNet, and GPT-2. They also tried to combine the predicted token $s_{i+1}$ of all four models to outperform any of them.
Evaluation
For evaluation, we use the bilingual evaluation understudy (BLEU) score [papineni2002bleu]. We compare the performance of GPT-3 [joseph2023multilingual], Flan-T5 [joseph2023multilingual], and mT5 [joseph2023multilingual] on MultiCochrane (English) [joseph2023multilingual] (please see Table B.1). For AutoMeTS dataset [van2020automets], we use the accuracy of the next token $s_{i+1}$ as we discussed and we compare the performance of BERT, RoBERTa [van2020automets], XLNet [van2020automets], GPT-2 [van2020automets], and their combination [van2020automets] (please see Table B.1). The citations in this paragraph correspond to the sources that provide data on the performance of the models. For MultiCochrane, T5 models have much better performance than other baselines. For AutoMeTS, a combination of different models actually outperforms any of them, but overall, none of those models have a high accuracy or BLEU score.
B.3 Text Generation
Task Description
Text generation is a broad task that includes or relates to multiple specific tasks such as question answering (Sec A.1) and text summarization (Sec B.1). Here we focus on data-to-text tasks with open (instead of well-defined) answers, which involves taking structured data (e.g. a table) and producing text as output that describes these data. For example, generating patient clinic letters, radiology reports, and medical notes [liu2023utility].
Datasets, Models, and Discussion
Yermakov et al. [yermakov2021biomedical] introduced a new dataset BioLeaflets and evaluated the performance of multiple LLMs on the data-to-text generation task. They found that T5 [T5] and BARD [manyika2023overview] achieve the best performance for this task. However, multiple questions remain. Current LLMs may generate texts with typos, hallucinations, and repetitious words. Also, LLMs are not mature enough to produce coherent long text so far. Ranjit et al. [ranjit2023retrieval] studied the task of generating reports for chest X-ray images. They also investigated and discussed the hallucinations that occurred in the GPT-generated texts. There are concerns about using model-generated pseudo-text for attacking since humans without expert knowledge cannot easily see the factual errors in the generated text. Rodriguez et al. [rodriguez2022cross] works on preventing attacks in the biomedical domain.